### v1

# AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

Abstract
Multimodal Chain-of-Thought (CoT) has emerged as a powerful technique for enhancing the vision-language reasoning with interleaved information. However, existing methods often rely on simplistic heuristics for constructing interleaved CoT, typically depending on attention maps, which our empirical analysis reveals can be unreliable. What’s more, the shortcomings of their passive and purposeless selection strategies and their arbitrary triggering mechanisms in capturing the model’s cognitive need for information are further amplified. In this paper, we propose AIMCoT, an Active Information-driven Multi-modal Chain-of-Thought framework that addresses these fundamental limitations. AIMCoT introduces three synergistic components: (1) Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity imbalance, thereby producing more reliable attention maps as a foundation. (2) Active Visual Probing (AVP), which replaces passive selection with a proactive, goal-oriented strategy grounded in information theory to select image regions that help answer the questions maximally. (3) Dynamic Attention-shifting Trigger (DAT), which intelligently determines the optimal moments to insert visual information by monitoring the model’s text-to-vision attention shifts. Extensive experiments on three challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods across different settings. By actively foraging for information and dynamically structuring its reasoning process, AIMCoT represents a critical step towards more robust, effective, and human-like multimodal reasoning. Our code is available at https://anonymous.4open.science/r/AIMCoT.
1 Introduction
The advent of Chain-of-Thought (CoT) prompting has marked a significant milestone in the reasoning capabilities of Large Language Models (LLMs) Wei et al. (2022); Wang et al. (2022); Zhang et al. (2022); Suzgun et al. (2022); Li et al. (2025); Yao et al. (2023); Besta et al. (2024), enabling them to deconstruct complex problems into a series of intermediate, interpretable steps. This paradigm has been naturally extended to Vision-Language Models (VLMs), where early efforts Mitra et al. (2024); Zheng et al. (2023); Lei et al. (2024); Zhang et al. (2023) focused on generating text-only rationales to articulate the model’s reasoning process over visual inputs. A pivotal advancement in this domain was the introduction of Interleaved-modal Chain-of-Thought Gao et al. (2025), which pioneered the direct integration of visual patches into the reasoning chain. By pairing textual rationales with corresponding image regions, the interleaved CoT demonstrated a superior ability to ground language in visual evidence, setting a new standard for multimodal reasoning.
However, the efficacy of existing research Ge et al. (2025); Gao et al. (2025) is fundamentally predicated on the reliability of their underlying mechanisms for selecting and integrating visual information. These methods typically rely on a passive, attention-driven strategy: they select the Top-K regions with the highest attention scores and insert them at predefined moments, e.g., the appearance of a newline character, a practice with little theoretical or empirical justification for its timing. This reliance exposes critical vulnerabilities that limit their full potential. Our empirical analysis reveals that high-attention regions are often redundant or, more alarmingly, fail to capture the most crucial visual details, especially when a granularity mismatch exists between the textual query and the visual evidence. This raises three fundamental questions: (1) What information source can reliably identify truly salient visual regions? (2) How can we accurately select useful regions in a proactive and purposeful manner rather than passively relying on potentially unreliable attention scores? (3) When is the optimal moment to insert visual evidence into the reasoning process?
In this work, we postulate that a more robust multimodal reasoning framework requires a shift from passive attention-following to an active, information-seeking paradigm. We propose AIMCoT: Active Information-driven Multi-modal Chain-of-Thought, a novel framework that directly addresses the limitations of prior work, reframing the selection of visual evidence from a passive, attention-based retrieval task to an active, goal-oriented probing process. Inspired by the principles of information foraging Pirolli & Card (1999); Broadbent (2013), AIMCoT operates on the premise that the most valuable visual rationale is one that maximally reduces the model’s uncertainty about the subsequent step in its reasoning chain. Rather than simply asking "Where is the model looking?", we prompt the model to actively ask "Which piece of visual information will be most helpful for me to see right now?". This shift is realized through three synergistic components:
  1. 1.
    Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity disparity to improve the reliability of attention map for better identifying salient regions by generating a context-aware description of the image.
  2. 2.
    Active Visual Probing (AVP), which implements a proactive, goal-oriented selection strategy grounded in information theory, selecting a set of visual regions that provide the highest possible information gain for the task at hand.
  3. 3.
    Dynamic Attention-shifting Trigger (DAT), an intelligent triggering mechanism that carefully captures the critical moments when model’s cognitive focus shifts significantly from text to vision, and inserts visual information precisely.
Our contributions are as follows:
  • We introduce AIMCoT, a novel training-free framework that reframes the construction of multimodal CoT as an active information-foraging process, moving beyond the limitations of static, passive, and purposeless region selection.
  • We propose a system comprised of three complementary methods (CAG, AVP, DAT) that collectively enable VLMs to proactively forage for informative visual evidence and strategically integrate it into their reasoning process at the detected critical moments.
  • We present a comprehensive set of empirically-grounded motivations that inspire the designs within our AIMCoT framework. Furthermore, we provide substantial theoretical analysis and design targeted experiments to explore the properties of these designs, covering key aspects such as the deployability of AIMCoT, the interplay between the CAG and AVP modules, and the necessity of incorporating an exploratory candidate pool.
  • Through extensive experiments conducted on two backbones (Chameleon-7B and Qwen2-VL-7B) across challenging benchmarks including M3CoT, ScienceQA, and LLaVA-W, we demonstrate that AIMCoT significantly outperforms state-of-the-art baselines and advances the frontier of vision-language reasoning.
2 Related Work
The success of CoT prompting in LLMs has naturally extended to VLMs, aiming to make their reasoning processes explicit and interpretable. Early efforts focused on generating text-only rationales. MMCoT Zhang et al. (2023) first generates a rationale from the input and then uses this rationale along with the original multimodal data to infer the final answer. CCoT Mitra et al. (2024) prompts the VLM to generate an intermediate scene graph to structure its understanding. DDCoT Zheng et al. (2023) decomposes complex problems into simpler sub-questions and leverages external models to fill information gaps. Alternatively, SCAFFOLD Lei et al. (2024) overlays a coordinate grid on the image, enabling the VLM to reference spatial regions explicitly in its textual reasoning.
A pivotal advancement is the introduction of interleaved-modal CoT, which integrates visual evidence directly into the reasoning chain. The leading approach, ICoT Gao et al. (2025), selects the top-K regions from the VLM’s attention map and inserts them at predefined moments. MRFD Ge et al. (2025) meticulously selects the salient visual regions to foster more reliable reasoning, alleviating VLM hallucination. However, our analysis reveals that such passive, attention-driven strategies are fundamentally limited by the unreliability of raw attention maps, especially during text-vision granularity mismatches, and their arbitrary insertion points fail to capture the model’s dynamic cognitive needs. Our proposed AIMCoT addresses these critical vulnerabilities by shifting the paradigm from passive selection to active information foraging, employing goal-oriented visual probing and dynamic, attention-shift-triggered integration.
3 Motivation
3.1 Identifying the Reliability of the Attention Map
The utilization of the attention map in multimodal learning has been explored by recent research Gao et al. (2025); Ge et al. (2025); Xie et al. (2022); Wang et al. (2023); Liu et al. (2022), inspiring us to also take it as a source from which the salient visual regions are selected. However, considering their strong dependence on attention map, we question its reliability by posing two fundamental questions:
  1. (1)
    Do all significant regions on the attention map help the VLMs answer questions correctly?
  2. (2)
    Does the attention map comprehensively capture all visual regions that are instrumental to the VLM’s correct prediction?
Figure 1: The images of the 22nd question on LLaVA-W benchmark, which is a close-up photo of a meal at ICHIRAN. The left and right figures are respectively the original image and the first three sets of regions selected by the Top-K strategy (red, purple, and blue, respectively). A detailed explanation is shown in Appendix B.
To acquire a holistic comprehension of these two questions, our empirical analysis is conducted on the popular Visual Question Answering (VQA) benchmark, LLaVA-W Liu et al. (2024), employing ICoT Gao et al. (2025) as a baseline model, which selects the Top-K regions from the attention map to construct a text-vision interleaved CoT. Chameleon-7B Team (2024) serves as the backbone.
Table 1: Performance degradation of the baseline model (ICoT, 0-shot) when the Top $K_{mask}$ regions on the attention map are masked.
$K_{mask}$ 0 1 5 10 20
Degradation 0% 0.26% 1.43% 3.93% 2.44%
We quantitatively investigate the role of high-attention regions by masking the top-$K_{mask}$ regions identified by ICoT (0-shot). On LLaVA-W, this masking leads to a minor performance drop of just 3.93% (top 10) and 2.44% (top 20), as shown in Table 1. Notably, performance increases when the mask is expanded from the top 10 to the top 20 regions. These results strongly suggest that not all high-attention regions contribute significantly to the model’s prediction: some have negligible impact or even introduce detrimental signals.
In response to the second question, we manually inspect the content of high-attention regions to understand their role in VLM’s predictions. In a challenging LLaVA-W example (Figure 1), the answer lies in a small area, i.e., the inner rim of a ramen bowl, among rich visual information. The top two most attended image patch sets entirely miss this crucial detail, with only a negligible portion appearing in the third set. This indicates a potential misalignment between high-attention regions and key visual information, particularly when the text-vision granularity disparity is significant.
3.2 Foraging for the Most Information to Guide Region Selection
The analysis in Section 3.1 empirically reveals that the efficacy of the attention-driven Top-K strategy is notably constrained, particularly when faced with significant text-vision granularity disparities. Furthermore, fundamentally, since attention maps merely reflect token correlations, a static approach without an explicit goal like Top-K, which relies solely on attention scores, is inherently suboptimal. This insight directly motivates us to explore a selection method that is more proactive and purposeful.
Existing recognized research Pirolli & Card (1999); Broadbent (2013); Oaksford & Chater (1994); Friston (2010) supports that people, when possible, will maximize their rate of gaining valuable information, as it yields more useful information per unit cost. Inspired by this, we explore providing the VLM with image regions that yield the highest information gain, thereby maximally reducing the model’s uncertainty in answering a given question. Intuitively, the model’s uncertainty can be quantified by the entropy of its probability distribution over the vocabulary given the current context, while the information gain of an image region is measured based on model’s entropy when the region is included in the context. The definitions are meticulously detailed in Section 4.3.
Figure 2: The visualization of regions selected by the information gain-guided strategy.
To continue with the example in Section 3.1, Figure 2 illustrates the top three regions ranked by the information gain for VLM. Evidently, in contrast to the regions selected via Top-K shown in Figure 1, the information gain-based selection accurately guides the VLM to first focus on the inner rim of the bowl (red), where the critical information is contained. Although the VLM does not yield the final answer in this region, we note that this indicates a correct line of reasoning, as the ground truth is situated in a highly similar area (a nearby location also on the inner wall of the bowl). Subsequently, the region ranked third (blue) precisely encompasses a large portion of the restaurant’s name, which is just the answer to the question. This suggests that even for a challenging case where the text-vision granularity is highly disparate, information gain serves as a better foundation for region selection.
3.3 Inserting Visual Information at Key Moments
Although existing research Gao et al. (2025) has attempted to construct interleaved CoT, they often ignore a reasonable triggering mechanism to capture the critical moment of inserting visual information. For example, ICoT Gao et al. (2025) simply uses a newline character as a trigger signal. This motivates an in-depth investigation into reliable indicators for identifying these key moments for inserting visual content.
To this end, we conduct a comparative case study on the LLaVA-W benchmark, with ROUGE-L as the evaluation metric. We collect all predictions generated by the baseline model (ICoT) and partition them into high- and low-scoring groups based on ROUGE-L scores. Given that the model’s attention map has served as an important foundation for existing works Gao et al. (2025); Ge et al. (2025); Xie et al. (2022); Wang et al. (2023); Liu et al. (2022), which is not only readily accessible but also effectively reflects the model’s aggregate attention to the textual and visual parts of the input, we are inspired to monitor the attention shifts between these two modalities throughout the prediction process. The experiments are detailed in Appendix F, from which two key observations are revealed: (1) Correlation analysis: inserting visual data precisely when the model’s attention pivots towards the visual modality is strongly correlated with higher scores and (2) Group analysis: this phenomenon further serves as a crucial characteristic that distinguishes high-scoring from low-scoring outputs.
4 AIMCoT
In this section, we begin by briefly reviewing the background of multimodal learning. Then, we detail the proposed methods motivated by the following key insights derived from Section 3:
  1. (1)
    First, the high-scoring attention regions are not always beneficial for question-answering, and crucial visual evidence can be missed particularly in cases of text-vision granularity mismatch.
  2. (2)
    Second, for a given set of candidate regions, selection based on information gain significantly outperforms the conventional attention-driven Top-K method.
  3. (3)
    Finally, capturing the critical moments to insert visual information improves the construction of multimodal CoT, and text-to-vision attention shifts serve as an important indicator.
Accordingly, we propose AIMCoT, which encompasses three key methods: (1) Context-enhanced Attention-map Generation (CAG), which generates a fine-grained description for the image to alleviate text-vision disparity. (2) Active Visual Probing (AVP), which proactively and purposefully selects regions that are most helpful for answering the question. (3) Dynamic Attention-shifting Trigger (DAT), which triggers vision insertion into CoT when the model’s cognitive focus is significantly shifted from text to vision.
4.1 Preliminaries
Vision-Language Model. A VLM typically fuses a vision encoder for preprocessing visual input and a generative language model, which jointly enable it to respond in a human-like manner as follows:
$answer=\text{VLM}(I,x),$ (1)
where $I$ and $x$ are the image and query, respectively.
Multimodal CoT. Compared to the direct response shown in Equation 1, multimodal CoT encourages the VLM to output the thought process, called rationales, before outputting the final answer. Existing work Gao et al. (2025) has been devoted to generalizing rationales from pure text to text-vision interleaved form.
4.2 Context-enhanced Attention-map Generation (CAG)
In this component, the VLM is prompted to carefully generate an explanatory description of the given image within the question before the process of VQA, with the explicit goal of helping a potential respondent correctly answer the question. In this process, the VLM acts as a facilitator who interprets the image in the context of the given question to guide the respondent’s thought process. Formally, it is expressed as follows:
$\mathcal{D}_{CAG}=\text{VLM}(I,x,\mathcal{P}_{CAG}),$ (2)
where $\text{VLM},I,x$ are the used VLM, the given image and question, respectively. $\mathcal{P}_{CAG}$ is the prompt provided for the model to generate the description. Then, to compensate for the sparsity of textual information within the context, the generated description $\mathcal{D}_{CAG}$ is concatenated to the question $x$ as follows:
$x^{\prime}=\text{concat}(x,\mathcal{D}_{CAG}).$ (3)
By enhancing the context to compensate for the sparsity of textual information, the disparity in text-vision granularity is effectively mitigated, unlocking the potential for the attention map to serve as a more reliable indicator of task-relevant regions. We provide an example in Appendix G.1 that details the template of $\mathcal{P}_{CAG}$, the entire process of CAG, and how the final attention map $A^{\prime}$ is generated, which is used in the next stage.
4.3 Active Visual Probing (AVP)
Built upon the theoretical foundation of information gain, AVP is designed to select the crucial regions from a set of candidate visual regions. Although AVP consists of three steps elaborated upon as follows, in terms of complexity, our provided analysis and empirical results in Appendices J and L suggest that the introduction of AVP still enables AIMCoT to strike a good balance between deployability and superior performance. A visualization of AVP is shown in Figure 3.
Figure 3: An overview of our AVP module, which iteratively selects $K$ most informative regions from a diversified candidate set $C$ to build an interleaved CoT that fosters vision-language reasoning.
Diversified Set of Candidate Regions Construction. Based on the first two insights summarized in Section 4, an inference can be obtained: relying solely on attention maps as the source may prevent the optimal region selection. Motivated by this, we propose to diversify the source of the candidate regions, thereby reducing the model’s dependency on attention maps alone. Specifically, we not only construct an attention-driven candidate set, $C_{attn}$, by selecting $N$ regions with the highest attention scores ($N\in\mathds{R}$), but also generate an exploratory candidate set, $C_{exp}$, by sampling $M$ grid regions uniformly at random from the input image ($M\in\mathds{R}$). Our empirical analysis in Appendix I shows that the incorporation of exploratory set $C_{exp}$ provides substantial salient visual regions for the VLM. Furthermore, we study the impact of introducing $C_{exp}$ on the performance of AIMCoT, and further compare random sampling, selective search Uijlings et al. (2013), and FastSAM Zhao et al. (2023) as methods for constructing $C_{exp}$ in Appendix M.
Formally, the process can be expressed as follows:
$C_{attn}$ $=\{R_{1},R_{2},\cdots,R_{N}\},~~~\text{s.t. $R_{i}$=Top-$i$ Region from the attention map $A^{\prime}$},$ (4)
$C_{exp}$ $=\{R_{N+1},R_{N+2},\cdots,R_{N+M}\},~~~~~C=C_{attn}\cup C_{exp},$ (5)
where $C_{attn},C_{exp}$, and $C$ are the attention-driven, exploratory, and total candidate sets, respectively. Subsequently, the goal is to select the most informative $K$ regions from $C$ ($K<|C|=N+M$).
Quantification of Information Gain. An ideal region of input image is the one that is most informative and critical for answering the question, i.e., the one that minimizes the model’s predictive uncertainty. Based on this idea, the region selection problem is transformed into a sequential information gain maximization task. To formalize this objective, we propose the concept of basic uncertainty $U_{B}$ defined as follows:
$U_{B}=H(Y|I,x,y_{<t})=-\sum_{y\in V}P(y|I,x,y_{<t})\text{log}_{2}P(y|I,x,y_{<t}),$ (6)
where $I,x$ and $y_{<t}$ are the input image, question and the predicted tokens, respectively. $P(y|I,x,y_{<t})$ is model’s probability distribution across the vocabulary $V$ in current context without the introduction of any region $R_{i}\in C$. It is seen that $U_{B}$ takes the form of entropy, thereby inherently capturing the model’s uncertainty when predicting the next token $t$ without incorporating any $R_{i}\in C$.
Similarly, the conditional uncertainty $U_{C,i}$, which illustrates the model’s uncertainty when an arbitrary region $R_{i}\in C$ is included into the context, is proposed and defined as follows:
$U_{C,i}=H(Y|I,x,y_{<t},R_{i})=-\sum_{y\in V}P(y|I,x,y_{<t},R_{i})\text{log}_{2}P(y|I,x,y_{<t},R_{i}),$ (7)
where $P(y|I,x,y_{<t},R_{i})$ is the model’s probability distribution across the vocabulary $V$ in the context combined with the region $R_{i}$. Eventually, the aforementioned conceptual formulation logically motivates the definition of our final key metric: the information gain of region $R_{i}$, which is formally defined as follows:
$IG(\{R_{i}\})=U_{B}-U_{C,i},~~~i=1,2,\cdots,N+M.$ (8)
Intuitively, $IG(\{R_{i}\})$ quantitatively characterizes how incorporating region $R_{i}$ reduces the model’s uncertainty for the subsequent token prediction.
Optimal Region Selection. To achieve the target of region selection, we formalize the problem as follows:
$\text{Maximize}~~F(S)=IG(S),~~~\text{s.t.}~S\subset C,|S|=K,$ (9)
where $S$ is the optimal set of selected regions and $K$ is the size of $S$ serving as a hyper-parameter. Motivated by two crucial insights, we propose a greedy algorithm detailed in Algorithm 1 to solve this problem: (1) According to recognized works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011), the greedy algorithm is a well-established and widely-used method for maximizing functions, especially for maximizing those that exhibit a tendency towards submodularity, even if they are not submodular theoretically; (2) Our experiments in Appendix H suggest that the information gain function $F$ empirically exhibits significant approximate submodularity.
Specifically, the selection is an iterative process consisting of $K$ steps. At each step, AVP proactively selects the most informative and not-yet-chosen region from the candidate set $C$ with the explicit goal of minimizing the uncertainty of the model’s answer to the question. The chosen region is then added to an intermediate set $R^{*}$. Notably, a key merit of AVP is its ability to bypass regions that, despite high attention scores, exhibit strong informational overlap with regions that have been selected. After $K$ rounds, the collection of all selected regions in $R^{*}$ constitutes the final optimal selection $S$. Substantial results in Appendices J.2 and L suggest that our framework’s average inference time is no more than 1.36 times that of an efficient baseline ICoT, and the framework scales efficiently with larger values of $K$ and $N_{C}=|C|$, respectively.
4.4 Dynamic Attention-shifting Trigger (DAT)
As motivated by the key observation in Section 3.3, it is vital to appropriately time the insertion of visual information when constructing a multimodal CoT, and furthermore, the shift of attention from the textual to the visual context serves as a crucial indicator. Motivated by these insights, we propose the DAT mechanism, which systematically evaluates the model’s attention scores on the visual context at every token $t$ generation step formulated as follows:
$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},$ (10)
where $C_{visual}$ is the visual information within the context; $\bar{a}_{t,i}$ is the average attention score of token $t$ towards the visual token with index $i$ across the last $N_{L}$ VLM layers. Drawing inspiration from the NLP community Jawahar et al. (2019); Tenney et al. (2019); Vig & Belinkov (2019), we restrict our focus to the model’s final layers, as they are responsible for capturing high-level semantic information, including semantic roles and coreference relations, where the shifting signal is presumed to be more reliable. In our implementation, we use the last 3 layers by default ($N_{L}=3$). Then, the shift of attention is formalized as follows:
$\Delta A_{visual}(t)=A_{visual}(t)-A_{visual}(t-1),$ (11)
which quantifies the model’s attention shift towards the visual context between generating the current token and the preceding one. Eventually, a hyper-parameter $\delta\in\mathds{R}$ is employed to delineate the point at which the attention shift $\Delta A_{visual}(t)$ is substantial enough to activate the AVP to insert essential visual information in the multimodal CoT. We detail a sensitivity analysis of the threshold $\delta$ with an emphasis on its impact on the frequency of triggering AVP and the performance of AIMCoT in Appendix K.
5 Experiments
5.1 Benchmarks and Baselines
In this study, we evaluate AIMCoT on three popular and challenging VQA benchmarks, including M3CoT Chen et al. (2024), ScienceQA Saikh et al. (2022), and LLaVA-Bench In-the-Wild (LLaVA-W) Liu et al. (2024). We provide detailed introductions in Appendix C.
To evaluate the performance of AIMCoT, we introduce the vanilla VLM w/o CoT (No-CoT) and a range of state-of-the-art methods as baseline models, including DDCoT Zheng et al. (2023), MMCoT Zhang et al. (2023), CCoT Mitra et al. (2024), and SCAFFOLD Lei et al. (2024), which generate text-only rationales. Furthermore, ICoT Gao et al. (2025), which constructs interleaved-modal CoT, is considered as well. The detailed introduction to them is listed in Appendix D. In presenting the results, we directly cite the performance reported in existing works where applicable.
5.2 Implementation Details
We implement AIMCoT and the baselines on Chameleon-7B Team (2024) and Qwen2-VL-7B-Instruct Wang et al. (2024) in two settings (both 0- and 1-shot), which aligns with the recent leading research Gao et al. (2025). The experiments are conducted on A6000 GPUs. The hyper-parameter settings for AIMCoT and the reproducibility statement are meticulously listed in Appendix G.2 and Section 7, respectively.
5.3 Performance Comparison
We evaluate the performance of AIMCoT against the state-of-the-art (SOTA) methods. The results shown in Table 2 clearly demonstrate the superiority of our proposed AIMCoT as it significantly outperforms all the baseline models under both 0- and 1-shot settings across all the datasets.
Specifically, AIMCoT surpasses all baselines that generate text-only rationales, confirming the efficacy of integrating salient visual information directly into CoT. When compared to ICoT, which also produces interleaved text-vision CoT, AIMCoT’s superior performance underscores the importance of our three key contributions: (1) a more reliable attention map as a foundation, (2) a proactive, goal-oriented mechanism for image region selection, and (3) an intelligent trigger for inserting visual information at critical moments.
Crucially, AIMCoT’s advantage is most pronounced on the open-ended LLaVA-W benchmark and in the 0-shot setting, which better simulate complex, real-world scenarios where the model must rely solely on its internal knowledge and reasoning. By emulating what can be seen as a more human-like cognitive process, AIMCoT unlocks the VLM’s foundational reasoning capabilities, enabling robust performance in novel and challenging situations.
Table 2: Performance comparison results on three widely-used benchmarks. The best performances are shown in bold. The metric for experiments on M3CoT and ScienceQA is Accuracy (ACC.), while on LLaVA-W, the metric ROUGE-L is adopted.
Backbone Method M3CoT (ACC.) ScienceQA (ACC.) LLaVA-W (ROUGE-L)
0-shot 1-shot 0-shot 1-shot 0-shot 1-shot
Chameleon-7B No-CoT 29.1 28.4 47.7 48.5 13.1 23.9
DDCoT 28.6 29.8 49.8 49.2 20.2 23.1
MMCoT 28.5 30.6 49.0 50.7 20.4 20.6
CCoT 29.4 31.4 50.2 51.3 22.1 24.5
SCAFFOLD 29.6 31.1 48.5 47.5 21.7 24.7
ICoT 29.8 32.3 51.0 53.4 25.2 27.6
AIMCoT (Ours) 31.4 32.8 53.1 54.5 29.8 32.0
Improvement 5.50% 1.47% 4.08% 2.04% 18.25% 15.94%
Qwen2-VL-7B No-CoT 43.6 45.4 56.3 64.4 32.7 33.5
MMCoT 40.1 42.5 51.3 58.3 30.7 31.4
CCoT 43.3 44.1 56.4 63.8 29.4 33.9
DDCoT 42.6 45.7 55.2 64.9 31.2 32.8
SCAFFOLD 41.7 44.9 53.7 62.5 31.8 33.1
ICoT 44.1 46.0 56.8 65.4 34.2 35.7
AIMCoT (Ours) 44.7 46.6 57.4 66.3 36.3 37.3
Improvement 1.4% 1.3% 1.1% 1.3% 6.2% 4.5%
5.4 Ablation Study
In this section, we conduct a series of ablation studies to verify the efficacy of each component within AIMCoT. The details of settings are as follows:
  • In w/o CAG, the VLM is directly prompted with the raw question $x$ and the paired image $I$;
  • In w/o AVP, the AVP is replaced by the attention-driven Top-K strategy by following existing works Gao et al. (2025); Ge et al. (2025), which selects the regions with Top-K attention scores on the model’s attention map;
  • In w/o DAT, following existing research Gao et al. (2025), the insertion of visual information in CoT is triggered when the model outputs the signal token, which is a line break by default.
Ablation results in Table 3 validate the contributions of our core components. First, CAG provides essential context enhancement, proving crucial for generating high-quality CoT, particularly when text queries are sparse (M3CoT, LLaVA-W). Second, the consistent, significant performance drop when replacing AVP with a Top-K baseline underscores the substantial superiority of our proactive, information-oriented method. Furthermore, the removal of DAT also results in a considerable performance decline, highlighting the critical importance of the timing of image insertion.
Table 3: Ablation study of AIMCoT conducted on Chameleon-7B under 0-shot setting.
Dataset AIMCoT w/o CAG w/o AVP w/o DAT
M3CoT (ACC.) 31.4 30.5 (-0.9) 30.6 (-0.8) 30.8 (-0.6)
ScienceQA (ACC.) 53.1 52.8 (-0.3) 52.3 (-0.8) 52.7 (-0.4)
LLaVA-W (ROUGE-L) 29.8 26.8 (-3.0) 26.2 (-3.6) 27.3 (-2.5)
The performance gains from AIMCoT are most pronounced on LLaVA-W, a challenging benchmark requiring open-ended generation. This large improvement starkly demonstrates our model’s advanced capability to comprehend intricate multimodal information and tackle demanding, unconstrained tasks.
Table 4: Ablation study of the baseline model (BM) on Chameleon-7B under 0-shot setting.
Dataset BM BM w/ CAG BM w/ AVP BM w/ CAG, AVP
M3CoT (ACC.) 29.8 30.3 (+0.5) 30.2 (+0.4) 30.8 (+1.0)
ScienceQA (ACC.) 51.0 52.0 (+1.0) 51.9 (+0.9) 52.7 (+1.7)
LLaVA-W (ROUGE-L) 25.2 25.8 (+0.6) 26.4 (+1.2) 27.3 (+2.1)
5.5 In-depth Analysis: the Interplay between CAG and AVP
Our proposed components collaborate organically to foster the construction of multimodal CoT: as a preceding module, CAG enriches the context to benefit the construction of the candidate set; subsequently, AVP triggered by DAT proactively selects the most salient regions from the candidate set.
In this section, we investigate the interaction between CAG and AVP via an ablation study. Starting from the baseline ICoT model (BM), i.e., AIMCoT stripped of all proposed modules, we sequentially add CAG and then AVP. Table 4 presents the results, from which we derive the following insights:
  • The compatibility between CAG and AVP: both CAG and AVP individually provide significant gains, but their combination synergistically improves the construction of the interleaved CoT.
  • The consistent superiority of AVP over Top-K selection: AVP consistently and significantly outperforms the standard Top-K selection method for choosing image regions, both with and without the presence of the CAG module.
  • The interplay between CAG and AVP: The average performance improvement of AVP over Top-K selection increases from 2.62% to 2.94% when CAG is introduced. This suggests CAG enhances the source attention map, providing a more reliable set of candidate regions and thereby unlocking AVP’s full potential to select the most salient visual evidence.
6 Conclusion
In this paper, we propose AIMCoT, a novel framework that reframes the construction of interleaved-modal CoT as an active, information-foraging process, addressing the limitations in existing methods, which often rely on passive, heuristic-driven mechanisms for selecting and inserting visual information at suboptimal moments. Our extensive experiments on three popular and challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods in both 0- and 1-shot settings (up to 18%). By dynamically structuring its reasoning and actively seeking the most informative visual cues, AIMCoT achieves a more proactive, goal-oriented, and human-like approach to vision-language reasoning.
Despite the strong performance, AIMCoT presents avenues for future exploration. The AVP module, while highly effective and optimized, introduces a slight computational overhead compared to simpler attention-based selection. Future work could explore lightweight, learnable policies for region selection to further enhance its deployability. We also plan to extend our evaluation to a broader range of VLM architectures and more complex, long-form reasoning tasks to further probe the generalizability and limits of our active information-seeking paradigm.
7 Reproducibility Statement
To ensure the reproducibility of our work, we provide the complete source code for our AIMCoT framework and all experiments in the supplementary materials, accessible via an anonymized link: https://anonymous.4open.science/r/AIMCoT. The architectural details and theoretical underpinnings of our proposed components, including Context-enhanced Attention-map Generation (CAG), Active Visual Probing (AVP), and Dynamic Attention-shifting Trigger (DAT), are thoroughly described in Section 4 of the main paper. The specific greedy algorithm employed by AVP is detailed in Algorithm 1 (Appendix E). All datasets used in our evaluation are publicly available benchmarks, as detailed in Section 5.1 and Appendix C. We provide a comprehensive list of all hyper-parameter settings used to achieve the reported results for each benchmark in Appendix G.2 (Table 5). Furthermore, extensive ablation studies (Section 5.4), and in-depth analyses (Section 5.5, Appendices F,H, I, K, and L), are provided to allow for a complete replication of our findings.
Appendix A LLM Usage Statement
In the preparation of this manuscript, we utilized a large language model (LLM) as a general-purpose writing assistance tool. The primary uses of the LLM were twofold:
  • Grammatical Correction: The LLM was employed to proofread the manuscript for grammatical errors, spelling mistakes, and awkward phrasing. This helped improve the overall clarity and readability of our paper.
  • Text Condensation: To adhere to the page limits of the conference, we used the LLM to help condense and rephrase certain paragraphs and sentences.
All suggestions provided by the LLM were carefully reviewed, critically evaluated, and manually edited by the authors to ensure that the scientific integrity and intended meaning of our work were preserved. Notably, the LLM was not used for core research activities, including the ideation of the AIMCoT framework, the design of experiments, the analysis of results, or the generation of the primary scientific claims. The final intellectual content and all contributions presented in this paper are entirely our own.
Appendix B Explanation for Figure 1
Figure 1 shows the images of the 22nd question on LLaVA-W benchmark. The query for this question is: "What’s the name of the restaurant serving these dishes?" and the image is a close-up photo of a meal at ICHIRAN. The left figure is the original image, and the right figure visualizes the regions selected by the Top-K strategy. The num_selected_patchesnum\_selected\_patches regions with the highest, second-highest, and third-highest scores are designated as the first, second, and third sets, colored red, purple, and blue, respectively. The first set is utilized by the baseline model according to its default design (num_selected_patches=72num\_selected\_patches=72).
Appendix C Introduction to the Benchmarks
M3CoT Chen et al. (2024) is a novel multimodal CoT benchmark, which introduces complex, multi-step problems across science, mathematics, and commonsense domains, comprising 11,459 samples in total. M3CoT is characterized by succinct textual queries (<15 tokens on average) paired with intricate problems. This inherent text-vision imbalance makes it an ideal platform to validate the efficacy of our proposed CAG in mitigating this issue and the superiority of AVP in proactively selecting the salient visual regions.
ScienceQA Saikh et al. (2022) is a popular benchmark for multiple-choice question answering with explanations on scholarly articles, comprising over 100,000 context-question-answer triples to address data scarcity in scientific machine reading comprehension.
LLaVA-Bench In-the-Wild (LLaVA-W) Liu et al. (2024) is a challenging open-ended benchmark designed to evaluate the real-world capabilities of VLMs by mimicking the unpredictability of real-world scenarios. The answers generated by GPT-4v Achiam et al. (2023) serve as the labels. LLaVA-W is exceptionally well-suited for evaluating the capability of our proposed framework to address complex, open-ended problems by generating a multimodal CoT, attending to salient regions within the image, and meticulously parsing the query.
Appendix D Introduction to the Baseline Models
No-CoT prompts the VLM to answer questions directly based on the input query and image. In the 1-shot setting, an example containing the query, image, and corresponding answer is attached.
DDCoT Zheng et al. (2023) deconstructs a multimodal problem into reasoning and recognition sub-questions, uses negative-space prompting to identify and fill visual information gaps with external models, and then integrates all information for a final joint reasoning step to generate rationales.
MMCoT Zhang et al. (2023) first generates a rationale from fused language and vision inputs, and then uses this rationale along with the original multimodal data to infer the final answer.
CCoT Mitra et al. (2024) first prompts the VLM to generate a scene graph from an image and then uses it as an intermediate reasoning step to produce the final response.
SCAFFOLD Lei et al. (2024) promotes vision-language coordination in the VLM by overlaying a dot matrix with coordinates onto an image, which then serves as a visual anchor that can be explicitly referenced in the textual prompt.
ICoT Gao et al. (2025) leverages the attention maps of the VLM to select relevant patches from the input image and insert them into the reasoning process, thereby generating sequential steps of paired visual and textual rationales.
Appendix E Greedy Algorithm within AVP Module
The complete process of the greedy algorithm within AVP is shown in Algorithm 1.
Input: total candidate set $C$, size of optimal selection $K$
Output: optimal selection $S$
1 $R^{*}\leftarrow\varnothing$
2 for $k\leftarrow 1,2,\cdots,K$ do
3 $U_{B}\leftarrow H(Y|I,x,y_{<t},R^{*})$
4 for $i\leftarrow 1,2,\cdots,N+M$ do
5    $U_{C,i}\leftarrow H(Y|I,x,y_{<t},R^{*}\cup\{R_{i}\})$
6    $IG(\{R_{i}\})\leftarrow U_{B}-U_{C,i}$
7 $R_{next}\leftarrow\text{argmax}_{R_{i}\in C\backslash R^{*}}\{IG(\{R_{i}\})\}$
8 $R^{*}\leftarrow R^{*}\cup\{R_{next}\}$
9
[…]
11 return $S$
Algorithm 1 Greedy Algorithm for Optimal Region Selection
Appendix F Detailed Analysis of Key Moments to Insert Visual Information
Experimental Setup. We take ICoT Gao et al. (2025) as a baseline model, which is required to answer all questions from the LLaVA-W benchmark in a 0-shot setting, with ROUGE-L used as the evaluation metric. The hyper-parameters follow the default settings of the open-source implementation for ICoT, and all experiments are conducted with the Chameleon-7B backbone.
Formal Definition of Attention Shifts. To analyze attention shifts, we examine the averaged attention maps across all attention heads in the last three layers of the VLM during the prediction of each token $t$. The model’s total attention scores allocated to the visual and text components of the input are respectively measured as follows:
$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},\hskip 14.22636ptA_{text}(t)=\sum_{j\in\text{indices of $C_{text}$}}\bar{a}_{t,j},$ (12)
where $C_{visual},C_{text}$ are the visual and text information within the context, respectively. Then, the shift in attention from the textual to the visual modality while generating token $t$ is defined as follows:
$\delta_{t}=A_{visual}(t)-A_{visual}(t-1).$ (13)
$\Delta_{k}=[\delta_{1},\delta_{2},\cdots,\delta_{|\Delta_{k}|}]$ encompasses the model’s attention shifts for each token when answering the arbitrary $k$-th question, where $|\Delta_{k}|$ is the number of tokens for answering the $k$-th question.
Formal Definition of Scores. for the predictions generated by the baseline model, the ROUGE-L scores are given by $List\_R=[R_{1},R_{2},\cdots,R_{|List\_R|}]$, where $R_{k}$ is the score for the model’s response to the $k$-th question, and |List_R||List\_R| is the number of questions within the benchmark.
Based on these concepts, we design a two-part experiment:
Experiment 1: Correlation Analysis. We investigate the relationship between the proportion of visual insertions under significant attention shifts and the score of the corresponding generated prediction.
First, to identify whether a visual insertion is conducted during a significant attention shift, we define a high attention growth threshold, $\delta_{k}^{(h)}$ for the $k$-th response ($\delta_{k}^{(h)}$ is set to the 80% upper quantile of $\Delta_{k}$ by default). An insertion is considered to have been conducted under a significant shift and referred to as a synchronized insertion if and only if its corresponding attention shift value exceeds the threshold $\delta_{k}^{(h)}$.
Next, since the model can conduct multiple insertions per response for a question, we calculate $P_{k}$, the proportion of synchronized insertions out of the total number of insertions for the $k$-th question.
Finally, since the proportions of synchronized insertions $P_{1},P_{2},\cdots,P_{|List\_R|}]$ and the ROUGE-L scores for all the questions $R_{1},R_{2},\cdots,R_{|List\_R|}]$ are obtained, the Pearson Correlation coefficient can be computed. Specifically, the Pearson Correlation is 0.2166 with a p-value of 0.048, which suggests that the proportions of the synchronized insertions and the corresponding score are significantly positively related to each other.
Experiment 2: Group Analysis. We investigate the relationship between the proportion of synchronized insertions and the quality of the model’s response.
To group the generated predictions according to response quality, we establish high- and low-scoring groups. All predictions are ranked in descending order by their ROUGE-L scores. The top 30% form $G_{h}$, the high-scoring group (high-quality responses), and $G_{l}$, the bottom 30% form the low-scoring group (low-quality responses).
Then, we calculate the mean proportion of synchronized insertions for groups $G_{h},G_{l}$, which are denoted as $\bar{P}_{h},\bar{P}_{l}$, respectively.
Finally, the means of the two groups $\bar{P}_{h},\bar{P}_{l}$ are compared, and a T-test is performed to assess the statistical significance of the difference. Specifically, we find that $\bar{P}_{h}=0.8889,\bar{P}_{l}=0.5000$, which suggests that in the high-scoring group, approximately 89% of insertions are the synchronized insertions with significant attention shift from textual input to visual information; in contrast, in the low-scoring group, only about half of the insertions are synchronized insertions. Besides, the P-value of T-test is as low as $0.0019$, which demonstrates that the result is highly statistically significant.
Appendix G Model Settings
G.1 Template of $\mathcal{P}_{CAG}$
Figure 4 provides an intuitive example showing the template of $\mathcal{P}_{CAG}$ and how it is used to prompt the VLM to carefully generate a guiding description for the input image. In particular, for multiple-choice questions, such as those in M3CoT and ScienceQA, we prepend the following brief explanation to $\mathcal{P}_{CAG}$ to aid the VLM in better understanding its designated task: "This is a multiple-choice question. The question is based on the image provided."
Figure 4: An illustration of the entire process of context enhancement by the CAG module, using problem physical-commonsense-1398 from the M3CoT benchmark as an example. This showcases both the template and usage of $\mathcal{P}_{CAG}$.
Then, the cross-attention weight matrix based on the enhanced context $x^{\prime}$ and $I$ can be obtained as follows:
$A^{\prime}=\text{softmax}(\frac{(H_{T}W^{Q})(H_{V}W^{K})^{T}}{\sqrt{d_{K}}}),$ (14)
where $H_{T}\in\mathds{R}^{n_{T}\times d},H_{V}\in\mathds{R}^{n_{V}\times d}$ are the hidden states of the textual and visual input, respectively. $W^{Q},W^{K}$ are the weight matrices of the linear transformation layers for query and key, respectively.
G.2 Setting of Hyper-parameters
The hyper-parameter settings are presented in Table 5. For easier understanding, we provide further explanation of the hyper-parameters $s_{r}$ and $s_{g}$ as follows: in the AVP module, the input image is first divided into $s_{g}\times s_{g}$ grids according to the set "Grid size" $s_{g}$. Then, each region finally selected by AVP consists of $s_{r}\times s_{r}$ grids.
Table 5: Hyper-parameter settings across three datasets.
Parameter M3CoT ScienceQA LLaVA-W
$N_{C}$ 8 8 6
$K$ 3 3 3
NN 4 4 2
$M$ 4 4 1
Region size for AVP $s_{r}$ (grid) 1 1 1
Grid size for AVP $s_{g}$ 4 4 4
$\delta$ 0.5 0.2 0.2
Temperature 0.7 0.7 0.7
Do sample True True True
Top_p 0.9 0.9 0.9
Repetition_penalty 1.2 1.2 1.2
Min_new_tokens 32 32 32
Max_new_tokens 512 1024 1024
Appendix H Empirical Analysis of the Approximate Submodularity of Function $F$
To offer a more comprehensive insight into the motivation for employing a greedy algorithm, this section provides a thorough analysis. We emphasize that for functions that are not theoretically submodular, a greedy approach remains one of the conventional methods for addressing their maximization, as established in recognized works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011). In this part, we conduct meticulously designed experiments to investigate the extent to which the information gain function $F$ approximates submodularity. The experimental results reveal that $F$ empirically exhibits significant submodular characteristics. This finding motivates us to follow established works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011) to propose a greedy algorithm to solve the maximization problem for $F$. The analysis is detailed as follows:
Firstly, we would like to introduce the definition of a submodular function. According to existing research Nemhauser et al. (1978), a function $f$ is a submodular function if it satisfies
$f(A\cup\{R_{i}\})-f(A)\geq f(B\cup\{R_{i}\})-f(B)$ (15)
for any sets $A\subseteq B\subset C$ and any element that satisfies $R_{i}\in C\backslash B$. In our scenario, the Inequality 15 is written as
$F(A\cup\{R_{i}\})-F(A)\geq F(B\cup\{R_{i}\})-F(B)$ (16)
for any $A\subset B\subset C$ and any $R_{i}\in C\backslash B$, which means that the information gain from incorporating a visual region exhibits a diminishing returns property.
To demonstrate this empirically, we design the experiment detailed as follows, aiming to show that for two sets of regions of different sizes, $S_{small}\subset S_{large}\subset C$, the information gain from incorporating a given visual region $R_{test}\in C\setminus S_{large}$ into the context of a VLM is greater when $R_{test}$ is added to $S_{small}$ than when it is added to $S_{large}$, ceteris paribus.
Experimental Setup. In our experimental design, each time the AVP process is triggered to select salient regions, we first execute it to select $K_{small}$ regions from the total candidate pool $C$ to form the set $S_{small}$. Subsequently, building upon $S_{small}$, we select an additional $K_{large}-K_{small}$ regions to construct the set $S_{large}$, where $K_{small}$ and $K_{large}$ are the respective set sizes. This construction inherently ensures that $S_{small}\subset S_{large}$.
Next, to compute the information gain contributed by a given region, we randomly sample a region $R_{test}$ from $C\setminus S_{large}$. We then calculate the VLM’s information content, which are denoted as $U_{s},U_{s}^{*},U_{l},\text{and }U_{l}^{*}$, when the context incorporates (1) $S_{small}$, (2) $S_{small}\cup\{R_{test}\}$, (3) $S_{large}$, and (4) $S_{large}\cup\{R_{test}\}$, respectively. We expect to observe in the majority of cases that:
$U_{s}^{*}-U_{s}\geq U_{l}^{*}-U_{l}.$ (17)
We conduct experiments on the M3CoT and LLaVA-W benchmarks, setting $K_{small}\in\{2,3,4,5\}$ and $K_{large}=K_{small}+1$ for simplicity. In terms of evaluation, we record the proportion of instances for which the inequality $U_{s}^{*}-U_{s}\geq U_{l}^{*}-U_{l}$ holds, and further introduce a Binomial Test to rigorously examine the significance of the results.
Experimental Results. The experimental results are presented in Table 6. As we can see, the Inequality 15 holds in most instances across all settings and datasets. Furthermore, to confirm the significance of the obtained results, we introduce the Binomial Test, an exact statistical procedure for assessing the extent to which experimental outcomes with a binary structure are attributable to chance alone. The p-values, presented in Table 6, are all substantially below the 0.05 significance level. This demonstrates that the information gain function $F$ behaves in a manner that is empirically near-submodular, which motivates us to follow existing research Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011) where greedy algorithms are proposed to solve the problem of maximizing approximately submodular functions.
Table 6: Proportions of instances on M3CoT and LLaVA-W benchmarks for which the approximate submodularity of information gain function $F$ is manifested. The backbone is Chameleon-7B and the model is our proposed AIMCoT. $K_{large}$ is set to $K_{small}+1$ for simplicity. The significance levels of these results are listed below them.
$K_{small}$ 2 3 4 5
M3CoT (n=2318) 72.00% 62.99% 67.04% 61.09%
P-value <1e-6 <1e-6 <1e-6 <1e-6
LLaVA-W (n=60) 61.67% 68.33% 61.67% 63.33%
P-value 0.0462 0.0031 0.0462 0.0249
Appendix I Analysis of the Selected Regions’ Source
In this section, we examine the distribution of sources for the visual regions selected by the AVP module of AIMCoT. These regions are drawn from two sets, $C_{attn}$ and $C_{exp}$, with their respective selection proportions denoted as $P_{attn}$ and $P_{exp}$. Intuitively, $P_{exp}$ reflects the significance of incorporating the exploratory set $C_{exp}$ to construct a better multimodal CoT. A larger value of $P_{exp}$ indicates that the exploratory set $C_{exp}$ makes a greater contribution by providing informative salient regions to AIMCoT, and vice versa.
Experimental Setup The experiments are conducted on the M3CoT and LLaVA-W benchmarks. Our proposed AIMCoT is implemented with the Chameleon-7B backbone under a default 0-shot setting. To ensure the reliability of the results, we repeat each experiment three times on both benchmarks.
Results and Analysis As presented in Table 7, although the value of $P_{exp}$ fluctuates across different experimental runs on the same benchmark, it remains consistently around 20% on M3CoT and 30% on LLaVA-W. This indicates that the influence of stochastic factors on the source distribution of the selected regions is limited, which validates our rationale of using this metric as a reflection of the relative importance of $C_{attn}$ and $C_{exp}$. Furthermore, we observe that $P_{exp}$ is significantly greater than zero. This demonstrates that the exploratory set $C_{exp}$ consistently serves as a critical component of the total candidate set $C$, contributing a substantial portion of the informative regions for AIMCoT.
Table 7: Proportion of salient regions selected by the AVP module of our proposed AIMCoT from the exploratory set $C_{exp}$.
Experiment Number 1 2 3
M3CoT 17.25% 20.44% 27.27%
LLaVA-W 31.33% 25.77% 26.67%
Appendix J Deployment of AIMCoT
J.1 Analysis of the Complexity of AVP Module
Overview of the AVP Module
The Active Visual Probing (AVP) module’s primary purpose is to dynamically and intelligently select salient sub-regions of an image during the text generation process. This is achieved by calculating the “information gain” that each potential sub-region offers, thereby allowing the model to “zoom in” on relevant visual details and generate more informed and contextually aware text.
The AVP logic is primarily encapsulated in three key methods:
  1. 1.
    forward: The main entry point where the AVP process is triggered based on changes in visual attention.
  2. 2.
    _generate_candidate_regions: Generates a diverse set of potential image regions (candidates) for evaluation.
  3. 3.
    _calculate_information_gain_iterative: The core of the AVP module. It iteratively evaluates candidate regions and selects the combination that maximizes the reduction in uncertainty (entropy) for the next token prediction.
Definition of the Notations
Let’s define the key variables that will be used in the complexity analysis:
  • NN: The current sequence length of the input tokens.
  • $N_{C}$: The total number of candidate regions generated (avp_num_candidates).
  • $K$: The number of regions to be selected in each AVP cycle (avp_num_regions_to_select).
  • $G$: The grid size of the vision model’s feature map (e.g., model_vision_grid_size, which is 4 by default, making the total number of patches $G^{2}=16$).
  • $V_{sub}$: The number of visual tokens (“vokens”) generated for a single cropped sub-image region.
  • $\Delta N$: The length added per selected region, where $\Delta N=V_{sub}+2$ (accounting for the boi and eoi tokens).
  • $L$: The number of layers in the transformer model.
  • $H$: The hidden size of the model.
  • $V_{vocab}$: The size of the model’s vocabulary.
AVP Triggering in the forward Method
The AVP mechanism is not activated on every forward pass; instead, it is triggered conditionally based on the change in attention directed towards the visual tokens. Specific to its operational process, in terms of attention calculation, the code calculates latest_vattns—which refers to the sum of attention scores from the last token to all visual patch tokens—and this step requires iterating through the attention matrices. Meanwhile, regarding the trigger condition, the core logic is if delta_vattns > config[’delta’], where delta_vattns represents the difference between the current and previous visual attention sums. In conclusion, the cost of this trigger check per token generation is minimal; it primarily involves retrieving and summing pre-computed attention scores. The complexity is approximately $O(L\cdot N)$ to extract and sum the relevant attention weights to the $G^{2}$ visual patches, but this is dwarfed by the main model’s complexity.
_generate_candidate_regions Method
This method generates $N_{C}$ candidate regions from the image’s attention map, using a hybrid strategy that combines attention-based and random sampling. Specifically, for attention-based candidates, it first flattens the $G\times G$ attention map, then uses torch.topk to find the indices of the avp_num_attention_based patches with the highest attention—with the complexity of topk on a tensor of size $G^{2}$ being $O(G^{2}\log(\texttt{avp\_num\_attention\_based}))$—and subsequently creates bounding boxes around these top patches, which is a constant time operation for each of the avp_num_attention_based candidates; for random candidates, it generates the remaining $N_{C}-\texttt{avp\_num\_attention\_based}$ candidates by randomly selecting coordinates, an operation with complexity $O(N_{C}-\texttt{avp\_num\_attention\_based})$. As a result, the time complexity in this part is $O(G^{2}\log(\texttt{avp\_num\_attention\_based})+N_{C})$, and since avp_num_attention_based is a small constant and $G^{2}$ is fixed (e.g., 16), this can be considered approximately $O(N_{C})$.
_calculate_information_gain_iterative Method
This is the most computationally intensive part of the AVP module. It employs a greedy, iterative approach to select the $K$ best regions out of $N_{C}$ candidates. The method consists of an outer loop that runs $K$ times (for each region to be selected). Inside this loop, it evaluates the remaining candidates to pick the one that provides the highest immediate information gain.
Let’s analyze a single iteration of this outer loop (e.g., the $k$-th iteration, where $k$ ranges from 0 to K1K-1):
First, for the initial entropy calculation, it performs one forward pass through the base model (self.model) with the current sequence of tokens (which includes tokens from $k$ previously selected regions), where the sequence length at this stage is $N_{k}=N+k\cdot\Delta N$; with Key-Value (KV) caching from previous iterations, the cost can be incremental: $O(L\cdot\Delta N\cdot N_{k-1}\cdot H)$ for updating the cache with the last selected region’s tokens, rather than a full $O(L\cdot N_{k}^{2}\cdot H)$.
Next, for the batch preparation for lookahead analysis, the code iterates through the remaining $N_{C}-k$ candidate regions, and for each candidate, it performs cropping and vokenization (cropping the image pixels and passing them to self.model.get_image_tokens, which involves a forward pass through the vision encoder with complexity approximately $O(G^{2})$ per crop, negligible compared to the transformer) and tensor concatenation (creating a new input sequence by appending the new vokens, with the length of this new sequence being $N_{k}+\Delta N$), with this loop running $N_{C}-k$ times.
Subsequently, for the batch forward pass (lookahead), the $N_{C}-k$ new input sequences are padded and batched together, a single batched forward pass is performed on these $N_{C}-k$ sequences, the maximum sequence length in this batch is $N_{k}+\Delta N$, and since all lookahead sequences share the same prefix of length $N_{k}$, KV caching for the prefix can be reused across the batch; the complexity is then self-attention among suffix tokens: $O(L\cdot(N_{C}-k)\cdot\Delta N^{2}\cdot H)$ and cross-attention to prefix: $O(L\cdot(N_{C}-k)\cdot\Delta N\cdot N_{k}\cdot H)$, with the dominant term (when $N_{k}\gg\Delta N$) being $O(L\cdot(N_{C}-k)\cdot\Delta N\cdot N_{k}\cdot H)$, which is the key optimization from batch processing and caching, reducing from quadratic to linear dependence on $N_{k}$.
Additionally, for the information gain calculation, it calculates the entropy for each of the $N_{C}-k$ outputs from the lookahead pass, which involves a softmax over the vocabulary and is $O((N_{C}-k)\cdot V_{vocab})$.
Finally, for the selection and state update, torch.argmax finds the best candidate in $O(N_{C}-k)$ time, and the base input sequence is updated for the next iteration.
Overall Complexity of the Method
We must sum the complexity over the K iterations of the outer loop. The most computationally intensive operation is the batched lookahead forward pass, which is significantly optimized through batch processing and Key-Value caching. The overall time complexity can be expressed as follows:
$\sum_{k=0}^{K-1}O\left(L\cdot(N_{C}-k)\cdot\Delta N\cdot(N+k\cdot\Delta N)\cdot H\right).$ (18)
This formula highlights that, thanks to KV caching, the complexity scales linearly with sequence length N—a substantial improvement over the standard quadratic dependence. While the cost is also linear with respect to the number of candidates $N_{C}$ and selections $K$, our framework maintains strong deployability. This is because the AVP module is highly efficient at extracting crucial visual information from the candidate set. Our empirical results demonstrate that AIMCoT achieves exceptional performance (as shown in Table 2) even when these hyper-parameters are kept at low levels (e.g., $K=3,N_{C}=6$, which is the default setting). In conclusion, we can approximate the complexity of AVP as follows:
$O\left(K\cdot N_{C}\cdot L\cdot\Delta N\cdot(N+K\cdot\Delta N)\cdot H\right).$ (19)
This demonstrates that the combination of architectural optimizations (batching and KV caching) and the high extractive efficiency of AVP ensures the module’s practicality, making it efficient and readily deployable in practice. As a comparison, the attention-driven selection method within the baseline model ICoT is with a complexity of $O(logK\cdot\Delta N+L\cdot N\cdot H)$, which also scales linearly with sequence length N. Although this is acknowledged to be lower than the complexity of AVP shown in Equation 19, our empirical results detailed in Appendix J.2 suggest that AVP’s average inference time is no more than 1.36 times that of this method, while achieving performance far superior to it as analyzed in Section 5.5.
J.2 Empirical Analysis of the Deployment of AIMCoT
In this section, we empirically investigate the deployability of the AIMCoT framework and the temporal overhead introduced by the AVP module. For the experimental setup, we utilized Chameleon-7B as the backbone in a 0-shot setting to compare the average inference time of AIMCoT against ICoT Gao et al. (2025). ICoT, as a key baseline model in this study, employs a Top-K strategy to simultaneously select regions with the highest attention scores for constructing the multimodal CoT, thereby expected to possess a relatively lower time complexity compared to AIMCoT. Consequently, the comparison with ICoT serves as a direct indicator of AIMCoT’s deployability.
Table 8: Comparison of the average time to process each instance between AIMCoT and the baseline model (ICoT).
Dataset AIMCoT ICoT
M3CoT 13.37s 11.62s
LLaVA-W 11.65s 8.58s
The experimental results are presented in Table 8. Two key observations can be drawn. First, the AVP module does not introduce significant temporal costs to the AIMCoT framework, an efficiency attributable to batch processing and the KV Cache mechanism. Second, AIMCoT achieves substantially superior performance as shown in Table 2, at a time cost comparable to that of the efficient baseline, which is less than $1.36$ times that of ICoT (specifically, $1.15$ and $1.36$ times on M3CoT and LLaVA-W benchmarks, respectively). This suggests that our proposed AIMCoT framework achieves a favorable trade-off between performance and deployability.
Appendix K Sensitivity Analysis of $\delta$
The hyper-parameter $\delta$ within the DAT module serves as a crucial threshold to trigger the AVP module, which inserts salient visual regions to improve the construction of the multimodal CoT. In this section, we detail a sensitivity analysis of $\delta$ by adjusting $\delta$ across the range of $[0.1,0.125,0.15,0.175,0.2,0.225]$ and examining not only (1) the performance of AIMCoT, but also (2) the number of times the AVP is triggered. The experiments are conducted under 0-shot setting on Chameleon-7B backbone and LLaVA-W benchmark. The experimental results are shown in Figure 5.
The left figure illustrates that AIMCoT exhibits limited performance when the threshold, $\delta$, is set too low. This underscores the importance of inserting visual information at critical moments: excessively frequent or inopportune visual insertions can disrupt the VLM’s reasoning process, leading to suboptimal performance. As $\delta$ increases, the model’s performance progressively improves, reaching its peak at $\delta=0.2$ (our default setting), which corresponds to a ROUGE-L score of 0.2983. However, a further increase in $\delta$ results in a slight performance degradation. This highlights the criticality of visual information insertion for constructing an interleaved Chain of Thought: an overly stringent threshold excessively impedes the incorporation of visual data, preventing the AVP from supplying the model with necessary visual supplementation in a timely manner.
Conversely, the right figure demonstrates a consistent decrease in the number of times the AVP is triggered as $\delta$ is raised. This showcases the efficacy of $\delta$ as a threshold for modulating the activation frequency of the AVP.
Figure 5: Experimental results of the sensitivity analysis of the hyper-parameter $\delta$. The left figure illustrates the performance of AIMCoT when $\delta$ takes different values, while the right one shows the number of times the AVP module within AIMCoT is triggered.
Table 9: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on M3CoT benchmark, which illustrate the average time of AIMCoT to process an instance under different settings. Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
$N_{C}$ 1 2 3 4 5 6 7
K=1 12.78s 12.80s 12.83s 12.87s 12.90s 12.93s 12.97s
K=3 - - 13.18s 13.23s 13.30s 13.37s 13.42s
K=5 - - - - 13.41s 13.49s 13.56s
K=7 - - - - - - 13.66s
Table 10: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on LLaVA-W benchmark, which illustrate the average time of AIMCoT to process an instance under different settings. Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
$N_{C}$ 1 2 3 4 5 6 7
K=1 10.61s 10.70s 10.75s 10.84s 10.92s 11.00s 11.08s
K=3 - - 11.21s 11.36s 11.51s 11.65s 11.80s
K=5 - - - - 11.79s 11.96s 12.14s
K=7 - - - - - - 12.38s
Appendix L Sensitivity Analysis of $K,N_{C}$
Our proposed AIMCoT incorporates the design of AVP module. In contrast to existing research, AIMCoT, benefiting from the AVP module, does not simply select the top-$K$ regions with the highest attention scores from the attention map. Instead, it meticulously selects $K$ regions from a total set $C$ of $N_{C}$ candidate regions to construct the multimodal CoT ($N_{C}=N+M$). However, this approach may inevitably raise concerns regarding the deployability of AIMCoT, particularly as the hyper-parameters $N_{C}$ and $K$ increase.
To investigate this, we conduct experiments to explore the average processing time per instance for AIMCoT with larger values of $N_{C}$ and $K$. For the experimental setup, we implement AIMCoT with the Chameleon-7B backbone on the M3CoT and LLaVA-W datasets under various combinations of $N_{C}$ and $K$. The results are presented in Tables 9 and 10, from which we derive two key insights:
  • Insensitivity to the growth of the candidate set size $N_{C}$: observing any row with a fixed $K$ (e.g., $K=3$ on Table 9), as $N_{C}$ increases from 3 to 7, the total processing time rises from 13.18s to 13.42s, a marginal increase of only 0.24s. This implies that each additional candidate region introduces an average overhead of less than 0.06s. This strongly demonstrates that the performance of the AVP module does not degrade sharply with a moderate expansion of the candidate pool, indicating excellent scalability.
  • Diminishing marginal cost with the increase in the number of selections $K$: considering a fixed column for $N_{C}$ (e.g., $N_{C}=7$ on Table 10), as $K$ increases from $1\rightarrow 3$, $3\rightarrow 5$, and $5\rightarrow 7$, the processing time increases by 0.72s, 0.34s, and 0.24s, respectively, which shows a notable decrease in incremental cost. This suggests that the primary computational overhead of the AVP algorithm lies in the initiation of the iterative search (the jump from $K=1$ to $K=3$). Once the iteration begins, the cost of subsequent selection steps is remarkably low, benefiting from efficient mechanisms such as the batch processing and KV Cache.
Based on these key insights, our proposed AVP module demonstrates high computational efficiency and robustness when faced with increased computational complexity (i.e., larger $N_{C}$ and $K$ values).
Table 11: Performance comparison AIMCoT variants on the basis of different constructions of the candidate set $C$.
Construction of $C$ M3CoT (ACC.) LLaVA-W (ROUGE-L)
$C_{attn}\cup C_{rand}~~~(C_{exp}=C_{rand})$ 31.4 29.8
$C_{attn}\cup C_{ss}~~~(C_{exp}=C_{ss})$ 31.2 29.5
$C_{attn}\cup C_{fsam}~~~(C_{exp}=C_{fsam})$ 31.0 29.6
$C_{attn}$ 30.8 28.9
$C_{rand}$ 30.4 28.6
$C_{ss}$ 30.3 28.7
$C_{fsam}$ 29.9 27.7
Appendix M Ablation Study on the Construction of Candidate Set $C$
In this section, we investigate the influence of different compositions of the total set $C$ on the performance of our proposed AIMCoT. We specifically examine two primary configurations:
Constructing $C$ using only $C_{attn}$ or $C_{exp}$. For the latter, we evaluate four distinct construction methodologies for $C_{exp}$: (a) $C_{rand}$: uniform random sampling; (b) $C_{ss}$: the selective search algorithm Uijlings et al. (2013), which is the seminal region proposal method utilized in R-CNN Girshick et al. (2014); (c) $C_{fsam}$: FastSAM Zhao et al. (2023), a computationally efficient variant of the foundational vision segmentation model, SAM Kirillov et al. (2023).
Constructing $C$ using both $C_{attn}$ and $C_{exp}$. Similarly, as for $C_{exp}$, we also consider its diversified construction, including $C_{rand},C_{ss}$, and $C_{fsam}$.
It is worth noting that although the construction is diverse, the size of $C$ remains consistent. When $C$ is composed of $C_{attn}$ and $C_{exp}$, the two each account for half. The experimental results are shown in Table 11.
As observed, the combination of the two sets (i.e., $C=C_{attn}\cup C_{exp}$) invariably yields superior performance for AIMCoT compared to configurations where either $C=C_{attn}$ or $C=C_{exp}$ is used exclusively. This highlights the importance of diversifying the sources of candidate visual regions.
When comparing the different construction methods for $C_{exp}$, the performance gap among models is marginal when used in conjunction with $C_{attn}$. Specifically, despite its simplicity, random sampling achieves highly competitive results, which motivates our choice to adopt it as the default method for constructing $C_{exp}$. Intuitively, the advantage of random sampling lies in its ability to provide regions across different parts of the image unbiasedly with maximal spatial diversity.

### v2

@@
-AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning
+AIM-CoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning
 Abstract
-Multimodal Chain-of-Thought (CoT) has emerged as a powerful technique for enhancing the vision-language reasoning with interleaved information.
-However, existing methods often rely on simplistic heuristics for constructing interleaved CoT, typically depending on attention maps, which our empirical analysis reveals can be unreliable.
-What’s more, the shortcomings of their passive and purposeless selection strategies and their arbitrary triggering mechanisms in capturing the model’s cognitive need for information are further amplified.
-In this paper, we propose AIMCoT, an Active Information-driven Multi-modal Chain-of-Thought framework that addresses these fundamental limitations.
-AIMCoT introduces three synergistic components: (1) Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity imbalance, thereby producing more reliable attention maps as a foundation.
-(2) Active Visual Probing (AVP), which replaces passive selection with a proactive, goal-oriented strategy grounded in information theory to select image regions that help answer the questions maximally.
-(3) Dynamic Attention-shifting Trigger (DAT), which intelligently determines the optimal moments to insert visual information by monitoring the model’s text-to-vision attention shifts.
-Extensive experiments on three challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods across different settings.
-By actively foraging for information and dynamically structuring its reasoning process, AIMCoT represents a critical step towards more robust, effective, and human-like multimodal reasoning.
+Interleaved-Modal Chain-of-Thought (I-MCoT) advances vision-language reasoning, such as Visual Question Answering (VQA).
+This paradigm integrates specially selected visual evidence from the input image into the context of Vision-Language Models (VLMs), enabling them to ground their reasoning logic in these details.
+Accordingly, the efficacy of an I-MCoT framework relies on identifying what to see (evidence selection) and when to see it (triggering of insertions).
+However, existing methods fall short in both aspects.
+First, for selection, they rely on attention signals, which are unreliable—particularly under severe granularity imbalance between the brief textual query and the informative image.
+Second, for triggering, they adopt static triggers, which fail to capture the VLMs’ dynamic needs for visual evidence.
+To this end, we propose a novel I-MCoT framework, Active Information-driven Multi-modal Chain-of-Thought (AIM-CoT), which aims to improve both evidence selection and insertion triggering via: (1) Context-enhanced Attention-map Generation (CAG) to mitigate granularity imbalance via textual context enhancement; (2) Active Visual Probing (AVP) to proactively select the most informative evidence via an information foraging process; and (3) Dynamic Attention-shift Trigger (DAT) to precisely activate insertions when VLM’s attention shifts from text to visual context.
+Experiments across three benchmarks and four backbones demonstrate AIM-CoT’s consistent superiority.
 Our code is available at https://anonymous.4open.science/r/AIMCoT.
+AIM-CoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning
+Xiping Li1, Jianghong Ma2thanks: Corresponding Author 1The Chinese University of Hong Kong 2Harbin Institute of Technology (Shenzhen) lihsiping@gmail.com, majianghong@hit.edu.cn
 1 Introduction
-The advent of Chain-of-Thought (CoT) prompting has marked a significant milestone in the reasoning capabilities of Large Language Models (LLMs) Wei et al.
+Chain-of-Thought (CoT) prompting, initially established for Large Language Models (LLMs) Wei et al.
 (2022); Wang et al.
-(2022); Zhang et al.
+; Zhang et al.
 (2023); Suzgun et al.
 (2023); Li et al.
-(2025); Yao et al.
-(2023); Besta et al.
-(2024), enabling them to deconstruct complex problems into a series of intermediate, interpretable steps.
-This paradigm has been naturally extended to Vision-Language Models (VLMs), where early efforts Mitra et al.
+(2025); Wang et al.
+(2023); Diao et al.
+(2024); Chu et al.
+(2024), has been naturally adapted to Vision-Language Models (VLMs) Zhang et al.
+(2025); Chen et al.
+(2024b); Cheng et al.
+(2025); Xu et al.
+(2024), empowering them to tackle vision-related tasks via a series of intermediate reasoning steps.
+Visual Question Answering (VQA) serves as a representative scenario targeted in this study.
+In a typical VQA scenario, the VLM is presented with a multimodal question comprising a textual query and an associated image, and is then prompted to answer the question.
+Early efforts in this domain Mitra et al.
 (2024); Zheng et al.
@@
 (2025); Zhang et al.
-(2023) focused on generating text-only rationales to articulate the model’s reasoning process over visual inputs.
-A pivotal advancement in this domain was the introduction of Interleaved-modal Chain-of-Thought Gao et al.
-(2025), which pioneered the direct integration of visual patches into the reasoning chain.
-By pairing textual rationales with corresponding image regions, the interleaved CoT demonstrated a superior ability to ground language in visual evidence, setting a new standard for multimodal reasoning.
-However, the efficacy of existing research Ge et al.
-(2025); Gao et al.
-(2025) is fundamentally predicated on the reliability of their underlying mechanisms for selecting and integrating visual information.
-These methods typically rely on a passive, attention-driven strategy: they select the Top-K regions with the highest attention scores and insert them at predefined moments, e.g., the appearance of a newline character, a practice with little theoretical or empirical justification for its timing.
-This reliance exposes critical vulnerabilities that limit their full potential.
-Our empirical analysis reveals that high-attention regions are often redundant or, more alarmingly, fail to capture the most crucial visual details, especially when a granularity mismatch exists between the textual query and the visual evidence.
-This raises three fundamental questions: (1) What information source can reliably identify truly salient visual regions?
-(2) How can we accurately select useful regions in a proactive and purposeful manner rather than passively relying on potentially unreliable attention scores?
-(3) When is the optimal moment to insert visual evidence into the reasoning process?
-In this work, we postulate that a more robust multimodal reasoning framework requires a shift from passive attention-following to an active, information-seeking paradigm.
-We propose AIMCoT: Active Information-driven Multi-modal Chain-of-Thought, a novel framework that directly addresses the limitations of prior work, reframing the selection of visual evidence from a passive, attention-based retrieval task to an active, goal-oriented probing process.
-Inspired by the principles of information foraging Pirolli & Card (1999); Broadbent (2013), AIMCoT operates on the premise that the most valuable visual rationale is one that maximally reduces the model’s uncertainty about the subsequent step in its reasoning chain.
-Rather than simply asking "Where is the model looking?", we prompt the model to actively ask "Which piece of visual information will be most helpful for me to see right now?".
-This shift is realized through three synergistic components:
-  1. 1.
-    Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity disparity to improve the reliability of attention map for better identifying salient regions by generating a context-aware description of the image.
-  2. 2.
-    Active Visual Probing (AVP), which implements a proactive, goal-oriented selection strategy grounded in information theory, selecting a set of visual regions that provide the highest possible information gain for the task at hand.
-  3. 3.
-    Dynamic Attention-shifting Trigger (DAT), an intelligent triggering mechanism that carefully captures the critical moments when model’s cognitive focus shifts significantly from text to vision, and inserts visual information precisely.
-Our contributions are as follows:
-  • We introduce AIMCoT, a novel training-free framework that reframes the construction of multimodal CoT as an active information-foraging process, moving beyond the limitations of static, passive, and purposeless region selection.
-  • We propose a system comprised of three complementary methods (CAG, AVP, DAT) that collectively enable VLMs to proactively forage for informative visual evidence and strategically integrate it into their reasoning process at the detected critical moments.
-  • We present a comprehensive set of empirically-grounded motivations that inspire the designs within our AIMCoT framework.
-Furthermore, we provide substantial theoretical analysis and design targeted experiments to explore the properties of these designs, covering key aspects such as the deployability of AIMCoT, the interplay between the CAG and AVP modules, and the necessity of incorporating an exploratory candidate pool.
-  • Through extensive experiments conducted on two backbones (Chameleon-7B and Qwen2-VL-7B) across challenging benchmarks including M3CoT, ScienceQA, and LLaVA-W, we demonstrate that AIMCoT significantly outperforms state-of-the-art baselines and advances the frontier of vision-language reasoning.
+(2024) focused on generating text-only CoT.
+A pivotal advancement is the paradigm of Interleaved-Modal Chain-of-Thought (I-MCoT) Gao et al.
+(2025).
+Unlike text-only methods, this paradigm aims to provide fine-grained visual evidence for model reasoning.
+Specifically, it first selects the salient sub-regions from the input image, and then inserts them as visual tokens into the context of reasoning chain.
+However, to construct an effective I-MCoT framework, two critical questions should be well addressed: (1) What to see?
+(Selection: identifying the specific image regions that support the current reasoning step) and (2) When to see it?
+(Triggering: determining the precise moment visual insertion is needed).
+Existing research Gao et al.
+(2025) falls short in these two aspects due to its reliance on passive and static heuristics.
+Regarding selection, it adopts attention-based selection, which relies on the quality of cross-attention maps of VLM layers.
+Specifically, the visual regions that receive the most attention from tokens in the text context are selected, assumed to be most helpful for the subsequent generation.
+However, as revealed in Section 3.1, the raw VLM attention fails to accurately identify the salient regions, particularly when there is a granularity imbalance.
+Specifically, in the input context, the rich details in the image overwhelm the brief textual query.
+Therefore, with limited semantic anchors, the query cannot effectively steer the cross-attention towards the crucial visual regions.
+Regarding triggering, prior works often insert visual information at fixed, predefined moments (e.g., upon generating a newline character Gao et al.
+(2025)).
+As analyzed in Section 3.3, such a static mechanism fails to align with the model’s dynamic cognitive need for fine-grained visual information.
+To overcome these limitations, we propose a novel I-MCoT framework, Active Information-driven Multi-modal Chain-of-Thought (AIM-CoT), which aims to address both what to see and when to see it by shifting VLM reasoning from passive, static perception to active, dynamic exploration.
+Grounded in Information Foraging Theory Pirolli and Card (1999); Broadbent (2013), AIM-CoT detects the VLM’s momentary need for visual cues and proactively forages for the most informative evidence.
+Integrated into the context window, this evidence improves subsequent generation.
+AIM-CoT realizes this shift through three synergistic components: (1) Context-enhanced Attention-map Generation (CAG): To refine the VLM’s cross-modal attention distribution, CAG elicits a query-conditioned description of the image and appends it to the textual context.
+Rather than serving as an auxiliary caption, CAG provides semantic anchors that align textual and visual granularities.
+(2) Active Visual Probing (AVP): To address the “What to see” problem, AVP quantifies the information gain of candidate regions, integrating those that provide the highest utility into the context for subsequent reasoning and generation.
+(3) Dynamic Attention-shift Trigger (DAT): To address the “When to see it” problem, DAT dynamically and precisely triggers visual insertion when a significant shift from textual to visual focus is detected, indicating the VLM’s cognitive demand for visual evidence.
+Our contributions are summarized as follows:
+  • We introduce AIM-CoT, a unified system (CAG, AVP, DAT) that enables VLMs to proactively forage for informative visual evidence and dynamically capture the critical triggering moments.
+  • Our analyses demonstrate that AIM-CoT (1) selects truly salient visual evidence via information gain, (2) withstands noisy attention and descriptions, and (3) remains deployment-friendly.
+  • Experiments conducted on four backbones across three VQA benchmarks demonstrate the consistently superior performance of AIM-CoT over state-of-the-art baselines.
 2 Related Work
-The success of CoT prompting in LLMs has naturally extended to VLMs, aiming to make their reasoning processes explicit and interpretable.
-Early efforts focused on generating text-only rationales.
-MMCoT Zhang et al.
-(2023) first generates a rationale from the input and then uses this rationale along with the original multimodal data to infer the final answer.
-CCoT Mitra et al.
-(2024) prompts the VLM to generate an intermediate scene graph to structure its understanding.
-DDCoT Zheng et al.
-(2023) decomposes complex problems into simpler sub-questions and leverages external models to fill information gaps.
-Alternatively, SCAFFOLD Lei et al.
-(2024) overlays a coordinate grid on the image, enabling the VLM to reference spatial regions explicitly in its textual reasoning.
-A pivotal advancement is the introduction of interleaved-modal CoT, which integrates visual evidence directly into the reasoning chain.
+Multimodal CoT has been widely adopted in VLM research.
+Early text-only efforts enhance reasoning by decomposing questions Zheng et al.
+(2023), generating intermediate scene graphs Mitra et al.
+(2024), or overlaying coordinate grids for spatial referencing Lei et al.
+(2025).
+A pivotal advancement is the I-MCoT paradigm, which integrates visual evidence directly into reasoning chains to improve subsequent reasoning/generation.
 The leading approach, ICoT Gao et al.
-(2025), selects the top-K regions from the VLM’s attention map and inserts them at predefined moments.
-MRFD Ge et al.
-(2025) meticulously selects the salient visual regions to foster more reliable reasoning, alleviating VLM hallucination.
-However, our analysis reveals that such passive, attention-driven strategies are fundamentally limited by the unreliability of raw attention maps, especially during text-vision granularity mismatches, and their arbitrary insertion points fail to capture the model’s dynamic cognitive needs.
-Our proposed AIMCoT addresses these critical vulnerabilities by shifting the paradigm from passive selection to active information foraging, employing goal-oriented visual probing and dynamic, attention-shift-triggered integration.
+(2025), selects the visual patches receiving the highest attention scores from the text tokens in the context.
+Then, this visual evidence is integrated when the VLM outputs a newline token.
+However, our analysis in Section 3 highlights the limitations of existing research.
+Therefore, while we adopt the established I-MCoT paradigm (similar to ICoT Gao et al.
+(2025)), our contributions lie in shifting from the passive selection and static trigger to the proactive information-foraging selection and dynamic attention-shifting trigger.
 3 Motivation
-3.1 Identifying the Reliability of the Attention Map
-The utilization of the attention map in multimodal learning has been explored by recent research Gao et al.
-(2025); Ge et al.
-(2025); Xie et al.
-(2022); Wang et al.
-(2023); Liu et al.
-(2022), inspiring us to also take it as a source from which the salient visual regions are selected.
-However, considering their strong dependence on attention map, we question its reliability by posing two fundamental questions:
+We focus on two pivotal questions for I-MCoT: What and When to see.
+First, we expose the limitations of passive attention-based selection in Section 3.1.
+To bridge this, we explore an active paradigm shift, investigating the potential of information gain for selection (Section 3.2) and dynamic attention shifts for triggering (Section 3.3).
+3.1 Moving Beyond Attention Maps: The Reliability Gap in Passive Visual Selection
+Existing I-MCoT methods adopt attention-based visual selection. We examine the reliability of attention for visual selection by asking:
   1. (1)
-    Do all significant regions on the attention map help the VLMs answer questions correctly?
+    Sufficiency: Do highly attended patches always benefit question answering?
   2. (2)
-    Does the attention map comprehensively capture all visual regions that are instrumental to the VLM’s correct prediction?
-Figure 1: The images of the 22nd question on LLaVA-W benchmark, which is a close-up photo of a meal at ICHIRAN.
-The left and right figures are respectively the original image and the first three sets of regions selected by the Top-K strategy (red, purple, and blue, respectively).
-A detailed explanation is shown in Appendix B.
-To acquire a holistic comprehension of these two questions, our empirical analysis is conducted on the popular Visual Question Answering (VQA) benchmark, LLaVA-W Liu et al.
-(2024), employing ICoT Gao et al.
-(2025) as a baseline model, which selects the Top-K regions from the attention map to construct a text-vision interleaved CoT.
-Chameleon-7B Team (2024) serves as the backbone.
-Table 1: Performance degradation of the baseline model (ICoT, 0-shot) when the Top $K_{mask}$ regions on the attention map are masked.
-$K_{mask}$ 0 1 5 10 20
-Degradation 0% 0.26% 1.43% 3.93% 2.44%
-We quantitatively investigate the role of high-attention regions by masking the top-$K_{mask}$ regions identified by ICoT (0-shot).
-On LLaVA-W, this masking leads to a minor performance drop of just 3.93% (top 10) and 2.44% (top 20), as shown in Table 1.
-Notably, performance increases when the mask is expanded from the top 10 to the top 20 regions.
-These results strongly suggest that not all high-attention regions contribute significantly to the model’s prediction: some have negligible impact or even introduce detrimental signals.
-In response to the second question, we manually inspect the content of high-attention regions to understand their role in VLM’s predictions.
-In a challenging LLaVA-W example (Figure 1), the answer lies in a small area, i.e., the inner rim of a ramen bowl, among rich visual information.
-The top two most attended image patch sets entirely miss this crucial detail, with only a negligible portion appearing in the third set.
-This indicates a potential misalignment between high-attention regions and key visual information, particularly when the text-vision granularity disparity is significant.
-3.2 Foraging for the Most Information to Guide Region Selection
-The analysis in Section 3.1 empirically reveals that the efficacy of the attention-driven Top-K strategy is notably constrained, particularly when faced with significant text-vision granularity disparities.
-Furthermore, fundamentally, since attention maps merely reflect token correlations, a static approach without an explicit goal like Top-K, which relies solely on attention scores, is inherently suboptimal.
-This insight directly motivates us to explore a selection method that is more proactive and purposeful.
-Existing recognized research Pirolli & Card (1999); Broadbent (2013); Oaksford & Chater (1994); Friston (2010) supports that people, when possible, will maximize their rate of gaining valuable information, as it yields more useful information per unit cost.
-Inspired by this, we explore providing the VLM with image regions that yield the highest information gain, thereby maximally reducing the model’s uncertainty in answering a given question.
-Intuitively, the model’s uncertainty can be quantified by the entropy of its probability distribution over the vocabulary given the current context, while the information gain of an image region is measured based on model’s entropy when the region is included in the context.
-The definitions are meticulously detailed in Section 4.3.
-Figure 2: The visualization of regions selected by the information gain-guided strategy.
-To continue with the example in Section 3.1, Figure 2 illustrates the top three regions ranked by the information gain for VLM.
-Evidently, in contrast to the regions selected via Top-K shown in Figure 1, the information gain-based selection accurately guides the VLM to first focus on the inner rim of the bowl (red), where the critical information is contained.
-Although the VLM does not yield the final answer in this region, we note that this indicates a correct line of reasoning, as the ground truth is situated in a highly similar area (a nearby location also on the inner wall of the bowl).
-Subsequently, the region ranked third (blue) precisely encompasses a large portion of the restaurant’s name, which is just the answer to the question.
-This suggests that even for a challenging case where the text-vision granularity is highly disparate, information gain serves as a better foundation for region selection.
-3.3 Inserting Visual Information at Key Moments
-Although existing research Gao et al.
-(2025) has attempted to construct interleaved CoT, they often ignore a reasonable triggering mechanism to capture the critical moment of inserting visual information.
-For example, ICoT Gao et al.
-(2025) simply uses a newline character as a trigger signal.
-This motivates an in-depth investigation into reliable indicators for identifying these key moments for inserting visual content.
-To this end, we conduct a comparative case study on the LLaVA-W benchmark, with ROUGE-L as the evaluation metric.
-We collect all predictions generated by the baseline model (ICoT) and partition them into high- and low-scoring groups based on ROUGE-L scores.
-Given that the model’s attention map has served as an important foundation for existing works Gao et al.
-(2025); Ge et al.
-(2025); Xie et al.
-(2022); Wang et al.
-(2023); Liu et al.
-(2022), which is not only readily accessible but also effectively reflects the model’s aggregate attention to the textual and visual parts of the input, we are inspired to monitor the attention shifts between these two modalities throughout the prediction process.
-The experiments are detailed in Appendix F, from which two key observations are revealed: (1) Correlation analysis: inserting visual data precisely when the model’s attention pivots towards the visual modality is strongly correlated with higher scores and (2) Group analysis: this phenomenon further serves as a crucial characteristic that distinguishes high-scoring from low-scoring outputs.
-4 AIMCoT
-In this section, we begin by briefly reviewing the background of multimodal learning. Then, we detail the proposed methods motivated by the following key insights derived from Section 3:
-  1. (1)
-    First, the high-scoring attention regions are not always beneficial for question-answering, and crucial visual evidence can be missed particularly in cases of text-vision granularity mismatch.
-  2. (2)
-    Second, for a given set of candidate regions, selection based on information gain significantly outperforms the conventional attention-driven Top-K method.
-  3. (3)
-    Finally, capturing the critical moments to insert visual information improves the construction of multimodal CoT, and text-to-vision attention shifts serve as an important indicator.
-Accordingly, we propose AIMCoT, which encompasses three key methods: (1) Context-enhanced Attention-map Generation (CAG), which generates a fine-grained description for the image to alleviate text-vision disparity.
-(2) Active Visual Probing (AVP), which proactively and purposefully selects regions that are most helpful for answering the question.
-(3) Dynamic Attention-shifting Trigger (DAT), which triggers vision insertion into CoT when the model’s cognitive focus is significantly shifted from text to vision.
+    Necessity: Do truly crucial patches always receive high attention?
+To investigate these two questions, we randomly sample 500 instances from M3CoT and 500 from ScienceQA, respectively.
+Regarding sufficiency, we mask the top-$K_{mask}$ most attended regions and evaluate the performance drop.
+Notably, the masking is only for evaluating the importance of these regions as visual evidence, not for AIM-CoT itself.
+As shown in Table 8 (Appendix C.1), masking highly attended regions leads to only minor degradation.
+Although information redundancy may explain this, our upcoming necessity check suggests that attention peaks rarely coincide with truly salient regions, failing to be sufficient signals for evidence selection.
+Regarding necessity, we examine how well the most attended regions $R_{attn}$ cover the truly crucial regions $R_{true}$ in Section 5.5, which is quantified by Intersection over Union (IoU).
+Across 1,000 instances, over 75% have an IoU even below 50%.
+This shows that attention peaks rarely align with truly critical visual evidence, i.e., crucial regions do not necessarily emerge as the most attended ones.
+Moreover, Appendix C.2 suggests that alleviating the text-vision granularity imbalance is an effective way to improve this alignment.
+In conclusion, attention alone is unreliable for evidence selection.
+Therefore, we do not treat attention as a selection rule.
+However, improving attention distribution remains crucial, as it (1) underlies autoregressive generation in VLMs and (2) provides candidate regions for subsequent information-driven selection.
+Accordingly, we introduce CAG, which assists VLMs to attend to critical visual content precisely by enriching brief textual queries with semantic anchors (Section 4.2).
+3.2 Identifying What to See: From Semantic Alignment to Information Gain
+The empirical limitations observed in Section 3.1 can be attributed to a fundamental misalignment: cross-attention maps primarily capture the semantic correlations among tokens, whereas the ultimate goal of the selection phase in I-MCoT is to identify visual evidence that provides rich visual information for subsequent reasoning and generation.
+This insight motivates a shift from attention-based selection to an information-centric approach.
+However, two fundamental questions arise regarding the design of such a mechanism: (1) Metric: What constitutes a theoretical measure of “information” in the context of VLM reasoning?
+(2) Process: How should the selection mechanism mimic the cognitive process of information gathering, as opposed to static Top-K selection?
+To answer these, we adopt Information Foraging Theory (IFT) Pirolli and Card (1999); Broadbent (2013); Oaksford and Chater (1994); Friston (2010) as a guiding principle for our framework design.
+First, regarding the metric, IFT suggests that information is valuable only when it reduces the agent’s uncertainty.
+This motivates us to ground our selection criterion in uncertainty reduction (i.e., information gain), which can be quantified by the entropy of the VLM’s predictive distribution over its vocabulary.
+Second, regarding the process, IFT characterizes foraging as a sequential trajectory where each step depends on previous knowledge.
+This suggests that the optimal strategy is not a one-off Top-K selection, but a dynamic, iterative process where the model updates its belief state after every glimpse, i.e., the selected regions are inserted into VLM’s context in the form of visual tokens.
+(a) Regions selected by attention-based strategy.
+(b) Regions selected by information-driven strategy.
+Figure 1: A comparison of regions selected by different strategies. Details are provided in Appendix D.1.
+These two theoretical insights directly shape the architecture of our AVP module (Section 4.3).
+Figure 1 compares AVP-selected regions with those from attention-based method, which is further detailed in Appendix D.1.
+3.3 Identifying When to See: From Static Heuristics to Dynamic Attention Shifts
+The fundamental goal of I-MCoT is to ground the VLM’s reasoning in inserted fine-grained visual evidence.
+Therefore, the insertions should align with the model’s fluctuating need for visual evidence across reasoning steps.
+However, existing methods rely on static triggers like newlines Gao et al.
+(2025), failing to capture critical moments when the VLM actively seeks visual evidence to support subsequent generation.
+To address this, we propose that the attention shift from textual to visual contexts (i.e., the text-to-vision attention shift) serves as a superior, dynamic indicator.
+Intuitively, a significant pivot in attention towards the visual modality suggests the VLM’s cognitive demand for visual grounding.
+Unlike static triggers, monitoring these shifts allows for determining the precise “when” based on the model’s real-time internal state.
+We empirically validate this intuition through an in-depth analysis of the baseline model on the LLaVA-W benchmark (detailed in Appendix F).
+Our investigation correlates the timing of visual insertions with model performance, revealing two key observations: (1) Correlation analysis: Synchronizing visual insertions with salient text-to-vision attention shifts strongly correlates with better generation quality.
+(2) Group analysis: This dynamic shift pattern distinguishes high-quality responses from low-quality ones.
+In conclusion, although attention is unreliable as a decision rule for selection, its shift is still a reliable diagnostic signal revealing when the model is seeking visual information.
+Motivated by this, we present DAT, an attention-shift trigger that dynamically activates visual insertions (Section 4.4).
+Figure 2: An overview of the AIM-CoT framework.
+CAG initially enhances the query to refine attention distributions, thereby supporting reasoning and supplying reliable candidates for AVP.
+During reasoning, AVP executes visual insertion by iteratively selecting the most informative regions as evidence conditioned on the current context.
+This insertion is triggered once DAT (the trigger) detects a significant text-to-vision shift in the VLM’s internal attention state.
+See Appendix D.2 for a concise reading guide.
+4 AIM-CoT
+In this section, we begin by briefly reviewing the background of vision-language reasoning.
+Then, based on the motivations detailed in Section 3, we present AIM-CoT.
+As a training-free framework, AIM-CoT models the interleaved-modal reasoning as an information-foraging process with three synergistic components: (1) CAG, which pre-processes the input to generate a fine-grained description, mitigating text-vision granularity imbalance.
+(2) AVP, which proactively selects regions that maximize information gain whenever activated.
+(3) DAT, which triggers AVP to insert visual evidence into the CoT precisely when the model’s cognitive focus shifts from text to vision.
+A visualization of AIM-CoT is shown in Figure 2.
+It is important to distinguish AIM-CoT from attribution methods that rely on masking or deleting parts of the input image.
+AIM-CoT operates on a frozen VLM and adopts a “Trigger-Select-Insert” paradigm: although sub-regions are integrated as visual evidence, the input context, including the original image is fully preserved.
+This ensures the VLM operates on a coherent, augmented context rather than a disrupted one Khorram et al.
+(2021).
 4.1 Preliminaries
-Vision-Language Model. A VLM typically fuses a vision encoder for preprocessing visual input and a generative language model, which jointly enable it to respond in a human-like manner as follows:
-$answer=\text{VLM}(I,x),$ (1)
-where $I$ and $x$ are the image and query, respectively.
-Multimodal CoT.
-Compared to the direct response shown in Equation 1, multimodal CoT encourages the VLM to output the thought process, called rationales, before outputting the final answer.
-Existing work Gao et al.
-(2025) has been devoted to generalizing rationales from pure text to text-vision interleaved form.
+Due to space constraints, we defer the definitions of important concepts to Appendix B, including (1) the Vision-Language Models (VLMs), (2) the Context Window of a VLM, (3) Patches and Regions (i.e., Visual Tokenization), and (4) I-MCoT.
 4.2 Context-enhanced Attention-map Generation (CAG)
-In this component, the VLM is prompted to carefully generate an explanatory description of the given image within the question before the process of VQA, with the explicit goal of helping a potential respondent correctly answer the question.
-In this process, the VLM acts as a facilitator who interprets the image in the context of the given question to guide the respondent’s thought process.
-Formally, it is expressed as follows:
-$\mathcal{D}_{CAG}=\text{VLM}(I,x,\mathcal{P}_{CAG}),$ (2)
-where $\text{VLM},I,x$ are the used VLM, the given image and question, respectively.
-$\mathcal{P}_{CAG}$ is the prompt provided for the model to generate the description.
-Then, to compensate for the sparsity of textual information within the context, the generated description $\mathcal{D}_{CAG}$ is concatenated to the question $x$ as follows:
-$x^{\prime}=\text{concat}(x,\mathcal{D}_{CAG}).$ (3)
-By enhancing the context to compensate for the sparsity of textual information, the disparity in text-vision granularity is effectively mitigated, unlocking the potential for the attention map to serve as a more reliable indicator of task-relevant regions.
-We provide an example in Appendix G.1 that details the template of $\mathcal{P}_{CAG}$, the entire process of CAG, and how the final attention map $A^{\prime}$ is generated, which is used in the next stage.
+As revealed in Section 3.1, attention is an unreliable basis for visual selection, particularly under severe text-vision granularity imbalance.
+Although AIM-CoT shifts from passive attention-based selection to proactive information-driven selection, we argue that refining the VLM’s attention distribution remains meaningful for the interleaved-modal reasoning process.
+This is primarily for two reasons: (1) a more reliable attention distribution benefits reasoning, since it serves as the basis of model generation; and (2) crucially, the attention map serves as a source of candidate regions for subsequent selection (Section 4.3).
+To this end, we propose CAG to improve the VLM’s cross-attention distribution by mitigating the text-vision granularity imbalance.
+Specifically, before the VQA process begins, the VLM is elicited to carefully generate an explanatory description of the input image conditioned on the textual query.
+Rather than appending a caption, this process enriches the brief query with more semantic anchors that encode visual information within the image, thereby alleviating the granularity imbalance.
+Formally, this is expressed as follows:
+$\mathcal{D}_{\text{CAG}}$ $=\text{VLM}(I,x,\mathcal{P}_{\text{CAG}}),$ (1)
+$x^{\prime}$ $=\text{concat}(x,\mathcal{D}_{\text{CAG}}),$ (2)
+where $\mathcal{P}_{\text{CAG}}$ is the prompt for generating the description.
+$x^{\prime}$ is the updated textual query that integrates the description $\mathcal{D}_{\text{CAG}}$.
+A template of $\mathcal{P}_{\text{CAG}}$ is provided in Appendix A.3.
+To ensure the output $\mathcal{D}_{\text{CAG}}$ serves as a reliable textual anchor while preventing error propagation to downstream modules, it is vital to mitigate potential hallucinations within it.
+To this end, the prompt $\mathcal{P}_{\text{CAG}}$ is meticulously designed with negative constraints: it explicitly instructs the model to prioritize visible evidence and adopt a rigorous, cautious stance—skipping uncertain details rather than speculating.
+Experiments in Appendix L.1 validate the effectiveness of negative constraints.
 4.3 Active Visual Probing (AVP)
-Built upon the theoretical foundation of information gain, AVP is designed to select the crucial regions from a set of candidate visual regions.
-Although AVP consists of three steps elaborated upon as follows, in terms of complexity, our provided analysis and empirical results in Appendices J and L suggest that the introduction of AVP still enables AIMCoT to strike a good balance between deployability and superior performance.
-A visualization of AVP is shown in Figure 3.
-Figure 3: An overview of our AVP module, which iteratively selects $K$ most informative regions from a diversified candidate set $C$ to build an interleaved CoT that fosters vision-language reasoning.
-Diversified Set of Candidate Regions Construction.
-Based on the first two insights summarized in Section 4, an inference can be obtained: relying solely on attention maps as the source may prevent the optimal region selection.
-Motivated by this, we propose to diversify the source of the candidate regions, thereby reducing the model’s dependency on attention maps alone.
-Specifically, we not only construct an attention-driven candidate set, $C_{attn}$, by selecting $N$ regions with the highest attention scores ($N\in\mathds{R}$), but also generate an exploratory candidate set, $C_{exp}$, by sampling $M$ grid regions uniformly at random from the input image ($M\in\mathds{R}$).
-Our empirical analysis in Appendix I shows that the incorporation of exploratory set $C_{exp}$ provides substantial salient visual regions for the VLM.
-Furthermore, we study the impact of introducing $C_{exp}$ on the performance of AIMCoT, and further compare random sampling, selective search Uijlings et al.
-(2013), and FastSAM Zhao et al.
-(2023) as methods for constructing $C_{exp}$ in Appendix M.
-Formally, the process can be expressed as follows:
-$C_{attn}$ $=\{R_{1},R_{2},\cdots,R_{N}\},~~~\text{s.t. $R_{i}$=Top-$i$ Region from the attention map $A^{\prime}$},$ (4)
-$C_{exp}$ $=\{R_{N+1},R_{N+2},\cdots,R_{N+M}\},~~~~~C=C_{attn}\cup C_{exp},$ (5)
-where $C_{attn},C_{exp}$, and $C$ are the attention-driven, exploratory, and total candidate sets, respectively.
-Subsequently, the goal is to select the most informative $K$ regions from $C$ ($K<|C|=N+M$).
-Quantification of Information Gain.
-An ideal region of input image is the one that is most informative and critical for answering the question, i.e., the one that minimizes the model’s predictive uncertainty.
-Based on this idea, the region selection problem is transformed into a sequential information gain maximization task.
-To formalize this objective, we propose the concept of basic uncertainty $U_{B}$ defined as follows:
-$U_{B}=H(Y|I,x,y_{<t})=-\sum_{y\in V}P(y|I,x,y_{<t})\text{log}_{2}P(y|I,x,y_{<t}),$ (6)
-where $I,x$ and $y_{<t}$ are the input image, question and the predicted tokens, respectively.
-$P(y|I,x,y_{<t})$ is model’s probability distribution across the vocabulary $V$ in current context without the introduction of any region $R_{i}\in C$.
-It is seen that $U_{B}$ takes the form of entropy, thereby inherently capturing the model’s uncertainty when predicting the next token $t$ without incorporating any $R_{i}\in C$.
-Similarly, the conditional uncertainty $U_{C,i}$, which illustrates the model’s uncertainty when an arbitrary region $R_{i}\in C$ is included into the context, is proposed and defined as follows:
-$U_{C,i}=H(Y|I,x,y_{<t},R_{i})=-\sum_{y\in V}P(y|I,x,y_{<t},R_{i})\text{log}_{2}P(y|I,x,y_{<t},R_{i}),$ (7)
-where $P(y|I,x,y_{<t},R_{i})$ is the model’s probability distribution across the vocabulary $V$ in the context combined with the region $R_{i}$.
-Eventually, the aforementioned conceptual formulation logically motivates the definition of our final key metric: the information gain of region $R_{i}$, which is formally defined as follows:
-$IG(\{R_{i}\})=U_{B}-U_{C,i},~~~i=1,2,\cdots,N+M.$ (8)
-Intuitively, $IG(\{R_{i}\})$ quantitatively characterizes how incorporating region $R_{i}$ reduces the model’s uncertainty for the subsequent token prediction.
-Optimal Region Selection. To achieve the target of region selection, we formalize the problem as follows:
-$\text{Maximize}~~F(S)=IG(S),~~~\text{s.t.}~S\subset C,|S|=K,$ (9)
-where $S$ is the optimal set of selected regions and $K$ is the size of $S$ serving as a hyper-parameter.
-Motivated by two crucial insights, we propose a greedy algorithm detailed in Algorithm 1 to solve this problem: (1) According to recognized works Bian et al.
-(2017); Sener & Savarese (2018); Kim et al.
-(2016); Krause et al.
-(2008); Das & Kempe (2011), the greedy algorithm is a well-established and widely-used method for maximizing functions, especially for maximizing those that exhibit a tendency towards submodularity, even if they are not submodular theoretically; (2) Our experiments in Appendix H suggest that the information gain function $F$ empirically exhibits significant approximate submodularity.
-Specifically, the selection is an iterative process consisting of $K$ steps.
-At each step, AVP proactively selects the most informative and not-yet-chosen region from the candidate set $C$ with the explicit goal of minimizing the uncertainty of the model’s answer to the question.
-The chosen region is then added to an intermediate set $R^{*}$.
-Notably, a key merit of AVP is its ability to bypass regions that, despite high attention scores, exhibit strong informational overlap with regions that have been selected.
-After $K$ rounds, the collection of all selected regions in $R^{*}$ constitutes the final optimal selection $S$.
-Substantial results in Appendices J.2 and L suggest that our framework’s average inference time is no more than 1.36 times that of an efficient baseline ICoT, and the framework scales efficiently with larger values of $K$ and $N_{C}=|C|$, respectively.
-4.4 Dynamic Attention-shifting Trigger (DAT)
-As motivated by the key observation in Section 3.3, it is vital to appropriately time the insertion of visual information when constructing a multimodal CoT, and furthermore, the shift of attention from the textual to the visual context serves as a crucial indicator.
-Motivated by these insights, we propose the DAT mechanism, which systematically evaluates the model’s attention scores on the visual context at every token $t$ generation step formulated as follows:
-$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},$ (10)
-where $C_{visual}$ is the visual information within the context; $\bar{a}_{t,i}$ is the average attention score of token $t$ towards the visual token with index $i$ across the last $N_{L}$ VLM layers.
-Drawing inspiration from the NLP community Jawahar et al.
-(2019); Tenney et al.
-(2019); Vig & Belinkov (2019), we restrict our focus to the model’s final layers, as they are responsible for capturing high-level semantic information, including semantic roles and coreference relations, where the shifting signal is presumed to be more reliable.
-In our implementation, we use the last 3 layers by default ($N_{L}=3$).
-Then, the shift of attention is formalized as follows:
-$\Delta A_{visual}(t)=A_{visual}(t)-A_{visual}(t-1),$ (11)
-which quantifies the model’s attention shift towards the visual context between generating the current token and the preceding one.
-Eventually, a hyper-parameter $\delta\in\mathds{R}$ is employed to delineate the point at which the attention shift $\Delta A_{visual}(t)$ is substantial enough to activate the AVP to insert essential visual information in the multimodal CoT.
-We detail a sensitivity analysis of the threshold $\delta$ with an emphasis on its impact on the frequency of triggering AVP and the performance of AIMCoT in Appendix K.
-5 Experiments
-5.1 Benchmarks and Baselines
-In this study, we evaluate AIMCoT on three popular and challenging VQA benchmarks, including M3CoT Chen et al.
-(2024), ScienceQA Saikh et al.
-(2022), and LLaVA-Bench In-the-Wild (LLaVA-W) Liu et al.
-(2024).
-We provide detailed introductions in Appendix C.
-To evaluate the performance of AIMCoT, we introduce the vanilla VLM w/o CoT (No-CoT) and a range of state-of-the-art methods as baseline models, including DDCoT Zheng et al.
-(2023), MMCoT Zhang et al.
-(2023), CCoT Mitra et al.
-(2024), and SCAFFOLD Lei et al.
-(2024), which generate text-only rationales.
-Furthermore, ICoT Gao et al.
-(2025), which constructs interleaved-modal CoT, is considered as well.
-The detailed introduction to them is listed in Appendix D.
-In presenting the results, we directly cite the performance reported in existing works where applicable.
-5.2 Implementation Details
-We implement AIMCoT and the baselines on Chameleon-7B Team (2024) and Qwen2-VL-7B-Instruct Wang et al.
-(2024) in two settings (both 0- and 1-shot), which aligns with the recent leading research Gao et al.
-(2025).
-The experiments are conducted on A6000 GPUs.
-The hyper-parameter settings for AIMCoT and the reproducibility statement are meticulously listed in Appendix G.2 and Section 7, respectively.
-5.3 Performance Comparison
-We evaluate the performance of AIMCoT against the state-of-the-art (SOTA) methods.
-The results shown in Table 2 clearly demonstrate the superiority of our proposed AIMCoT as it significantly outperforms all the baseline models under both 0- and 1-shot settings across all the datasets.
-Specifically, AIMCoT surpasses all baselines that generate text-only rationales, confirming the efficacy of integrating salient visual information directly into CoT.
-When compared to ICoT, which also produces interleaved text-vision CoT, AIMCoT’s superior performance underscores the importance of our three key contributions: (1) a more reliable attention map as a foundation, (2) a proactive, goal-oriented mechanism for image region selection, and (3) an intelligent trigger for inserting visual information at critical moments.
-Crucially, AIMCoT’s advantage is most pronounced on the open-ended LLaVA-W benchmark and in the 0-shot setting, which better simulate complex, real-world scenarios where the model must rely solely on its internal knowledge and reasoning.
-By emulating what can be seen as a more human-like cognitive process, AIMCoT unlocks the VLM’s foundational reasoning capabilities, enabling robust performance in novel and challenging situations.
-Table 2: Performance comparison results on three widely-used benchmarks.
-The best performances are shown in bold.
-The metric for experiments on M3CoT and ScienceQA is Accuracy (ACC.), while on LLaVA-W, the metric ROUGE-L is adopted.
-Backbone Method M3CoT (ACC.) ScienceQA (ACC.) LLaVA-W (ROUGE-L)
+Drawing on the analyses in Section 3.2, we propose AVP, an information-driven mechanism designed to proactively select the regions with valuable information.
+Inspired by Information Foraging Theory (IFT), AVP operates through three systematic steps: (1) constructing a diverse candidate region set, (2) quantifying the information gain of each candidate region, and (3) executing a sequential, greedy selection process.
+Step 1: Candidate Region Set Construction.
+As introduced in Appendix B, each input image is partitioned into multiple regions with a fixed partitioning method.
+Although the partition for a given input image is deterministic, evaluating the information gain of each of these regions could be computationally prohibitive.
+Therefore, it is necessary to construct a candidate pool, which is a subset of the raw partition.
+We propose that although the candidate region set is a subset of the raw partition, it should remain sufficiently informative for subsequent selection.
+To this end, we derive candidate regions from diverse sources, covering both the VLM’s current focus and potential blind spots, as follows:
+  • Attention-driven Set ($C_{attn}$): We select the top-$N$ regions from the CAG-enhanced cross-attention map ($A^{\prime}$), as this subset is a real-time reflection of VLM’s internal state in the latest reasoning step.
+Although the attention-based signals can inherently carry noise, analysis in Appendix I.3 suggests that our subsequent information-driven selection effectively filters out non-informative high-attention regions (i.e., noise rejection).
+  • Exploratory Set ($C_{exp}$): To mitigate the “tunnel vision” of attention maps, we extract $M$ exploratory regions by uniformly sampling from all the partitioned regions.
+Despite its simplicity, this effective strategy outperforms complex alternatives (Appendix I.1) and provides essential regions overlooked by the attention-driven set $C_{attn}$ (Appendix I.2).
+Formally, the total candidate set $C$ is constructed as follows:
+$C_{attn}$ $=\{R_{1},R_{2},\dots,R_{N}\},$
+$C_{exp}$ $=\{R_{N+1},R_{N+2},\dots,R_{N+M}\},$
+$C$ $=C_{attn}\cup C_{exp},$
+where each item denotes a specific image sub-region. For $i\leq N$, $R_{i}$ is the $i$-th most attended region in $A^{\prime}$.
+Step 2: Information Gain Quantification.
+As analyzed in Section 3.2, IFT motivates us to identify the most informative regions.
+Therefore, it is necessary to quantify the information gain of each candidate region in $C$.
+To quantify Information Gain (IG) of a candidate region, we compare the VLM’s information content before and after integrating this specific candidate (as visual tokens) into its current context.
+These pre- and post-insertion information contents are termed basic uncertainty and conditional uncertainty, respectively.
+Intuitively, they measure the VLM’s uncertainty in predicting the next token (textual) during autoregression under these two states, respectively.
+Formally, Basic Uncertainty ($U_{B}$) is defined as the entropy of the VLM’s next-token distribution given the current context, and Conditional Uncertainty ($U_{C,i}$) as the entropy after explicitly introducing candidate region $R_{i}$.
+$U_{B}=H(Y|I,x,y_{<t})$
+$=-\sum_{y\in V}P(y|I,x,y_{<t})\log_{2}P(y|I,x,y_{<t}),$ (3)
+$U_{C,i}=H(Y|I,x,y_{<t},R_{i})$
+$=-\sum_{y\in V}P(y|I,x,y_{<t},R_{i})\log_{2}P(y|I,x,y_{<t},R_{i}),$
+where $V$ is the vocabulary, and $y_{<t}$ represents the generated tokens. The information gain is thus derived as:
+$IG(\{R_{i}\})=U_{B}-U_{C,i},\;i=1,\dots,N+M.$ (4)
+Although this calculation requires a forward pass, Appendix M shows that AIM-CoT remains highly deployable, with an average inference time under $1.36\times$ that of the highly efficient baseline.
+Step 3: Sequential Selection Process.
+Finally, as analyzed in Section 3.2, drawing on the iterative nature of information foraging, where each step depends on previous knowledge, we frame region selection as a sequential trajectory.
+Specifically, in each step, we propose to select the region that provides the maximum information gain conditioned on the context updated in the last step, as outlined in Algorithm 1.
+This is because this greedy algorithm (1) intuitively moves maximally toward providing the VLM with the most visual information at each step, and besides, (2) it has theoretical support: regarding such a subset selection problem, it yields an approximation to the global optimum (i.e., the maximum information gain) for the VLM.
+We provide comprehensive analyses of this theoretical property in Appendix I.4.
+1
+Input : Total candidate set $C$, Target selection size $K$
+Output : Optimal selection set $S$
+2
+[…]
+4 for $k\leftarrow 1$ to $K$ do
+5   Compute Basic Uncertainty: $U_{B}\leftarrow H(Y|I,x,y_{<t},R^{*})$ ;
+6 for $R_{i}\in C\setminus R^{*}$ do
+7      Compute Conditional Uncertainty: $U_{C,i}\leftarrow H(Y|I,x,y_{<t},R^{*}\cup\{R_{i}\})$ ;
+8      Calculate Gain: $IG(\{R_{i}\})\leftarrow U_{B}-U_{C,i}$ ;
+9
+10  Select Best Region: $R_{next}\leftarrow\operatorname{argmax}_{R_{i}\in C\setminus R^{*}}\{IG(\{R_{i}\})\}$ ;
+11   Update Selected Set: $R^{*}\leftarrow R^{*}\cup\{R_{next}\}$ ;
+12
+[…]
+return $S$
+Algorithm 1 Greedy Algorithm for Optimal Region Selection
+4.4 Dynamic Attention-shift Trigger (DAT)
+As identified in Section 3.3, although attention remains an unreliable basis for selection, its shift serves as a reliable diagnostic signal for triggering.
+Motivated by this, we propose DAT, the trigger component of AIM-CoT.
+DAT tracks the VLM’s text-to-vision attention shifts across the autoregressive steps, as they reflect the VLM’s dynamic need for visual evidence.
+Then, based on the magnitude of each shift, DAT determines whether the visual evidence should be introduced via AVP at that step.
+Quantification of Text-to-vision Attention Shift.
+The text-to-vision attention shift is determined by the difference in attention distributions between steps.
+Intuitively, it tracks the flow of attention from the text context in the previous step to the visual context in the current step.
+Furthermore, since the sum of attention scores over these two parts (i.e., text and vision) is fixed in the VLM’s normalized attention map, it is sufficient to focus on the attention shift in the visual part.
+Formally, this can be quantified as follows:
+$\Delta A_{vision}(t)=A_{vision}(t)-A_{vision}(t-1),$ (5)
+where $A_{vision}(t)$ and $A_{vision}(t-1)$ represent the sum of attention scores allocated to all the visual tokens in the context by the currently predicted token (the $t$-th token) and the preceding token (the (t1)(t-1)-th token), respectively.
+Triggering Condition.
+DAT determines if a given shift $\Delta A_{vision}(t)$ is significant enough to warrant the provision of visual evidence via a qualitative criterion.
+In the criterion, a hyper-parameter $\delta\in\mathbb{R}$ is introduced to serve as the threshold.
+The AVP module is triggered to perform visual insertion after the $t$-th generation step if and only if $\Delta A_{vision}(t)>\delta$.
+In Appendix G.1, we provide a detailed sensitivity analysis of $\delta$ and introduce an adaptive thresholding strategy to set $\delta$ automatically.
+Critical Integration Strategy.
+A potential challenge inherent to the I-MCoT paradigm is the difficulty in eliminating noise that visual evidence might introduce.
+To address this, DAT does not naively trigger the visual evidence insertion.
+Instead, it employs a critical integration mechanism: the inserted regions are accompanied by a safety instruction, which cues the VLM to treat the visual evidence as “supplementary references”, encouraging the model to verify semantic consistency with the existing textual context.
+This effectively filters out irrelevant or hallucinated visual cues.
+We validate the effectiveness of this safety instruction in Appendix L.2.
+Table 1: Performance comparison results on three VQA benchmarks and four backbones. The best performances are shown in bold. We report Accuracy (ACC.) for M3CoT and ScienceQA, and ROUGE-L for LLaVA-W.
+Method M3CoT ScienceQA LLaVA-W
 0-shot 1-shot 0-shot 1-shot 0-shot 1-shot
-Chameleon-7B No-CoT 29.1 28.4 47.7 48.5 13.1 23.9
+[…]
+No-CoT 29.1 28.4 47.7 48.5 13.1 23.9
 DDCoT 28.6 29.8 49.8 49.2 20.2 23.1
@@
 ICoT 29.8 32.3 51.0 53.4 25.2 27.6
-AIMCoT (Ours) 31.4 32.8 53.1 54.5 29.8 32.0
-Improvement 5.50% 1.47% 4.08% 2.04% 18.25% 15.94%
-Qwen2-VL-7B No-CoT 43.6 45.4 56.3 64.4 32.7 33.5
+AIM-CoT (ours) 31.4 32.8 53.1 54.5 29.8 32.0
+Improv. +5.4% +1.5% +4.1% +2.1% +18.3% +15.9%
+[…]
+No-CoT 36.5 37.8 53.9 61.2 29.8 30.9
+DDCoT 36.8 38.6 53.2 61.6 29.0 30.4
+MMCoT 35.2 36.6 50.8 58.0 28.2 28.9
+CCoT 36.4 38.2 54.1 61.5 27.8 31.5
+SCAFFOLD 36.1 38.0 52.4 60.5 29.7 30.8
+ICoT 37.6 39.4 55.1 62.5 32.5 33.6
+AIM-CoT (ours) 39.7 41.5 56.9 64.9 35.5 36.6
+Improv. +5.6% +5.3% +3.3% +3.8% +9.2% +8.9%
+Method M3CoT ScienceQA LLaVA-W
+0-shot 1-shot 0-shot 1-shot 0-shot 1-shot
+[…]
+No-CoT 43.6 45.4 56.3 64.4 32.7 33.5
+DDCoT 42.6 45.7 55.2 64.9 31.2 32.8
 MMCoT 40.1 42.5 51.3 58.3 30.7 31.4
 CCoT 43.3 44.1 56.4 63.8 29.4 33.9
-DDCoT 42.6 45.7 55.2 64.9 31.2 32.8
 SCAFFOLD 41.7 44.9 53.7 62.5 31.8 33.1
 ICoT 44.1 46.0 56.8 65.4 34.2 35.7
-AIMCoT (Ours) 44.7 46.6 57.4 66.3 36.3 37.3
-Improvement 1.4% 1.3% 1.1% 1.3% 6.2% 4.5%
-5.4 Ablation Study
-In this section, we conduct a series of ablation studies to verify the efficacy of each component within AIMCoT. The details of settings are as follows:
-  • In w/o CAG, the VLM is directly prompted with the raw question $x$ and the paired image $I$;
-  • In w/o AVP, the AVP is replaced by the attention-driven Top-K strategy by following existing works Gao et al.
-(2025); Ge et al.
-(2025), which selects the regions with Top-K attention scores on the model’s attention map;
-  • In w/o DAT, following existing research Gao et al. (2025), the insertion of visual information in CoT is triggered when the model outputs the signal token, which is a line break by default.
-Ablation results in Table 3 validate the contributions of our core components.
-First, CAG provides essential context enhancement, proving crucial for generating high-quality CoT, particularly when text queries are sparse (M3CoT, LLaVA-W).
-Second, the consistent, significant performance drop when replacing AVP with a Top-K baseline underscores the substantial superiority of our proactive, information-oriented method.
-Furthermore, the removal of DAT also results in a considerable performance decline, highlighting the critical importance of the timing of image insertion.
-Table 3: Ablation study of AIMCoT conducted on Chameleon-7B under 0-shot setting.
-Dataset AIMCoT w/o CAG w/o AVP w/o DAT
+AIM-CoT (ours) 44.7 46.6 57.4 66.3 36.3 37.3
+Improv. +1.4% +1.3% +1.1% +1.4% +6.1% +4.5%
+[…]
+No-CoT 55.2 56.8 73.5 77.2 40.5 41.8
+DDCoT 54.8 57.5 72.9 77.8 39.6 41.3
+MMCoT 53.4 54.7 69.8 74.5 38.9 39.2
+CCoT 55.3 56.9 73.4 77.1 37.8 42.4
+SCAFFOLD 54.2 56.7 71.6 76.3 40.1 41.5
+ICoT 56.9 59.1 75.1 79.2 43.4 44.7
+AIM-CoT (ours) 58.7 61.2 76.8 81.3 46.5 49.1
+Improv. +3.2% +3.6% +2.3% +2.7% +7.1% +9.8%
+Statistical significance: $^{\dagger}p<0.05$, $^{\ddagger}p<0.01$ for AIM-CoT compared with the second-best method under the same backbone/setting (McNemar for ACC.
+on M3CoT/ScienceQA; Wilcoxon signed-rank for ROUGE-L on LLaVA-W).
+Full results are reported in Appendix E.
+5 Experiments
+5.1 Evaluation Setup
+Benchmark.
+We evaluate AIM-CoT on three widely used VQA benchmarks: M3CoT Chen et al.
+(2024a), ScienceQA Saikh et al.
+(2022), and LLaVA-W Liu et al.
+(2024).
+Detailed descriptions are provided in Appendix A.1.
+Baselines.
+We compare against several text-only baselines, including vanilla VLM without CoT (No-CoT), DDCoT Zheng et al.
+(2023), MMCoT Zhang et al.
+(2024), CCoT Mitra et al.
+(2024), and SCAFFOLD Lei et al.
+(2025).
+On top of these, the leading I-MCoT framework ICoT Gao et al.
+(2025), is also included.
+More details are available in Appendix A.2.
+When possible, we directly report figures from prior work.
+Backbones.
+The implementation is built on four mainstream VLM backbones with different architectures and scales: early-fusion Chameleon-7B Team (2024) and Janus-Pro-7B Chen et al.
+(2025), and late-fusion Qwen2-VL-7B Wang et al.
+(2024) and Qwen2.5-VL-32B Bai et al.
+(2025).
+We conduct experiments in both 0- and 1-shot settings, using the prompt template from the open-source ICoT implementation Gao et al.
+(2025).
+Hyper-parameter. Hyper-parameter settings are listed in Appendix A.4.
+Table 2: Ablation study of AIM-CoT conducted on Chameleon-7B under 0-shot setting.
+Dataset AIM-CoT w/o CAG w/o AVP w/o DAT
 M3CoT (ACC.) 31.4 30.5 (-0.9) 30.6 (-0.8) 30.8 (-0.6)
@@
 LLaVA-W (ROUGE-L) 29.8 26.8 (-3.0) 26.2 (-3.6) 27.3 (-2.5)
-The performance gains from AIMCoT are most pronounced on LLaVA-W, a challenging benchmark requiring open-ended generation.
-This large improvement starkly demonstrates our model’s advanced capability to comprehend intricate multimodal information and tackle demanding, unconstrained tasks.
+5.2 Performance Comparison
+Table 1 reports the experimental results.
+AIM-CoT consistently outperforms all baselines (both text-only and I-MCoT) across all benchmarks and backbones, under both 0- and 1-shot settings.
+This demonstrates the strong advantage of AIM-CoT for complex vision-language reasoning (besides, the analyses in Appendix M show that AIM-CoT maintains good deployability).
+More specifically, both I-MCoT methods (AIM-CoT and ICoT) surpass the text-only baselines, confirming the efficacy of explicit incorporation of visual evidence.
+Moreover, although AIM-CoT and ICoT follow the same I-MCoT paradigm, AIM-CoT achieves stronger performance.
+This gain comes from three key improvements: (1) AIM-CoT uses CAG to directly refine the VLM’s cross-attention distribution, instead of grounding in the unreliable signal used by ICoT; (2) it replaces fragile attention-based heuristics with information-driven selection (AVP); and (3) it replaces a static trigger with a dynamic and more precise trigger (DAT).
+Finally, the performance gains vary across backbones. In Appendix K, we analyze this effect in detail and attribute it to the interaction between model architecture and model scale.
+5.3 Ablation Study
+In this section, we conduct the ablation study to verify the efficacy of each component within AIM-CoT. The detailed settings are as follows:
+  • w/o CAG: The VLM operates under the image $I$ and the raw query $x$, instead of the CAG-enhanced query $x^{\prime}$.
+  • w/o AVP: The information-driven selection (AVP) is replaced by attention-based selection.
+  • w/o DAT: DAT is substituted with a static trigger (i.e., visual evidence is inserted whenever the VLM outputs a newline character).
+The results on Chameleon-7B are shown in Table 2 (more results and analyses on other backbones are deferred to Appendix J).
+First, refining the VLM’s attention distribution via CAG can strongly improve VLM’s performance on VQA tasks.
+This is because the attention is not only the basis of model generation but also one source of candidate evidence.
+Furthermore, (1) even without CAG, AIM-CoT still significantly outperforms ICoT (Table 1); and (2) removing AVP and DAT causes a larger performance drop than removing CAG.
+These results highlight the importance of tackling the what to see and when to see it questions.
+In response to these two issues, we propose AVP and DAT, respectively.
+To provide a comprehensive understanding of AIM-CoT, we go beyond merely validating component efficacy.
+In Sections 5.55.6, and Appendix H, we present both quantitative and visualized qualitative analyses of AVP, elucidating why information-driven regions are more effective than attention-based regions.
+5.4 Interplay between CAG and AVP
+In this section, we investigate the interaction between CAG and AVP. This is achieved by sequentially adding them to a basic model (BM) (i.e., AIM-CoT stripped of all its components).
+As shown in Table 3, while CAG and AVP yield individual gains, their combination exhibits a clear phenomenon of super-additivity, i.e., the joint improvement consistently exceeds the sum of their separate contributions.
+Specifically, on M3CoT, the joint gain ($1.0\%$) significantly surpasses the linear sum of individual gains ($0.3\%+0.4\%=0.7\%$).
+This trend holds across all benchmarks (e.g., $1.7\%>1.4\%$ on ScienceQA).
+This non-linear boost validates the interplay between the modules: CAG refines the VLM’s attention distribution, which amplifies AVP’s effectiveness; the AVP, in turn, precisely mines the salient visual evidence from the attention-driven candidates ($C_{attn}$) enhanced by CAG.
 Table 3: Ablation study of the baseline model (BM) on Chameleon-7B under 0-shot setting.
 Dataset BM BM w/ CAG BM w/ AVP BM w/ CAG, AVP
-M3CoT (ACC.) 29.8 30.3 (+0.5) 30.2 (+0.4) 30.8 (+1.0)
-ScienceQA (ACC.) 51.0 52.0 (+1.0) 51.9 (+0.9) 52.7 (+1.7)
+M3CoT (ACC.) 29.8 30.1 (+0.3) 30.2 (+0.4) 30.8 (+1.0)
+ScienceQA (ACC.) 51.0 51.5 (+0.5) 51.9 (+0.9) 52.7 (+1.7)
 LLaVA-W (ROUGE-L) 25.2 25.8 (+0.6) 26.4 (+1.2) 27.3 (+2.1)
-5.5 In-depth Analysis: the Interplay between CAG and AVP
-Our proposed components collaborate organically to foster the construction of multimodal CoT: as a preceding module, CAG enriches the context to benefit the construction of the candidate set; subsequently, AVP triggered by DAT proactively selects the most salient regions from the candidate set.
-In this section, we investigate the interaction between CAG and AVP via an ablation study.
-Starting from the baseline ICoT model (BM), i.e., AIMCoT stripped of all proposed modules, we sequentially add CAG and then AVP.
-Table 4 presents the results, from which we derive the following insights:
-  • The compatibility between CAG and AVP: both CAG and AVP individually provide significant gains, but their combination synergistically improves the construction of the interleaved CoT.
-  • The consistent superiority of AVP over Top-K selection: AVP consistently and significantly outperforms the standard Top-K selection method for choosing image regions, both with and without the presence of the CAG module.
-  • The interplay between CAG and AVP: The average performance improvement of AVP over Top-K selection increases from 2.62% to 2.94% when CAG is introduced.
-This suggests CAG enhances the source attention map, providing a more reliable set of candidate regions and thereby unlocking AVP’s full potential to select the most salient visual evidence.
+5.5 Quantitative Analysis: Semantic Relevance
+We further examine whether the visual regions selected by AIM-CoT match human-centered perception and the needs of the task.
+Our premise is that informative evidence should correspond to distinct and complete semantic entities rather than repetitive background patterns.
+We therefore use the Segment Anything Model (SAM) Kirillov et al.
+(2023) as a proxy to assess whether the selected regions align with meaningful object concepts.
+We conduct this evaluation on M3CoT and ScienceQA, randomly sampling 500 instances from each benchmark.
+Results are averaged over the four backbones: Chameleon-7B, Janus-Pro-7B, Qwen2-VL-7B, and Qwen2.5-VL-32B.
+To build a reference for visually relevant evidence, we use SAM to generate segmentation masks for the key entities mentioned in the question and ground-truth answer.
+We then define the Semantic Hit Rate (SHR) as the percentage of instances in which the selected regions substantially overlap with the corresponding SAM masks, measured by IoU greater than 0.5.
+Table 4: Comparison of SHR on M3CoT and ScienceQA (subset of 500 samples each). The results are averaged across four backbones.
+Method Selection Strategy SHR $\uparrow$
+M3CoT ScienceQA
+ICoT Gao et al. (2025) Attention-driven Top-K 18.7% 24.3%
+AIM-CoT (Ours) Info. Gain-driven AVP 65.2% 71.8%
+Table 4 shows that the attention-driven Top-K strategy used by ICoT struggles to recover semantically relevant evidence, with SHR dropping to 18.7% on M3CoT.
+This supports the observation that raw attention maps often drift toward high-contrast background patterns or generic salient objects, rather than the specific evidence required for reasoning.
+In contrast, AIM-CoT achieves substantially higher SHR scores, reaching 65.2% on M3CoT and 71.8% on ScienceQA.
+These results indicate that our active information-seeking paradigm filters out distractors and selects regions that are both structurally coherent and semantically aligned with the question, making them closer to human judgments of relevant evidence.
+5.6 Quantitative Analysis: Alignment with Human Intuition
+We next evaluate whether the visual regions selected by AIM-CoT better align with human intuition than those selected by the attention-driven baseline.
+Following prior work, we use GPT-4v Achiam et al.
+(2023) as an expert proxy judge, as strong multimodal judges have been shown to correlate well with human preferences Huang et al.
+(2025); Gera et al.
+(2025); D’Souza et al.
+(2025); Findeis et al.
+(2025); Zhen et al.
+(2025).
+We evaluate 500 randomly sampled instances from the M3CoT and LLaVA-W benchmarks.
+For each instance, GPT-4v is given the textual query, the original image, and two anonymized sets of selected regions: one produced by ICoT (Top-K) and the other by AIM-CoT (AVP).
+We randomize the order of the two sets to eliminate position bias.
+GPT-4v then compares them along three criteria: semantic relevance, object completeness, and overall helpfulness.
+The evaluation prompt is shown in Table 5.
+Table 5: The core prompt used for GPT-4v blind pairwise comparison.
+System Prompt: You are an expert judge.
+Compare two sets of image crops (Set A and Set B) extracted from the Original Image.
+Decide which set is more helpful for a human to answer the Question.
+Criteria: 1.
+Relevance: Does the set contain the specific objects or details mentioned in the question?
+2.
+Completeness: Are the objects complete, or are they meaningless background noise/fragments?
+3.
+Helpfulness: Which set would better help a human answer the question without seeing the full image?
+Output: ’Set A is better’, ’Set B is better’, or ’Tie’.
+Table 6: GPT-4v preference rates on 500 samples from M3CoT and LLaVA-W. “Win” denotes that AIM-CoT provides better visual evidence than ICoT. “Tie” indicates equal quality.
+Benchmark AIM-CoT Win Tie ICoT Win
+M3CoT (Reasoning) 76.4% 16.2% 7.4%
+LLaVA-W (In-the-Wild) 81.2% 13.6% 5.2%
+As shown in Table 6, AIM-CoT is strongly preferred across both benchmarks.
+It wins on 76.4% of M3CoT samples and 81.2% of LLaVA-W samples, while only a small fraction favor ICoT.
+These results suggest that AIM-CoT more consistently identifies complete and semantically informative evidence that matches human intuition.
+GPT-4v’s qualitative judgments further indicate that ICoT often focuses on noisy regions, whereas AIM-CoT more reliably localizes the full objects needed for reasoning.
 6 Conclusion
-In this paper, we propose AIMCoT, a novel framework that reframes the construction of interleaved-modal CoT as an active, information-foraging process, addressing the limitations in existing methods, which often rely on passive, heuristic-driven mechanisms for selecting and inserting visual information at suboptimal moments.
-Our extensive experiments on three popular and challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods in both 0- and 1-shot settings (up to 18%).
-By dynamically structuring its reasoning and actively seeking the most informative visual cues, AIMCoT achieves a more proactive, goal-oriented, and human-like approach to vision-language reasoning.
-Despite the strong performance, AIMCoT presents avenues for future exploration.
-The AVP module, while highly effective and optimized, introduces a slight computational overhead compared to simpler attention-based selection.
-Future work could explore lightweight, learnable policies for region selection to further enhance its deployability.
-We also plan to extend our evaluation to a broader range of VLM architectures and more complex, long-form reasoning tasks to further probe the generalizability and limits of our active information-seeking paradigm.
-7 Reproducibility Statement
-To ensure the reproducibility of our work, we provide the complete source code for our AIMCoT framework and all experiments in the supplementary materials, accessible via an anonymized link: https://anonymous.4open.science/r/AIMCoT.
-The architectural details and theoretical underpinnings of our proposed components, including Context-enhanced Attention-map Generation (CAG), Active Visual Probing (AVP), and Dynamic Attention-shifting Trigger (DAT), are thoroughly described in Section 4 of the main paper.
-The specific greedy algorithm employed by AVP is detailed in Algorithm 1 (Appendix E).
-All datasets used in our evaluation are publicly available benchmarks, as detailed in Section 5.1 and Appendix C.
-We provide a comprehensive list of all hyper-parameter settings used to achieve the reported results for each benchmark in Appendix G.2 (Table 5).
-Furthermore, extensive ablation studies (Section 5.4), and in-depth analyses (Section 5.5, Appendices F,H, I, K, and L), are provided to allow for a complete replication of our findings.
-Appendix A LLM Usage Statement
-In the preparation of this manuscript, we utilized a large language model (LLM) as a general-purpose writing assistance tool. The primary uses of the LLM were twofold:
-  • Grammatical Correction: The LLM was employed to proofread the manuscript for grammatical errors, spelling mistakes, and awkward phrasing.
-This helped improve the overall clarity and readability of our paper.
-  • Text Condensation: To adhere to the page limits of the conference, we used the LLM to help condense and rephrase certain paragraphs and sentences.
-All suggestions provided by the LLM were carefully reviewed, critically evaluated, and manually edited by the authors to ensure that the scientific integrity and intended meaning of our work were preserved.
-Notably, the LLM was not used for core research activities, including the ideation of the AIMCoT framework, the design of experiments, the analysis of results, or the generation of the primary scientific claims.
-The final intellectual content and all contributions presented in this paper are entirely our own.
-Appendix B Explanation for Figure 1
-Figure 1 shows the images of the 22nd question on LLaVA-W benchmark.
-The query for this question is: "What’s the name of the restaurant serving these dishes?" and the image is a close-up photo of a meal at ICHIRAN.
-The left figure is the original image, and the right figure visualizes the regions selected by the Top-K strategy.
-The num_selected_patchesnum\_selected\_patches regions with the highest, second-highest, and third-highest scores are designated as the first, second, and third sets, colored red, purple, and blue, respectively.
-The first set is utilized by the baseline model according to its default design (num_selected_patches=72num\_selected\_patches=72).
-Appendix C Introduction to the Benchmarks
+In this paper, we propose AIM-CoT, a novel I-MCoT framework that aims to frame the construction of interleaved-modal CoT as an active information-foraging process.
+Existing methods’ static triggers fail to capture the dynamic needs of the VLM for visual information in complex reasoning, and their attention-based selectors depend heavily on the VLM’s attention, which can be unreliable (for selection).
+In response to these challenges, AIM-CoT dynamically monitors the VLM’s cognitive need for fine-grained visual evidence for subsequent generation, and accordingly, selects salient visual evidence in an information-driven manner.
+Extensive experiments demonstrate that AIM-CoT outperforms the leading methods across three benchmarks and four VLM backbones.
+7 Acknowledgements
+This work was partially supported by the Guangdong Basic and Applied Basic Research Foundation under Grant No.
+2024A1515011949 and 2026A1515011672, the Shenzhen Science and Technology Program under Grant No.
+GXWD20231130110308001 and JCYJ20250604145617023.
+Appendix A Experimental Setup and Reproducibility
+A.1 Benchmarks
 M3CoT Chen et al.
-(2024) is a novel multimodal CoT benchmark, which introduces complex, multi-step problems across science, mathematics, and commonsense domains, comprising 11,459 samples in total.
+(2024a) is a novel multimodal CoT benchmark, which introduces complex, multi-step problems across science, mathematics, and commonsense domains, comprising 11,459 samples in total.
 M3CoT is characterized by succinct textual queries (<15 tokens on average) paired with intricate problems.
@@
 LLaVA-W is exceptionally well-suited for evaluating the capability of our proposed framework to address complex, open-ended problems by generating a multimodal CoT, attending to salient regions within the image, and meticulously parsing the query.
-Appendix D Introduction to the Baseline Models
+A.2 Baselines
 No-CoT prompts the VLM to answer questions directly based on the input query and image. In the 1-shot setting, an example containing the query, image, and corresponding answer is attached.
@@
 (2025) leverages the attention maps of the VLM to select relevant patches from the input image and insert them into the reasoning process, thereby generating sequential steps of paired visual and textual rationales.
-Appendix E Greedy Algorithm within AVP Module
-The complete process of the greedy algorithm within AVP is shown in Algorithm 1.
-Input: total candidate set $C$, size of optimal selection $K$
-Output: optimal selection $S$
-1 $R^{*}\leftarrow\varnothing$
-2 for $k\leftarrow 1,2,\cdots,K$ do
-3 $U_{B}\leftarrow H(Y|I,x,y_{<t},R^{*})$
-4 for $i\leftarrow 1,2,\cdots,N+M$ do
-5    $U_{C,i}\leftarrow H(Y|I,x,y_{<t},R^{*}\cup\{R_{i}\})$
-6    $IG(\{R_{i}\})\leftarrow U_{B}-U_{C,i}$
-7 $R_{next}\leftarrow\text{argmax}_{R_{i}\in C\backslash R^{*}}\{IG(\{R_{i}\})\}$
-8 $R^{*}\leftarrow R^{*}\cup\{R_{next}\}$
-9
-[…]
-11 return $S$
-Algorithm 1 Greedy Algorithm for Optimal Region Selection
-Appendix F Detailed Analysis of Key Moments to Insert Visual Information
-Experimental Setup.
-We take ICoT Gao et al.
-(2025) as a baseline model, which is required to answer all questions from the LLaVA-W benchmark in a 0-shot setting, with ROUGE-L used as the evaluation metric.
-The hyper-parameters follow the default settings of the open-source implementation for ICoT, and all experiments are conducted with the Chameleon-7B backbone.
-Formal Definition of Attention Shifts.
-To analyze attention shifts, we examine the averaged attention maps across all attention heads in the last three layers of the VLM during the prediction of each token $t$.
-The model’s total attention scores allocated to the visual and text components of the input are respectively measured as follows:
-$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},\hskip 14.22636ptA_{text}(t)=\sum_{j\in\text{indices of $C_{text}$}}\bar{a}_{t,j},$ (12)
-where $C_{visual},C_{text}$ are the visual and text information within the context, respectively.
-Then, the shift in attention from the textual to the visual modality while generating token $t$ is defined as follows:
-$\delta_{t}=A_{visual}(t)-A_{visual}(t-1).$ (13)
-$\Delta_{k}=[\delta_{1},\delta_{2},\cdots,\delta_{|\Delta_{k}|}]$ encompasses the model’s attention shifts for each token when answering the arbitrary $k$-th question, where $|\Delta_{k}|$ is the number of tokens for answering the $k$-th question.
-Formal Definition of Scores.
-for the predictions generated by the baseline model, the ROUGE-L scores are given by $List\_R=[R_{1},R_{2},\cdots,R_{|List\_R|}]$, where $R_{k}$ is the score for the model’s response to the $k$-th question, and |List_R||List\_R| is the number of questions within the benchmark.
-Based on these concepts, we design a two-part experiment:
-Experiment 1: Correlation Analysis.
-We investigate the relationship between the proportion of visual insertions under significant attention shifts and the score of the corresponding generated prediction.
-First, to identify whether a visual insertion is conducted during a significant attention shift, we define a high attention growth threshold, $\delta_{k}^{(h)}$ for the $k$-th response ($\delta_{k}^{(h)}$ is set to the 80% upper quantile of $\Delta_{k}$ by default).
-An insertion is considered to have been conducted under a significant shift and referred to as a synchronized insertion if and only if its corresponding attention shift value exceeds the threshold $\delta_{k}^{(h)}$.
-Next, since the model can conduct multiple insertions per response for a question, we calculate $P_{k}$, the proportion of synchronized insertions out of the total number of insertions for the $k$-th question.
-Finally, since the proportions of synchronized insertions $P_{1},P_{2},\cdots,P_{|List\_R|}]$ and the ROUGE-L scores for all the questions $R_{1},R_{2},\cdots,R_{|List\_R|}]$ are obtained, the Pearson Correlation coefficient can be computed.
-Specifically, the Pearson Correlation is 0.2166 with a p-value of 0.048, which suggests that the proportions of the synchronized insertions and the corresponding score are significantly positively related to each other.
-Experiment 2: Group Analysis. We investigate the relationship between the proportion of synchronized insertions and the quality of the model’s response.
-To group the generated predictions according to response quality, we establish high- and low-scoring groups.
-All predictions are ranked in descending order by their ROUGE-L scores.
-The top 30% form $G_{h}$, the high-scoring group (high-quality responses), and $G_{l}$, the bottom 30% form the low-scoring group (low-quality responses).
-Then, we calculate the mean proportion of synchronized insertions for groups $G_{h},G_{l}$, which are denoted as $\bar{P}_{h},\bar{P}_{l}$, respectively.
-Finally, the means of the two groups $\bar{P}_{h},\bar{P}_{l}$ are compared, and a T-test is performed to assess the statistical significance of the difference.
-Specifically, we find that $\bar{P}_{h}=0.8889,\bar{P}_{l}=0.5000$, which suggests that in the high-scoring group, approximately 89% of insertions are the synchronized insertions with significant attention shift from textual input to visual information; in contrast, in the low-scoring group, only about half of the insertions are synchronized insertions.
-Besides, the P-value of T-test is as low as $0.0019$, which demonstrates that the result is highly statistically significant.
-Appendix G Model Settings
-G.1 Template of $\mathcal{P}_{CAG}$
+A.3 Template of $\mathcal{P}_{CAG}$
 Figure 3 provides an intuitive example showing the template of $\mathcal{P}_{CAG}$ and how it is used to prompt the VLM to carefully generate a guiding description for the input image.
-In particular, for multiple-choice questions, such as those in M3CoT and ScienceQA, we prepend the following brief explanation to $\mathcal{P}_{CAG}$ to aid the VLM in better understanding its designated task: "This is a multiple-choice question.
-The question is based on the image provided."
+In particular, for multiple-choice questions, such as those in M3CoT and ScienceQA, we prepend the following brief explanation to $\mathcal{P}_{CAG}$ to aid the VLM in better understanding its designated task: “This is a multiple-choice question.
+The question is based on the image provided.
 Figure 3: An illustration of the entire process of context enhancement by the CAG module, using problem physical-commonsense-1398 from the M3CoT benchmark as an example.
@@
 $W^{Q},W^{K}$ are the weight matrices of the linear transformation layers for query and key, respectively.
-G.2 Setting of Hyper-parameters
-The hyper-parameter settings are presented in Table 5.
-For easier understanding, we provide further explanation of the hyper-parameters $s_{r}$ and $s_{g}$ as follows: in the AVP module, the input image is first divided into $s_{g}\times s_{g}$ grids according to the set "Grid size" $s_{g}$.
-Then, each region finally selected by AVP consists of $s_{r}\times s_{r}$ grids.
+A.4 Setting of Hyper-parameters
+The hyper-parameter settings are listed in Table 7.
 Table 7: Hyper-parameter settings across three datasets.
@@
 Max_new_tokens 512 1024 1024
-Appendix H Empirical Analysis of the Approximate Submodularity of Function $F$
+Appendix B Supplementary Important Definitions
+Vision-Language Model.
+A VLM typically fuses a vision encoder for preprocessing visual input and a generative language model, which jointly enable it to respond in a human-like manner:
+$answer=\text{VLM}(I,x),$ (7)
+where $I$ is the image and $x$ the query.
+Context Window of a VLM.
+An autoregressive generation of a VLM is conditioned on (1) the user’s initial input and (2) its interaction with the user.
+A context window is a structured representation of this conditional information.
+Specifically, a context window starts with the user’s multimodal input question, including a textual query $x$ and a paired image $I$.
+Built upon this, the VLM continuously extends the initial context window by iteratively incorporating each reasoning step and its corresponding visual evidence (if any) into this context, once they are acquired.
+This ensures that the prior part of the context window can consistently yet iteratively influence the subsequent generation and the final response.
+In implementation, this multimodal information, whether textual or visual context, is incorporated into the context window in the form of tokens.
+Visual Tokenization: Patches and Regions.
+In VLM’s visual encoder preprocessing, the input image $I$ is transformed into a sequence of patches and visual tokens.
+First, $I$ is divided into a fixed number of non-overlapping and equal-sized patches.
+Each patch is a local fragment of $I$, which is then projected into a corresponding visual token, serving as the fundamental atomic unit for the model’s attention mechanism.
+A region is a spatial crop comprised of multiple spatially contiguous (i.e., neighboring) patches.
+Compared to a single patch, a region covers a larger area to capture higher-level, concrete semantic information that a single atomic unit (i.e., patch) lacks.
+Compared to the entire image $I$, a selected region enables the provision of finer-grained visual details.
+Mapping from Patches to Regions.
+For a given input image, and given hyperparameters $s_{r}$ and $s_{g}$, the mapping from patches to regions is deterministic.
+However, this relies on an intermediate concept, the “grid”.
+Specifically, in the AVP module, the input image is first divided into $s_{g}\times s_{g}$ grids according to the set “Grid size” $s_{g}$.
+Each grid is comprised of multiple patches.
+For example, Chameleon Team (2024) divides the input image into $32\times 32=1024$ patches.
+Therefore, each grid contains $(32/s_{g})\times(32/s_{g})=1024/{s_{g}}^{2}$ patches.
+Furthermore, each region is a group of spatially continuous grids, consisting of $s_{r}\times s_{r}$ grids.
+Interleaved-Modal CoT (I-MCoT).
+Built upon a VLM backbone, the I-MCoT paradigm fundamentally relies on two key components: selection and triggering.
+Compared to direct response (Equation 7), the trigger mechanism temporarily pauses the textual autoregressive process when specific conditions are met.
+During this suspension, the I-MCoT method selects salient visual evidence from the input image, which is then tokenized and concatenated into VLM’s context to inform and support subsequent reasoning.
+Appendix C Supplementary Details Regarding Motivation in Section 3.1
+C.1 Experiment for Sufficiency Check
+In the sufficiency check in Section 3.1, we assess the impact of the most attended regions on the baseline model (i.e., ICoT Gao et al. (2025)) by masking them out.
+Experimental Setup.
+We conduct experiments with four VLM backbones (Chameleon-7B Team (2024), Janus-Pro-7B Chen et al.
+(2025), Qwen2-VL-7B Wang et al.
+(2024), and Qwen2.5-VL-32B Bai et al.
+(2025)) and report results averaged across all backbones.
+We adopt the zero-shot setting.
+Experimental Results.
+The results are shown in Table 8.
+As $K_{\text{mask}}$ increases, performance drops slightly, but the decline remains small.
+Even when $K_{\text{mask}}=50$, the model does not exhibit a substantial degradation in performance.
+Table 8: Performance difference of the baseline model (ICoT, 0-shot) when the Top $K_{\text{mask}}$ regions on the attention map are masked.
+$K_{\text{mask}}$ 0 10 20 30 50
+M3CoT 0% +0.05% -0.11% -0.36% -0.73%
+ScienceQA 0% -0.03% -0.08% -0.26% -0.55%
+C.2 Experiment for Text-vision Granularity Imbalance in Necessity Check
+As mentioned in the necessity check in Section 3.1 and analyzed in Section 5.5, in the vast majority of cases (specifically, over 75%), the VLM fails to distribute its attention toward the truly crucial regions, i.e., those containing salient task-relevant information.
+In this section, we present experimental evidence demonstrating that mitigating the text-vision granularity imbalance inherent in VQA benchmarks significantly enhances the VLM’s ability to accurately focus on salient visual content.
+Specifically, this is achieved by eliciting the VLM to extract the visual information from the informative image in textual form and transferring it to the otherwise brief raw query.
+Experimental Setup.
+We adhere to the experimental setup detailed in Section 5.5.
+The mitigation of granularity imbalance is implemented via CAG component proposed in Section 4.2.
+To assess whether mitigating granularity imbalance significantly improves the VLM’s attention distribution, we compare the Intersection over Union (IoU) between the most attended regions and the ground-truth crucial regions before and after the mitigation of imbalance.
+Furthermore, Wilcoxon signed-rank test is adopted to verify the statistical significance of the improvement.
+Experimental Results.
+The results are reported in Table 9.
+Evidently, by enriching the originally brief textual query with additional semantic anchors, the VLM is able to more accurately localize the critical regions within the input image.
+Moreover, we observe that this improvement becomes increasingly pronounced as the length of the CAG-generated description increases.
+The Wilcoxon signed-rank test confirms that these improvements are highly statistically significant ($p<0.001$).
+Table 9: Analysis of the impact of mitigating text-vision granularity imbalance on VLM’s attention alignment.
+Raw Query denotes the baseline using the brief textual query, while Enhanced Query incorporates the fine-grained image description generated by CAG.
+Input Context Mean IoU (%) $\Delta$ IoU p-value
+Raw Query (Baseline) 4.5 - -
+Enhanced Query (48 Tokens) 12.2 +7.7 <0.001<0.001
+Enhanced Query (96 Tokens) 18.6 +14.1 <0.001{<0.001}
+Enhanced Query (144 Tokens) 24.1 +19.6 <0.001{<0.001}
+Appendix D Explanation for Figures
+D.1 Explanation for Figure 1
+The example adopted in Section 3.2 is the 22nd question in the LLaVA-W benchmark.
+The query for this question is: “What’s the name of the restaurant serving these dishes?” and the associated image is a close-up photo of a meal at ICHIRAN, as shown in Figure 4(a).
+This is a representative and challenging example because the query is extremely brief, while the image contains highly dense information: although it includes ramen, side dishes, and tableware, the truly task-relevant visual evidence occupies only a tiny portion of the image—a small region on the left side of the bowl’s rim (Figure 4(b)).
+(a) Close-up meal photo (query: “What’s the name of the restaurant serving these dishes?”).
+(b) Ground-truth cue is localized to a tiny area on the bowl rim (red box).
+Figure 4: An intuitive example from LLaVA-W shows an ICHIRAN meal close-up.
+The left image is the raw image, while the small region containing the truly crucial information for answering the question is highlighted with the red box.
+Figure 1 compares the visual evidence selected by different methods.
+The left and right subfigures visualize the top-3 sets of attention-based regions and information-driven regions, respectively.
+The first, second, and third sets are colored red, purple, and blue, respectively; each set contains $72$ patches.
+As shown in Figure 1(a), the attention-based regions are scattered, making it difficult to convey complete and concrete visual information.
+More importantly, they fail to cover the small area that contains the truly crucial evidence.
+In contrast, as shown in Figure 1(b), the information gain-based selection guides the VLM to first focus on the inner rim of the bowl (red) accurately, where the critical information is contained.
+Although the VLM does not yield the final answer in this region, this indicates a correct line of reasoning, as the ground truth is situated in a highly similar area—a nearby location also on the inner wall of the bowl.
+Subsequently, the region ranked third (blue) precisely encompasses a large portion of the restaurant’s name, which is just the answer to the question.
+This suggests that even for a challenging case where the text-vision granularity is highly disparate, information gain serves as a better foundation for region selection.
+D.2 Quick Guide to Framework Overview
+Figure 2 illustrates AIM-CoT as a Trigger—Select—Insert loop over the VLM’s multimodal context window during Autoregressive Generation (blue, bottom).
+CAG (green, left) performs Text Context Enhancement by eliciting a fine-grained, query-conditioned image description and appending it to the original query.
+The enhanced query leads to Refined Attention, which both improves model reasoning and provides more reliable region candidates for subsequent selection.
+During decoding, DAT (yellow, top) tracks the Quantification of Text-to-vision Attention Shift; when the shift is significant (i.e., $\Delta A_{v}(t)>\delta$), the Triggering Condition is met and a visual insertion is activated.
+Once triggered, AVP (purple, middle) conducts Candidate Region Set Construction by merging attention-driven and exploratory regions, performs Information Gain (IG) Quantification (i.e., uncertainty reduction) for candidates, and runs a greedy Sequential Selection Process for $K$ steps.
+In each step, AVP selects the region with the highest information gain and inserts it into the context window; the next step’s selection is conditioned on the updated context.
+Finally, the inserted regions are accompanied by a Critical Integration Strategy (safety instruction) that treats this visual evidence as supplementary and requires consistency checking in the VLM.
+Once included in the context window, they help refine subsequent next-token predictions.
+Appendix E Statistical Significance Testing
+To rigorously validate the superiority of AIM-CoT, we conduct statistical significance tests comparing our method against the runner-up performance.
+The one-sided approach is chosen because our hypothesis specifically posits that AIM-CoT outperforms the baseline.
+M3CoT and ScienceQA.
+Since Accuracy (ACC) is derived from binary outcomes, we employ the one-sided McNemar’s Test.
+This focuses on assessing whether the number of instances where AIM-CoT corrects ICoT’s errors significantly exceeds the number of instances where it introduces new errors.
+LLaVA-W.
+For the ROUGE-L metric, which involves continuous scores, we utilize the one-sided Wilcoxon Signed-Rank Test (alternative hypothesis: AIM-CoT outperforms the runner-up).
+This non-parametric paired test evaluates whether the distribution of improvement scores is significantly positive.
+In Table 1, entries marked with $\dagger$ and $\ddagger$ denote statistical significance with $p<0.05$ and $p<0.01$, respectively. The detailed p-values and statistics are shown in Table 10.
+Table 10: Statistical significance tests comparing AIM-CoT against the runner-up under the same backbone/setting. Entries report one-sided $p$-values with test statistics in parentheses.
+Dataset 0/1-shot
+0-shot 1-shot
+[…]
+M3CoT (ACC) 0.0004(11.5) 0.0228(4)
+ScienceQA (ACC) 0.0000(22.7) 0.0246(3.87)
+LLaVA-W (R-L) 0.00001(1502) 0.00002(1477)
+[…]
+M3CoT (ACC) 0.0000(15.6) 0.0006(10.4)
+ScienceQA (ACC) 0.0003(11.7) 0.0000(23)
+LLaVA-W (R-L) 0.00012($14$) 0.00121(1327)
+Dataset 0/1-shot
+0-shot 1-shot
+[…]
+M3CoT (ACC) 0.0224(4.02) 0.0224(4.02)
+ScienceQA (ACC) 0.0223(4.03) 0.0241(3.91)
+LLaVA-W (R-L) 0.02035(1193) 0.03024(1170)
+[…]
+M3CoT (ACC) 0.0075(5.92) 0.0012(9.18)
+ScienceQA (ACC) 0.0028(7.67) 0.0107(5.3)
+LLaVA-W (R-L) 0.00039(1371) 0.00000(1538)
+Notes: $^{\dagger}p<0.05$, $^{\ddagger}p<0.01$ (one-sided).
+For ACC.
+(M3CoT/ScienceQA) we use McNemar’s test and report the $\chi^{2}$ statistic in parentheses; for ROUGE-L on LLaVA-W we use the Wilcoxon signed-rank test and report the $W$ statistic in parentheses.
+Appendix F Validation of Motivation in Section 3.3: Attention Shifts as Dynamic Triggers
+Experimental Setup.
+We take ICoT Gao et al.
+(2025) as a baseline model, which is required to answer all questions from the LLaVA-W benchmark in a 0-shot setting, with ROUGE-L used as the evaluation metric.
+The hyper-parameters follow the default settings of the open-source implementation for ICoT, and all experiments are conducted with the Chameleon-7B backbone.
+Formal Definition of Attention Shifts.
+To analyze attention shifts, we examine the averaged attention maps across all attention heads in the last three layers of the VLM during the prediction of each token $t$, following existing research Jawahar et al.
+(2019); Tenney et al.
+(2019); Vig and Belinkov (2019).
+The model’s total attention scores allocated to the visual and text components of the input are respectively measured as follows:
+$A_{vision}(t)$ $=\sum_{i\in\text{indices of $C_{vision}$}}\bar{a}_{t,i},$ (8)
+$A_{text}(t)$ $=\sum_{j\in\text{indices of $C_{text}$}}\bar{a}_{t,j},$ (9)
+where $C_{vision},C_{text}$ are the visual and text information within the context, respectively.
+Then, the shift in attention from the textual to the visual modality while generating token $t$ is defined as follows:
+$\delta_{t}=A_{vision}(t)-A_{vision}(t-1).$ (10)
+$\Delta_{k}=[\delta_{1},\delta_{2},\cdots,\delta_{|\Delta_{k}|}]$ encompasses the model’s attention shifts for each token when answering the arbitrary $k$-th question, where $|\Delta_{k}|$ is the number of tokens for answering the $k$-th question.
+Formal Definition of Scores.
+For the predictions generated by the baseline model, the ROUGE-L scores are given by $List\_R=[R_{1},R_{2},\cdots,R_{|List\_R|}]$, where $R_{k}$ is the score for the model’s response to the $k$-th question, and |List_R||List\_R| is the number of questions within the benchmark.
+Based on these concepts, we design a two-part experiment:
+Experiment 1: Correlation Analysis.
+We investigate the relationship between the proportion of visual insertions under significant attention shifts and the score of the corresponding generated prediction.
+First, to identify whether a visual insertion is conducted during a significant attention shift, we define a high attention growth threshold, $\delta_{k}^{(h)}$ for the $k$-th response ($\delta_{k}^{(h)}$ is set to the 80% upper quantile of $\Delta_{k}$ by default).
+An insertion is considered to have been conducted under a significant shift and referred to as a synchronized insertion if and only if its corresponding attention shift value exceeds the threshold $\delta_{k}^{(h)}$.
+Next, since the model can conduct multiple insertions per response for a question, we calculate $P_{k}$, the proportion of synchronized insertions out of the total number of insertions for the $k$-th question.
+Finally, since the proportions of synchronized insertions $P_{1},P_{2},\cdots,P_{|List\_R|}]$ and the ROUGE-L scores for all the questions $R_{1},R_{2},\cdots,R_{|List\_R|}]$ are obtained, the Pearson Correlation coefficient can be computed.
+Specifically, the Pearson Correlation is 0.2166 with a p-value of 0.048, which suggests that the proportions of the synchronized insertions and the corresponding score are significantly positively related to each other.
+Experiment 2: Group Analysis. We investigate the relationship between the proportion of synchronized insertions and the quality of the model’s response.
+To group the generated predictions according to response quality, we establish high- and low-scoring groups.
+All predictions are ranked in descending order by their ROUGE-L scores.
+The top 30% form the high-scoring group (high-quality responses) $G_{h}$, and the bottom 30% form the low-scoring group (low-quality responses) $G_{l}$.
+Then, we calculate the mean proportion of synchronized insertions for groups $G_{h},G_{l}$, which are denoted as $\bar{P}_{h},\bar{P}_{l}$, respectively.
+Finally, the means of the two groups $\bar{P}_{h},\bar{P}_{l}$ are compared, and a t-test is performed to assess the statistical significance of the difference.
+Specifically, we find that $\bar{P}_{h}=0.8889,\bar{P}_{l}=0.5000$, which suggests that in the high-scoring group, approximately 89% of insertions are the synchronized insertions with significant attention shift from textual input to visual information; in contrast, in the low-scoring group, only about half of the insertions are synchronized insertions.
+Besides, the p-value of t-test is as low as $0.0019$, which demonstrates that the result is highly statistically significant.
+Appendix G Hyper-parameter Sensitivity Analysis
+In this section, we conduct a comprehensive sensitivity analysis of hyper-parameters within AIM-CoT.
+Specifically, Appendix G.1 investigates the impact of the triggering threshold $\delta$ in the DAT module and introduces an adaptive thresholding strategy.
+Appendix G.2 examines the influence of the candidate set size $N_{C}$ and selection count $K$ in the AVP module on computational overhead and deployability.
+G.1 Analysis of $\delta$: Sensitivity and Adaptivity
+The hyper-parameter $\delta$ within the DAT module serves as a crucial threshold to trigger the AVP module, which inserts salient visual regions to improve the construction of the multimodal CoT.
+In this section, we provide a comprehensive analysis of this parameter from two perspectives: (1) a sensitivity analysis to understand its impact on performance and triggering frequency, and (2) an exploration of an adaptive thresholding strategy to demonstrate the framework’s robustness and generalizability without per-dataset tuning.
+Sensitivity Analysis
+We detail a sensitivity analysis of $\delta$ by adjusting it across the range of $[0.1,0.125,0.15,0.175,0.2,0.225]$ and examining not only the performance of AIM-CoT but also the number of times the AVP is triggered.
+The experiments are conducted under the 0-shot setting on the Chameleon-7B backbone and LLaVA-W benchmark.
+The experimental results are shown in Figure 5.
+The left figure illustrates that AIM-CoT exhibits limited performance when the threshold, $\delta$, is set too low.
+This underscores the importance of inserting visual information at critical moments: excessively frequent or inopportune visual insertions can disrupt the VLM’s reasoning process, leading to suboptimal performance.
+As $\delta$ increases, the model’s performance progressively improves, reaching its peak at $\delta=0.2$ (our default setting), which corresponds to a ROUGE-L score of 0.2983.
+However, a further increase in $\delta$ results in a slight performance degradation.
+This highlights the criticality of visual information insertion for constructing an interleaved Chain of Thought: an overly stringent threshold excessively impedes the incorporation of visual data, preventing the AVP from supplying the model with necessary visual supplementation in a timely manner.
+Conversely, the right figure demonstrates a consistent decrease in the number of times the AVP is triggered as $\delta$ is raised.
+This showcases the efficacy of $\delta$ as a threshold for modulating the activation frequency of the AVP.
+Figure 5: Experimental results of the sensitivity analysis of the hyper-parameter $\delta$.
+The left figure illustrates the performance of AIM-CoT when $\delta$ takes different values, while the right one shows the number of times the AVP module within AIM-CoT is triggered.
+Adaptive Thresholding Strategy
+While our main experiments utilize a fixed $\delta$ tuned for specific benchmarks (e.g., $\delta=0.5$ for M3CoT, and $\delta=0.2$ for ScienceQA and LLaVA-W), we acknowledge that the optimal threshold is intrinsically linked to the granularity of reasoning required by the task.
+M3CoT involves complex logic with sparse visual queries, leading to sharp, intense attention shifts (requiring a higher threshold to filter noise).
+In contrast, ScienceQA and LLaVA-W involve generating explanations or descriptions that require continuous visual grounding, resulting in smoother attention shifts (favoring a lower threshold).
+To address this variance and enhance the deployability of AIM-CoT without requiring per-dataset tuning, we propose and evaluate an Adaptive Z-Score Triggering mechanism.
+Instead of a fixed absolute threshold, this method employs a dynamic relative threshold based on the statistical properties of the attention shifts within the current generation context.
+Specifically, we calculate the Z-score of the current attention shift $\Delta A_{vision}(t)$ relative to the moving average of the previous $W$ tokens:
+$Z_{t}=\frac{\Delta A_{vision}(t)-\mu_{t-W:t-1}}{\sigma_{t-W:t-1}+\epsilon},$ (11)
+where $\mu$ and $\sigma$ are the mean and standard deviation of the attention shifts over a window $W$ (set to 5 by default), and $\epsilon$ is a small constant for numerical stability.
+The trigger is activated when $Z_{t}>\lambda$, where $\lambda$ is a universal sensitivity parameter (set to 3).
+Table 11: Performance comparison between the optimal Fixed Threshold (Tuned) and the Adaptive Z-Score Strategy across three backbones (0-shot setting).
+Dataset Method Setting Performance
+Chameleon-7B Janus-Pro-7B Qwen2.5-VL-32B
+M3CoT (ACC.) Fixed Threshold $\delta=0.5$ 31.4 39.7 58.7
+Adaptive Z-Score $\lambda=3$ 31.1 39.5 58.6
+ScienceQA (ACC.) Fixed Threshold $\delta=0.2$ 53.1 56.9 76.8
+Adaptive Z-Score $\lambda=3$ 52.9 56.7 76.8
+LLaVA-W (ROUGE-L) Fixed Threshold $\delta=0.2$ 29.8 35.5 46.5
+Adaptive Z-Score $\lambda=3$ 29.5 35.4 46.3
+As presented in Table 11, the Adaptive Z-Score strategy demonstrates remarkable robustness.
+On the powerful Qwen2.5-VL-32B backbone, the adaptive method matches the performance achieved with the tuned threshold on ScienceQA (76.8) and shows negligible deviation on M3CoT (-0.1), suggesting that stronger models exhibit more distinct attention patterns that are easier to capture dynamically.
+Even on smaller models like Janus-Pro-7B, the performance gap remains minimal (e.g., only 0.1 drop on LLaVA-W).
+This confirms that while a tuned fixed threshold offers a slight edge, the adaptive strategy provides a highly competitive and deployment-friendly alternative.
+Table 12: Unified sensitivity analysis of hyper-parameters $N_{C}$ and $K$ across 2 benchmarks (M3CoT and LLaVA-W) on 2 VLM backbones (Chameleon-7B and Qwen2.5-VL-32B).
+The table reports AIM-CoT’s average processing time (seconds) per instance.
+Entries where $N_{C}<K$ are marked with “-” as $N_{C}\geq K$ holds by definition.
+$N_{C}$ M3CoT LLaVA-W
+K=1 K=3 K=5 K=7 K=1 K=3 K=5 K=7
+Backbone: Chameleon-7B
+1 12.78 - - - 10.61 - - -
+2 12.80 - - - 10.70 - - -
+3 12.83 13.18 - - 10.75 11.21 - -
+4 12.87 13.23 - - 10.84 11.36 - -
+5 12.90 13.30 13.41 - 10.92 11.51 11.79 -
+6 12.93 13.37 13.49 - 11.00 11.65 11.96 -
+7 12.97 13.42 13.56 13.66 11.08 11.80 12.14 12.38
+Backbone: Qwen2.5-VL-32B
+1 52.12 - - - 42.45 - - -
+2 52.31 - - - 42.68 - - -
+3 52.48 55.08 - - 42.91 45.75 - -
+4 52.75 55.45 - - 43.15 46.12 - -
+5 52.94 55.72 56.95 - 43.52 46.48 47.95 -
+6 53.18 55.98 57.34 - 43.88 46.85 48.44 -
+7 53.42 56.23 57.68 58.71 44.15 47.27 48.95 50.10
+G.2 Sensitivity Analysis of $K,N_{C}$
+Our proposed AIM-CoT incorporates the design of AVP module.
+In contrast to existing research, AIM-CoT, benefiting from the AVP module, does not simply select the top-$K$ regions with the highest attention scores from the attention map.
+Instead, it meticulously selects $K$ regions from a total set $C$ of $N_{C}$ candidate regions to construct the multimodal CoT ($N_{C}=N+M$).
+However, this approach may inevitably raise concerns regarding the deployability of AIM-CoT, particularly as the hyper-parameters $N_{C}$ and $K$ increase.
+To investigate this, we conduct experiments to explore the average processing time per instance for AIM-CoT with larger values of $N_{C}$ and $K$.
+For the experimental setup, we implement AIM-CoT with both Chameleon-7B and Qwen2.5-VL-32B backbones on M3CoT and LLaVA-W datasets under different combinations of $N_{C}$ and $K$.
+The results are presented in Table 12.
+From the results, the following two key insights can be derived:
+  • Insensitivity to the growth of the candidate set size $N_{C}$: observing any row with a fixed $K$ (e.g., $K=3$ on M3CoT with Chameleon-7B), as $N_{C}$ increases from 3 to 7, the total processing time rises from 13.18s to 13.42s, a marginal increase of only 0.24s.
+This implies that each additional candidate region introduces an average overhead of less than 0.06s.
+This strongly demonstrates that the performance of the AVP module does not degrade sharply with a moderate expansion of the candidate pool, indicating excellent scalability.
+  • Diminishing marginal cost with the increase in the number of selections $K$: considering a fixed column for $N_{C}$ (e.g., $N_{C}=7$ on LLaVA-W with Chameleon-7B), as $K$ increases from $1\rightarrow 3$, $3\rightarrow 5$, and $5\rightarrow 7$, the processing time increases by 0.72s, 0.34s, and 0.24s, respectively, which shows a notable decrease in incremental cost.
+This suggests that the primary computational overhead of the AVP algorithm lies in the initiation of the iterative search (the jump from $K=1$ to $K=3$).
+Once the iteration begins, the cost of subsequent selection steps is remarkably low, benefiting from efficient mechanisms such as the batch processing and KV Cache.
+Based on these key insights, our proposed AVP module demonstrates high computational efficiency and robustness when faced with increased computational complexity (i.e., larger $N_{C}$ and $K$ values).
+Figure 6: Qualitative comparison on ScienceQA (Question 244) with Chameleon-7B.
+Left (Blue): AIM-CoT actively selects regions based on Information Gain (IG).
+It correctly focuses on the distance (Step 1) and magnet size (Step 2), which are the actual visual variables required to solve the physics question.
+Right (Yellow): The baseline ICoT, relying on raw Attention Scores (AS), fails to filter out noise.
+It attends to irrelevant text (“Pair 1”) and decorative letters (“N/S”) rather than the physical attributes of the objects, resulting in reasoning failure.
+Appendix H Qualitative Case Study: Information Gain vs. Attention
+To complement the quantitative analyses in Sections 5.5 and 5.6, we provide a step-by-step visualization of the reasoning process to demonstrate the efficacy of our proposed AVP module.
+Figure 6 presents a qualitative comparison between AIM-CoT (Ours) and the baseline ICoT on a challenging question from ScienceQA (Question ID: 244) using the Chameleon-7B Team (2024) backbone.
+As observed, the Information Gain-driven selection in AIM-CoT successfully identifies task-relevant visual evidence (e.g., magnet size and distance) that directly contributes to the correct reasoning chain.
+In contrast, the Attention-driven baseline is distracted by high-frequency but irrelevant features, such as the text labels (“Pair 1”) or high-contrast letters (“N”, “S”), leading to hallucinated reasoning and an incorrect answer.
+This visual evidence reinforces our hypothesis that high-attention regions do not necessarily equate to high-value information for reasoning.
+Appendix I Comprehensive Analysis of AVP Component
+In this appendix, we conduct a comprehensive analysis of the AVP component to validate its design choices and theoretical foundations. The detailed analyses are organized as follows:
+  • Appendix I.1 compares different strategies for constructing the candidate set to identify the most effective composition.
+  • Appendix I.2 quantifies the specific contribution of the exploratory set by analyzing the source distribution of the selected regions.
+  • Appendix I.3 evaluates the robustness of AVP against adversarial attention noise, demonstrating its intrinsic noise-rejection capability.
+  • Appendix I.4 provides a theoretical justification for (1) the approximate submodularity of the information gain function and (2) the optimality of our greedy selection algorithm.
+I.1 Construction of Candidate Set $C$: Method Comparison
+In this section, we investigate the influence of different compositions of the total set $C$ on the performance of our proposed AIM-CoT. We specifically examine two primary configurations:
+Constructing $C$ using only $C_{attn}$ or $C_{exp}$.
+For the latter, we evaluate three distinct construction methodologies for $C_{exp}$: (a) $C_{rand}$: uniform random sampling; (b) $C_{ss}$: the selective search algorithm Uijlings et al.
+(2013), which is the seminal region proposal method utilized in R-CNN Girshick et al.
+(2014); (c) $C_{fsam}$: FastSAM Zhao et al.
+(2023), a computationally efficient variant of the foundational vision segmentation model, SAM Kirillov et al.
+(2023).
+Constructing $C$ using both $C_{attn}$ and $C_{exp}$. Similarly, as for $C_{exp}$, we also consider its diversified construction, including $C_{rand},C_{ss}$, and $C_{fsam}$.
+It is worth noting that although the construction is diverse, the size of $C$ remains consistent.
+When $C$ is composed of $C_{attn}$ and $C_{exp}$, the two each account for half.
+The experimental results are shown in Table 13.
+As observed, the combination of the two sets (i.e., $C=C_{attn}\cup C_{exp}$) invariably yields superior performance for AIM-CoT compared to configurations where either $C=C_{attn}$ or $C=C_{exp}$ is used exclusively.
+This highlights the importance of diversifying the sources of candidate visual regions.
+When comparing the different construction methods for $C_{exp}$, the performance gap among models is marginal when used in conjunction with $C_{attn}$.
+Specifically, despite its simplicity, random sampling achieves highly competitive results, which motivates our choice to adopt it as the default method for constructing $C_{exp}$.
+Intuitively, the advantage of random sampling lies in its ability to provide regions across different parts of the image unbiasedly with maximal spatial diversity.
+Table 13: Performance comparison of AIM-CoT variants on the basis of different constructions of the candidate set $C$.
+Construction of $C$ M3CoT (ACC.) LLaVA-W (ROUGE-L)
+$C_{attn}\cup C_{rand}~~~(C_{exp}=C_{rand})$ 31.4 29.8
+$C_{attn}\cup C_{ss}~~~(C_{exp}=C_{ss})$ 31.2 29.5
+$C_{attn}\cup C_{fsam}~~~(C_{exp}=C_{fsam})$ 31.0 29.6
+$C_{attn}$ 30.8 28.9
+$C_{rand}$ 30.4 28.6
+$C_{ss}$ 30.3 28.7
+$C_{fsam}$ 29.9 27.7
+I.2 Source Distribution Analysis: Quantifying Contribution of Exploratory Set $C_{exp}$
+In this section, we examine the distribution of sources for the visual regions selected by the AVP module of AIM-CoT.
+These regions are drawn from two sets, $C_{attn}$ and $C_{exp}$, with their respective selection proportions denoted as $P_{attn}$ and $P_{exp}$.
+Intuitively, $P_{exp}$ reflects the significance of incorporating the exploratory set $C_{exp}$ to construct a better multimodal CoT.
+A larger value of $P_{exp}$ indicates that the exploratory set $C_{exp}$ makes a greater contribution by providing informative salient regions to AIM-CoT, and vice versa.
+Experimental Setup The experiments are conducted on the M3CoT and LLaVA-W benchmarks.
+Our proposed AIM-CoT is implemented with the Chameleon-7B backbone under a default 0-shot setting.
+To ensure the reliability of the results, we repeat each experiment three times on both benchmarks.
+Results and Analysis As presented in Table 14, although the value of $P_{exp}$ fluctuates across different experimental runs on the same benchmark, it remains consistently around 20% on M3CoT and 30% on LLaVA-W.
+This indicates that the influence of stochastic factors on the source distribution of the selected regions is limited, which validates our rationale of using this metric as a reflection of the relative importance of $C_{attn}$ and $C_{exp}$.
+Furthermore, we observe that $P_{exp}$ is significantly greater than zero.
+This demonstrates that the exploratory set $C_{exp}$ consistently serves as a critical component of the total candidate set $C$, contributing a substantial portion of the informative regions for AIM-CoT.
+Table 14: Proportion of salient regions selected by the AVP module of our proposed AIM-CoT from the exploratory set $C_{exp}$.
+Experiment Number 1 2 3
+M3CoT 17.25% 20.44% 27.27%
+LLaVA-W 31.33% 25.77% 26.67%
+I.3 Robustness of AVP Against Attention Noise
+In Section 4.3, we posit that AVP possesses an intrinsic noise-rejection capability, formalized in Equation 4, allowing it to filter out high-attention regions that yield negligible information gain.
+To empirically validate this resilience against error propagation (e.g., when upstream modules produce misleading attention maps), we conduct an adversarial noise injection experiment.
+Experimental Setup.
+We design an Adversarial Noise Injection protocol.
+For each image, we randomly select an irrelevant background patch (verified to have no overlap with ground-truth objects) and artificially override its attention score to be the global maximum in the attention map.
+We then compare the region selection behavior and downstream performance of two strategies using this corrupted attention map: (1) Top-K Selection (Baseline): Selects regions strictly based on attention scores.
+(2) AVP (Ours): Selects regions based on Information Gain (IG).
+Consistent with Appendix L.1, we evaluate on M3CoT, ScienceQA, and LLaVA-W using Chameleon-7B and Qwen2-VL-7B backbones.
+Evaluation Metrics.
+(1) Noise Rejection Rate (NRR): The percentage of instances where the model successfully avoids selecting the injected noise patch, despite it having the highest attention score.
+(2) Performance Drop ($\Delta$): The decline in task performance compared to the clean (non-adversarial) setting.
+Smaller $\Delta$ indicates higher robustness.
+Table 15: Robustness analysis against adversarial attention noise.
+NRR denotes Noise Rejection Rate ($\uparrow$ higher is better); $\Delta$ denotes performance drop ($\downarrow$ lower is better, closer to 0).
+Backbone Method M3CoT ScienceQA LLaVA-W
+NRR ($\%$) $\uparrow$ $\Delta$ Acc. $\downarrow$ NRR ($\%$) $\uparrow$ $\Delta$ Acc. $\downarrow$ NRR ($\%$) $\uparrow$ $\Delta$ ROUGE-L $\downarrow$
+Chameleon-7B Top-K 0.0 -4.2 0.0 -3.8 0.0 -5.1
+AVP (Ours) 96.4 -0.2 98.1 -0.1 95.5 -0.4
+Qwen2-VL-7B Top-K 0.0 -3.5 0.0 -2.9 0.0 -4.4
+AVP (Ours) 97.8 -0.1 99.2 0.0 96.7 -0.2
+As shown in Table 15, the contrast between the two methods is stark:
+First, Top-K is fundamentally vulnerable to attention errors.
+By definition, Top-K has an NRR of 0.0%, as it blindly trusts the manipulated attention scores.
+Consequently, it forces the inclusion of irrelevant visual noise, leading to significant performance degradation across all benchmarks (e.g., dropping 4.2% accuracy on M3CoT with Chameleon-7B).
+Second, AVP demonstrates exceptional noise immunity.
+Even when the noise patch is forced to have the highest attention priority, AVP successfully rejects it in over 95% of cases (NRR $95\%$).
+This confirms that AVP correctly identifies that the noise patch, despite its high attention score, provides minimal reduction in uncertainty (low IG).
+As a result, the performance drop is negligible ($\Delta\approx 0$), proving that AVP effectively breaks the chain of error propagation, ensuring robust reasoning even when the guidance from the attention map is unreliable.
+I.4 Theoretical Justification: Approximate Submodularity and Optimality of Greedy Selection
 To offer a more comprehensive insight into the motivation for employing a greedy algorithm, this section provides a thorough analysis.
 We emphasize that for functions that are not theoretically submodular, a greedy approach remains one of the conventional methods for addressing their maximization, as established in recognized works Bian et al.
-(2017); Sener & Savarese (2018); Kim et al.
+(2017); Sener and Savarese (2018); Kim et al.
 (2016); Krause et al.
-(2008); Das & Kempe (2011).
+(2008); Das and Kempe (2011).
 In this part, we conduct meticulously designed experiments to investigate the extent to which the information gain function $F$ approximates submodularity.
@@
 This finding motivates us to follow established works Bian et al.
-(2017); Sener & Savarese (2018); Kim et al.
+(2017); Sener and Savarese (2018); Kim et al.
 (2016); Krause et al.
-(2008); Das & Kempe (2011) to propose a greedy algorithm to solve the maximization problem for $F$.
+(2008); Das and Kempe (2011) to propose a greedy algorithm to solve the maximization problem for $F$.
 The analysis is detailed as follows:
@@
 $f(A\cup\{R_{i}\})-f(A)\geq f(B\cup\{R_{i}\})-f(B)$ (12)
-for any sets $A\subseteq B\subset C$ and any element that satisfies $R_{i}\in C\backslash B$. In our scenario, the Inequality 15 is written as
+for any sets $A\subseteq B\subset C$ and any element that satisfies $R_{i}\in C\backslash B$. In our scenario, the Inequality 12 is written as
 $F(A\cup\{R_{i}\})-F(A)\geq F(B\cup\{R_{i}\})-F(B)$ (13)
@@
 Next, to compute the information gain contributed by a given region, we randomly sample a region $R_{test}$ from $C\setminus S_{large}$.
-We then calculate the VLM’s information content, which are denoted as $U_{s},U_{s}^{*},U_{l},\text{and }U_{l}^{*}$, when the context incorporates (1) $S_{small}$, (2) $S_{small}\cup\{R_{test}\}$, (3) $S_{large}$, and (4) $S_{large}\cup\{R_{test}\}$, respectively.
+We then calculate the VLM’s information contents, which are denoted as $U_{s},U_{s}^{*},U_{l},\text{and }U_{l}^{*}$, when the context incorporates (1) $S_{small}$, (2) $S_{small}\cup\{R_{test}\}$, (3) $S_{large}$, and (4) $S_{large}\cup\{R_{test}\}$, respectively.
 We expect to observe in the majority of cases that:
@@
 The experimental results are presented in Table 16.
-As we can see, the Inequality 15 holds in most instances across all settings and datasets.
+As we can see, the Inequality 12 holds in most instances across all settings and datasets.
 Furthermore, to confirm the significance of the obtained results, we introduce the Binomial Test, an exact statistical procedure for assessing the extent to which experimental outcomes with a binary structure are attributable to chance alone.
@@
 This demonstrates that the information gain function $F$ behaves in a manner that is empirically near-submodular, which motivates us to follow existing research Bian et al.
-(2017); Sener & Savarese (2018); Kim et al.
+(2017); Sener and Savarese (2018); Kim et al.
 (2016); Krause et al.
-(2008); Das & Kempe (2011) where greedy algorithms are proposed to solve the problem of maximizing approximately submodular functions.
+(2008); Das and Kempe (2011) where greedy algorithms are proposed to solve the problem of maximizing approximately submodular functions.
 Table 16: Proportions of instances on M3CoT and LLaVA-W benchmarks for which the approximate submodularity of information gain function $F$ is manifested.
-The backbone is Chameleon-7B and the model is our proposed AIMCoT.
+The backbone is Chameleon-7B and the model is our proposed AIM-CoT.
 $K_{large}$ is set to $K_{small}+1$ for simplicity.
@@
 p-value 0.0462 0.0031 0.0462 0.0249
-Appendix I Analysis of the Selected Regions’ Source
-In this section, we examine the distribution of sources for the visual regions selected by the AVP module of AIMCoT.
-These regions are drawn from two sets, $C_{attn}$ and $C_{exp}$, with their respective selection proportions denoted as $P_{attn}$ and $P_{exp}$.
-Intuitively, $P_{exp}$ reflects the significance of incorporating the exploratory set $C_{exp}$ to construct a better multimodal CoT.
-A larger value of $P_{exp}$ indicates that the exploratory set $C_{exp}$ makes a greater contribution by providing informative salient regions to AIMCoT, and vice versa.
-Experimental Setup The experiments are conducted on the M3CoT and LLaVA-W benchmarks.
-Our proposed AIMCoT is implemented with the Chameleon-7B backbone under a default 0-shot setting.
-To ensure the reliability of the results, we repeat each experiment three times on both benchmarks.
-Results and Analysis As presented in Table 7, although the value of $P_{exp}$ fluctuates across different experimental runs on the same benchmark, it remains consistently around 20% on M3CoT and 30% on LLaVA-W.
-This indicates that the influence of stochastic factors on the source distribution of the selected regions is limited, which validates our rationale of using this metric as a reflection of the relative importance of $C_{attn}$ and $C_{exp}$.
-Furthermore, we observe that $P_{exp}$ is significantly greater than zero.
-This demonstrates that the exploratory set $C_{exp}$ consistently serves as a critical component of the total candidate set $C$, contributing a substantial portion of the informative regions for AIMCoT.
-Table 7: Proportion of salient regions selected by the AVP module of our proposed AIMCoT from the exploratory set $C_{exp}$.
-Experiment Number 1 2 3
-M3CoT 17.25% 20.44% 27.27%
-LLaVA-W 31.33% 25.77% 26.67%
-Appendix J Deployment of AIMCoT
-J.1 Analysis of the Complexity of AVP Module
+Appendix J Extensive Ablation Studies on More Backbones
+Section 5.3 presents an ablation study on the Chameleon-7B backbone. The results provide two key insights as follows:
+  • Each component (CAG, AVP, DAT) contributes substantially to the overall performance.
+  • Even without CAG, combining AVP and DAT is sufficient for AIM-CoT to outperform strong baseline models by a clear margin.
+This finding highlights the importance of addressing the “what to see” and “when to see it” questions.
+To test whether these insights generalize beyond Chameleon-7B, in this section, we extend the same ablation study to two additional VLM backbones: Janus-Pro-7B and Qwen2.5-VL-32B.
+As shown in Tables 17 and 18, the ablation trends are consistent with those in Section 5.3: removing any component leads to performance degradation across datasets and backbones, demonstrating the efficacy of our proposed components.
+More importantly, the results highlight the central thesis of this study.
+Even without CAG, the combination of AVP and DAT already yields strong performance and is sufficient to surpass strong baselines, including ICoT.
+This demonstrates that explicitly and effectively addressing what to see (AVP) and when to see it (DAT) is the key driver of AIM-CoT’s advantage.
+Meanwhile, CAG remains a non-trivial contributor: although its absolute gains are comparatively smaller, it consistently improves results across all settings, suggesting that mitigating the text-vision granularity mismatch further complements AVP+DAT.
+Table 17: Ablation study of AIM-CoT conducted on Janus-Pro-7B under 0-shot setting.
+Dataset AIM-CoT w/o CAG w/o AVP w/o DAT
+M3CoT (ACC.) 39.7 39.4 (-0.3) 38.6 (-1.1) 39.0 (-0.7)
+ScienceQA (ACC.) 56.9 56.7 (-0.2) 55.8 (-1.1) 56.3 (-0.6)
+LLaVA-W (ROUGE-L) 35.5 34.7 (-0.8) 33.5 (-2.0) 34.5 (-1.0)
+Table 18: Ablation study of AIM-CoT conducted on Qwen2.5-VL-32B under 0-shot setting.
+Dataset AIM-CoT w/o CAG w/o AVP w/o DAT
+M3CoT (ACC.) 58.7 58.4 (-0.3) 57.6 (-1.1) 58.1 (-0.6)
+ScienceQA (ACC.) 76.8 76.6 (-0.2) 75.9 (-0.9) 76.3 (-0.5)
+LLaVA-W (ROUGE-L) 46.5 46.1 (-0.4) 44.2 (-2.3) 45.0 (-1.5)
+Appendix K In-depth Analysis of Performance Gain Variations Across Backbones
+In Table 1, we observe that AIM-CoT yields varying degrees of improvement across different base models.
+Specifically, the gains on Chameleon-7B are substantially larger than those on Qwen2-VL-7B, while Qwen2.5-VL-32B exhibits higher improvements than its 7B counterpart.
+In this section, we analyze the underlying factors contributing to these phenomena from two perspectives: visual encoding mechanisms and model capability scaling.
+K.1 Cross-Family Analysis: Architecture and Visual Encoding
+The distinct performance gap between Chameleon-7B (up to 18.25% gain on LLaVA-W) and Qwen2-VL-7B (up to 6.14% gain) can be attributed to their fundamental differences in visual processing:
+  • Chameleon-7B (Early-Fusion & Fixed Resolution): As an early-fusion model utilizing a fixed tokenization strategy, Chameleon-7B often struggles with fine-grained visual details in high-resolution images, leading to the “myopic” attention behavior discussed in Section 3.1.
+AIM-CoT acts as a perceptual correction mechanism here.
+By actively cropping and zooming in on salient regions via the AVP module, AIM-CoT directly compensates for the model’s native resolution limitations.
+The improvement is transformative as it solves a “can’t see” problem, resulting in substantial gains.
+  • Qwen2-VL Series (NaViT & Dynamic Resolution): These models employ the NaViT architecture with dynamic resolution support, allowing them to process arbitrary aspect ratios and resolutions natively.
+Since Qwen2-VL can already perceive visual details relatively clearly, the marginal utility of “zooming in” is lower compared to Chameleon.
+For Qwen models, AIM-CoT serves more as an attention optimization tool rather than a vision repair tool, leading to more moderate base improvements.
+K.2 Intra-Family Analysis: Capability Dependence and Scaling Laws
+Within the Qwen family, we observe that the larger Qwen2.5-VL-32B benefits more from AIM-CoT (up to 9.84% gain) than the smaller Qwen2-VL-7B (up to 6.14% gain).
+This counter-intuitive trend (where a stronger baseline yields larger relative gains) highlights the capability-dependent nature of our framework:
+  • Quality of CAG Generation: AIM-CoT is a closed-loop system where the Context-enhanced Attention-map Generation (CAG) module relies on the model’s own capability to describe the image.
+The 32B model generates more precise, hallucination-free descriptions than the 7B model.
+This high-quality context leads to more accurate attention maps, enabling the AVP module to select significantly more valuable image regions.
+  • Interleaved Reasoning Capability: The efficacy of AIM-CoT also depends on how well the model can leverage the inserted visual patches.
+Larger models (32B) possess stronger long-context understanding and multi-step reasoning capabilities.
+They can effectively synthesize the fragmented information provided by the inserted patches to deduce the correct answer, whereas smaller models might struggle to integrate these additional visual cues coherently.
+In summary, AIM-CoT functions as a mechanism for Perceptual Correction on weaker architectures (Chameleon) and Reasoning Augmentation on stronger ones (Qwen 32B), demonstrating its adaptability across different model paradigms.
+Appendix L Safety Mechanisms in AIM-CoT
+To ensure the reliability of the reasoning chain and prevent error propagation, AIM-CoT incorporates explicit safety mechanisms at two critical stages.
+First, regarding Source Quality (Input), we apply Negative Constraints within the CAG module to strictly ground the generated context and mitigate hallucinations.
+Second, regarding Integration Safety (Process), we employ Safety Instructions during the visual insertion, guiding the model to verify relevance and filter out potential noise.
+In this section, we analyze the individual contributions of these safeguards to the framework’s overall robustness.
+L.1 Effectiveness of Negative Constraints in CAG Generation
+In Section 4.2, we introduce negative constraints within the CAG prompt $\mathcal{P}_{CAG}$ to ensure the utmost precision of the generated descriptions.
+To rigorously assess the necessity of this design and the potential risk of error propagation, we conduct an ablation study focusing on the hallucination rate of the generated descriptions $\mathcal{D}_{CAG}$ and their impact on downstream performance.
+Experimental Setup.
+We contrast the standard AIM-CoT against AIM-CoT w/o Constraints, a variant where the cautionary instructions are removed from $\mathcal{P}_{CAG}$.
+To demonstrate the generalizability of our design, we implement the framework on two representative backbones: Chameleon-7B and Qwen2-VL-7B.
+The experiments are conducted on M3CoT, ScienceQA, and LLaVA-W benchmarks under the 0-shot setting.
+Evaluation Metrics.
+Two metrics are incorporated for evaluation: (1) Hallucination Rate (HR): We employ GPT-4v as an external evaluator to identify factual inconsistencies between the image content and the generated description $\mathcal{D}_{CAG}$.
+HR is reported as the percentage of descriptions containing details unsupported by the visual input.
+(2) Downstream Performance: We report Accuracy (ACC.) for M3CoT and ScienceQA, and ROUGE-L for LLaVA-W to measure the final reasoning outcome.
+Table 19: Ablation study on the effectiveness of Negative Constraints across two backbones.
+HR denotes Hallucination Rate ($\downarrow$ lower is better); Perf.
+denotes downstream performance ($\uparrow$ higher is better).
+Backbone Method M3CoT ScienceQA LLaVA-W
+HR ($\%$) $\downarrow$ Acc. $\uparrow$ HR ($\%$) $\downarrow$ Acc. $\uparrow$ HR ($\%$) $\downarrow$ ROUGE-L $\uparrow$
+Chameleon-7B w/o Constraints 5.8 30.1 4.5 52.5 8.4 27.9
+AIM-CoT (Ours) 1.9 31.4 1.2 53.1 2.7 29.8
+Qwen2-VL-7B w/o Constraints 3.5 43.9 2.8 56.9 5.1 35.1
+AIM-CoT (Ours) 0.8 44.7 0.5 57.4 1.5 36.3
+The results are summarized in Table 19, from which two key observations are derived:
+First, the CAG module exhibits high intrinsic reliability across backbones.
+Even without explicit negative constraints, the hallucination rates remain relatively low (e.g., $6\%$ on M3CoT for both models).
+This suggests that the foundational strategy of using VLM-generated descriptions as context is robust and not inherently prone to generating misleading information, regardless of the underlying model architecture.
+Second, negative constraints provide a critical layer of safety.
+Despite the solid baseline, the introduction of negative constraints consistently suppresses residual hallucinations across all datasets and backbones.
+Notably, for Chameleon-7B on M3CoT, the constraints reduce the hallucination rate by approximately 67% (from 5.8% to 1.9%).
+This demonstrates that while the models are generally capable, the constraints effectively filter out subtle noise and “over-interpretations.” By driving the hallucination rate down to a negligible level, we ensure that the subsequent AVP module operates on a virtually noise-free foundation, thereby maximizing overall reasoning performance.
+L.2 Effectiveness of Safety Instructions in DAT Integration
+In Section 4.4, we introduce a robust integration strategy within the DAT module, utilizing specific Safety Instructions (e.g., instructing the model to treat inserted regions as “supplementary references” and verify semantic consistency) to mitigate the risk of integrating irrelevant or hallucinated visual cues.
+To validate the efficacy of this textual safeguard, we conduct a stress test involving Irrelevant Visual Injection.
+Experimental Setup.
+We simulate a worst-case scenario where the upstream selection process fails completely.
+Whenever the DAT triggers a visual insertion, instead of inserting the semantically relevant region selected by AVP, we deliberately inject a random, irrelevant image patch from a different dataset sample.
+We then measure the model’s resilience under this interference using two configurations: (1) Naive Injection (w/o Safety): The irrelevant patch is inserted directly into the reasoning chain without any accompanying safety prompts.
+(2) AIM-CoT Integration (w/ Safety): The irrelevant patch is inserted accompanied by our proposed safety instructions.
+We employ the same backbones (Chameleon-7B, Qwen2-VL-7B) and benchmarks (M3CoT, ScienceQA, LLaVA-W) as in previous sections.
+Evaluation Metrics.
+We report the Performance under Noise and the Performance Drop ($\Delta$) relative to the clean, standard AIM-CoT performance (where accurate regions are inserted).
+A smaller magnitude of $\Delta$ indicates greater robustness against misleading visual information.
+Table 20: Ablation study on Safety Instructions under irrelevant visual injection.
+$\Delta$ indicates the performance drop compared to the standard AIM-CoT ($\downarrow$ lower drop/magnitude is better).
+Backbone Method M3CoT ScienceQA LLaVA-W
+Acc. $\Delta$ Acc. $\Delta$ ROUGE-L $\Delta$
+Chameleon-7B Naive Injection (w/o Safety) 24.5 -6.9 46.2 -6.9 22.1 -7.7
+AIM-CoT (w/ Safety) 29.9 -1.5 51.8 -1.3 28.4 -1.4
+Qwen2-VL-7B Naive Injection (w/o Safety) 38.2 -6.5 50.9 -6.5 30.5 -5.8
+AIM-CoT (w/ Safety) 43.5 -1.2 56.1 -1.3 35.1 -1.2
+The results in Table 20 provide compelling evidence for the necessity of safety instructions:
+First, unprotected visual insertion is highly risky.
+When irrelevant patches are naively injected, the models suffer severe performance degradation (e.g., $\sim$7% drop on M3CoT for Chameleon-7B).
+This confirms that without guidance, VLMs tend to over-trust visual tokens, attempting to reason over noise and consequently hallucinating or diverging from the correct logic path.
+Second, Safety Instructions act as an effective semantic firewall.
+By simply instructing the model to verify consistency, the performance drop is drastically reduced to a marginal level (e.g., only -1.5% on M3CoT).
+This demonstrates that the safety instruction successfully cues the model to exercise critical judgment: when the inserted visual evidence conflicts with the textual context (as in this random injection case), the model chooses to disregard the visual noise and rely on its internal knowledge, thereby maintaining robust performance even in the presence of upstream errors.
+Appendix M Deployment of AIM-CoT
+M.1 Analysis of the Complexity of AVP Module
 Overview of the AVP Module
-The Active Visual Probing (AVP) module’s primary purpose is to dynamically and intelligently select salient sub-regions of an image during the text generation process.
+The AVP module’s primary purpose is to dynamically and intelligently select salient sub-regions of an image during the text generation process.
 This is achieved by calculating the “information gain” that each potential sub-region offers, thereby allowing the model to “zoom in” on relevant visual details and generate more informed and contextually aware text.
@@
 It iteratively evaluates candidate regions and selects the combination that maximizes the reduction in uncertainty (entropy) for the next token prediction.
-Definition of the Notations
+Definition of Notations
 Let’s define the key variables that will be used in the complexity analysis:
@@
 This is because the AVP module is highly efficient at extracting crucial visual information from the candidate set.
-Our empirical results demonstrate that AIMCoT achieves exceptional performance (as shown in Table 2) even when these hyper-parameters are kept at low levels (e.g., $K=3,N_{C}=6$, which is the default setting).
+Our empirical results demonstrate that AIM-CoT achieves exceptional performance (as shown in Table 1) even when these hyper-parameters are kept at low levels (e.g., $K=3,N_{C}=6$, which is the default setting).
 In conclusion, we can approximate the complexity of AVP as follows:
@@
 As a comparison, the attention-driven selection method within the baseline model ICoT is with a complexity of $O(logK\cdot\Delta N+L\cdot N\cdot H)$, which also scales linearly with sequence length N.
-Although this is acknowledged to be lower than the complexity of AVP shown in Equation 19, our empirical results detailed in Appendix J.2 suggest that AVP’s average inference time is no more than 1.36 times that of this method, while achieving performance far superior to it as analyzed in Section 5.5.
-J.2 Empirical Analysis of the Deployment of AIMCoT
-In this section, we empirically investigate the deployability of the AIMCoT framework and the temporal overhead introduced by the AVP module.
-For the experimental setup, we utilized Chameleon-7B as the backbone in a 0-shot setting to compare the average inference time of AIMCoT against ICoT Gao et al.
+Although this is acknowledged to be lower than the complexity of AVP shown in Equation 16, our empirical results detailed in Appendix M.2 suggest that AVP’s average inference time is no more than 1.36 times that of this method, while achieving performance far superior to it as analyzed in Section 5.4.
+M.2 Empirical Analysis of the Deployment of AIM-CoT
+In this section, we empirically investigate the deployability of the AIM-CoT framework and the temporal overhead introduced by the AVP module.
+For the experimental setup, to ensure the reliability of the experimental results, we introduced two representative VLM backbones as the basis for the experiment: Chameleon-7B and Qwen2.5-VL-32B.
+They not only differ significantly in the number of parameters, but also employ different early-fusion and late-fusion strategies, respectively.
+In the 0-shot setting, we compare the average inference time of AIM-CoT against ICoT Gao et al.
 (2025).
-ICoT, as a key baseline model in this study, employs a Top-K strategy to simultaneously select regions with the highest attention scores for constructing the multimodal CoT, thereby expected to possess a relatively lower time complexity compared to AIMCoT.
-Consequently, the comparison with ICoT serves as a direct indicator of AIMCoT’s deployability.
-Table 8: Comparison of the average time to process each instance between AIMCoT and the baseline model (ICoT).
-Dataset AIMCoT ICoT
-M3CoT 13.37s 11.62s
-LLaVA-W 11.65s 8.58s
+ICoT, as a key baseline model in this study, employs a Top-K strategy to simultaneously select regions with the highest attention scores for constructing the multimodal CoT, and is therefore expected to possess a relatively lower time complexity compared to AIM-CoT.
+Consequently, the comparison with ICoT serves as a direct indicator of AIM-CoT’s deployability.
+Table 21: Comparison of average inference time (seconds). Values in parentheses denote the relative time against ICoT.
+Dataset Method Chameleon-7B Qwen2.5-VL-32B
+M3CoT ICoT 11.62 51.21
+[…]
+LLaVA-W ICoT 8.58 36.95
+[…]
 The experimental results are presented in Table 21.
 Two key observations can be drawn.
-First, the AVP module does not introduce significant temporal costs to the AIMCoT framework, an efficiency attributable to batch processing and the KV Cache mechanism.
-Second, AIMCoT achieves substantially superior performance as shown in Table 2, at a time cost comparable to that of the efficient baseline, which is less than $1.36$ times that of ICoT (specifically, $1.15$ and $1.36$ times on M3CoT and LLaVA-W benchmarks, respectively).
-This suggests that our proposed AIMCoT framework achieves a favorable trade-off between performance and deployability.
-Appendix K Sensitivity Analysis of $\delta$
-The hyper-parameter $\delta$ within the DAT module serves as a crucial threshold to trigger the AVP module, which inserts salient visual regions to improve the construction of the multimodal CoT.
-In this section, we detail a sensitivity analysis of $\delta$ by adjusting $\delta$ across the range of $[0.1,0.125,0.15,0.175,0.2,0.225]$ and examining not only (1) the performance of AIMCoT, but also (2) the number of times the AVP is triggered.
-The experiments are conducted under 0-shot setting on Chameleon-7B backbone and LLaVA-W benchmark.
-The experimental results are shown in Figure 5.
-The left figure illustrates that AIMCoT exhibits limited performance when the threshold, $\delta$, is set too low.
-This underscores the importance of inserting visual information at critical moments: excessively frequent or inopportune visual insertions can disrupt the VLM’s reasoning process, leading to suboptimal performance.
-As $\delta$ increases, the model’s performance progressively improves, reaching its peak at $\delta=0.2$ (our default setting), which corresponds to a ROUGE-L score of 0.2983.
-However, a further increase in $\delta$ results in a slight performance degradation.
-This highlights the criticality of visual information insertion for constructing an interleaved Chain of Thought: an overly stringent threshold excessively impedes the incorporation of visual data, preventing the AVP from supplying the model with necessary visual supplementation in a timely manner.
-Conversely, the right figure demonstrates a consistent decrease in the number of times the AVP is triggered as $\delta$ is raised.
-This showcases the efficacy of $\delta$ as a threshold for modulating the activation frequency of the AVP.
-Figure 5: Experimental results of the sensitivity analysis of the hyper-parameter $\delta$.
-The left figure illustrates the performance of AIMCoT when $\delta$ takes different values, while the right one shows the number of times the AVP module within AIMCoT is triggered.
-Table 9: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on M3CoT benchmark, which illustrate the average time of AIMCoT to process an instance under different settings.
-Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
-$N_{C}$ 1 2 3 4 5 6 7
-K=1 12.78s 12.80s 12.83s 12.87s 12.90s 12.93s 12.97s
-K=3 - - 13.18s 13.23s 13.30s 13.37s 13.42s
-K=5 - - - - 13.41s 13.49s 13.56s
-K=7 - - - - - - 13.66s
-Table 10: Experimental results of sensitivity analysis of hyper-parameters $N_{C},K$ on LLaVA-W benchmark, which illustrate the average time of AIMCoT to process an instance under different settings.
-Since $N_{C},K$ inherently satisfies $N_{C}\geq K$, any entry corresponding to a setting that satisfies $N_{C}<K$ is filled with the symbol "-".
-$N_{C}$ 1 2 3 4 5 6 7
-K=1 10.61s 10.70s 10.75s 10.84s 10.92s 11.00s 11.08s
-K=3 - - 11.21s 11.36s 11.51s 11.65s 11.80s
-K=5 - - - - 11.79s 11.96s 12.14s
-K=7 - - - - - - 12.38s
-Appendix L Sensitivity Analysis of $K,N_{C}$
-Our proposed AIMCoT incorporates the design of AVP module.
-In contrast to existing research, AIMCoT, benefiting from the AVP module, does not simply select the top-$K$ regions with the highest attention scores from the attention map.
-Instead, it meticulously selects $K$ regions from a total set $C$ of $N_{C}$ candidate regions to construct the multimodal CoT ($N_{C}=N+M$).
-However, this approach may inevitably raise concerns regarding the deployability of AIMCoT, particularly as the hyper-parameters $N_{C}$ and $K$ increase.
-To investigate this, we conduct experiments to explore the average processing time per instance for AIMCoT with larger values of $N_{C}$ and $K$.
-For the experimental setup, we implement AIMCoT with the Chameleon-7B backbone on the M3CoT and LLaVA-W datasets under various combinations of $N_{C}$ and $K$.
-The results are presented in Tables 9 and 10, from which we derive two key insights:
-  • Insensitivity to the growth of the candidate set size $N_{C}$: observing any row with a fixed $K$ (e.g., $K=3$ on Table 9), as $N_{C}$ increases from 3 to 7, the total processing time rises from 13.18s to 13.42s, a marginal increase of only 0.24s.
-This implies that each additional candidate region introduces an average overhead of less than 0.06s.
-This strongly demonstrates that the performance of the AVP module does not degrade sharply with a moderate expansion of the candidate pool, indicating excellent scalability.
-  • Diminishing marginal cost with the increase in the number of selections $K$: considering a fixed column for $N_{C}$ (e.g., $N_{C}=7$ on Table 10), as $K$ increases from $1\rightarrow 3$, $3\rightarrow 5$, and $5\rightarrow 7$, the processing time increases by 0.72s, 0.34s, and 0.24s, respectively, which shows a notable decrease in incremental cost.
-This suggests that the primary computational overhead of the AVP algorithm lies in the initiation of the iterative search (the jump from $K=1$ to $K=3$).
-Once the iteration begins, the cost of subsequent selection steps is remarkably low, benefiting from efficient mechanisms such as the batch processing and KV Cache.
-Based on these key insights, our proposed AVP module demonstrates high computational efficiency and robustness when faced with increased computational complexity (i.e., larger $N_{C}$ and $K$ values).
-Table 11: Performance comparison AIMCoT variants on the basis of different constructions of the candidate set $C$.
-Construction of $C$ M3CoT (ACC.) LLaVA-W (ROUGE-L)
-$C_{attn}\cup C_{rand}~~~(C_{exp}=C_{rand})$ 31.4 29.8
-$C_{attn}\cup C_{ss}~~~(C_{exp}=C_{ss})$ 31.2 29.5
-$C_{attn}\cup C_{fsam}~~~(C_{exp}=C_{fsam})$ 31.0 29.6
-$C_{attn}$ 30.8 28.9
-$C_{rand}$ 30.4 28.6
-$C_{ss}$ 30.3 28.7
-$C_{fsam}$ 29.9 27.7
-Appendix M Ablation Study on the Construction of Candidate Set $C$
-In this section, we investigate the influence of different compositions of the total set $C$ on the performance of our proposed AIMCoT. We specifically examine two primary configurations:
-Constructing $C$ using only $C_{attn}$ or $C_{exp}$.
-For the latter, we evaluate four distinct construction methodologies for $C_{exp}$: (a) $C_{rand}$: uniform random sampling; (b) $C_{ss}$: the selective search algorithm Uijlings et al.
-(2013), which is the seminal region proposal method utilized in R-CNN Girshick et al.
-(2014); (c) $C_{fsam}$: FastSAM Zhao et al.
-(2023), a computationally efficient variant of the foundational vision segmentation model, SAM Kirillov et al.
-(2023).
-Constructing $C$ using both $C_{attn}$ and $C_{exp}$. Similarly, as for $C_{exp}$, we also consider its diversified construction, including $C_{rand},C_{ss}$, and $C_{fsam}$.
-It is worth noting that although the construction is diverse, the size of $C$ remains consistent.
-When $C$ is composed of $C_{attn}$ and $C_{exp}$, the two each account for half.
-The experimental results are shown in Table 11.
-As observed, the combination of the two sets (i.e., $C=C_{attn}\cup C_{exp}$) invariably yields superior performance for AIMCoT compared to configurations where either $C=C_{attn}$ or $C=C_{exp}$ is used exclusively.
-This highlights the importance of diversifying the sources of candidate visual regions.
-When comparing the different construction methods for $C_{exp}$, the performance gap among models is marginal when used in conjunction with $C_{attn}$.
-Specifically, despite its simplicity, random sampling achieves highly competitive results, which motivates our choice to adopt it as the default method for constructing $C_{exp}$.
-Intuitively, the advantage of random sampling lies in its ability to provide regions across different parts of the image unbiasedly with maximal spatial diversity.
+First, the AVP module does not introduce significant temporal costs to the AIM-CoT framework, an efficiency attributable to batch processing and the KV Cache mechanism.
+Second, AIM-CoT achieves substantially superior performance as shown in Table 1, at a time cost comparable to that of the efficient baseline, which is less than $1.36$ times that of ICoT.
+This suggests that our proposed AIM-CoT framework achieves a favorable trade-off between performance and deployability.